### v1

# Memorization to Generalization: Emergence of Diffusion Models from Associative Memory

Abstract
Hopfield networks are associative memory (AM) systems, designed for storing and retrieving patterns as local minima of an energy landscape. In the classical Hopfield model, an interesting phenomenon occurs when the amount of training data reaches its critical memory load — spurious states, or unintended stable points, emerge at the end of the retrieval dynamics, leading to incorrect recall. In this work, we examine diffusion models, commonly used in generative modeling, from the perspective of AMs. The training phase of diffusion model is conceptualized as memory encoding (training data is stored in the memory). The generation phase is viewed as an attempt of memory retrieval. In the small data regime the diffusion model exhibits a strong memorization phase, where the network creates distinct basins of attraction around each sample in the training set, akin to the Hopfield model below the critical memory load. In the large data regime, a different phase appears where an increase in the size of the training set fosters the creation of new attractor states that correspond to manifolds of the generated samples. Spurious states appear at the boundary of this transition and correspond to emergent attractor states, which are absent in the training set, but, at the same time, have distinct basins of attraction around them. Our findings provide: a novel perspective on the memorization-generalization phenomenon in diffusion models via the lens of AMs, theoretical prediction of existence of spurious states, empirical validation of this prediction in commonly-used diffusion models. denotes equal contribution.
1 Introduction
Hopfield networks are the simplest energy-based models that conceptualize memories as attractor states corresponding to local minima of their energy function, and the memory retrieval as dynamical convergence towards those attractor states [1, 2, 3]. Recently, they have seen a resurgence of interest due to advances in their memorization capacity. Notably, Dense Associative Memories (DenseAMs), which are extensions of Hopfield networks with super-linear memory capacity [4, 5] have paved the way for more sophisticated associative memory systems [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], driven in part by their connection to the attention mechanism in transformers [19, 20, 13] and to neurobiology [20, 21]. Simultaneously, generative diffusion models [22] have gained considerable popularity, due to their flexibility and accuracy in modeling high-dimensional distributions for a variety of domains —such as image generation [23, 24, 25], audio [26, 27, 28], video synthesis [29, 30, 31, 32], and other scientific applications. However, despite their effectiveness, diffusion models pose challenges related to privacy and security, as concerns grow about their tendency to replicate training data [33, 34, 35]. Such matters consequently emphasize the need for further understanding of memorization and generalization behaviors in diffusion models.
Recent works [36, 37, 38] have begun establishing theoretical connections between DenseAMs and generative diffusion models, offering a foundation for bridging the two fields. Specifically, it was shown that the logarithm of the probability of the generated samples in diffusion models can be interpreted as the energy function of a commonly used DenseAM model with the softmax activation function. This makes it possible to apply theoretical tools developed for associative memory to better understand the computational properties of diffusion models.
A possible key that can further bridge these two systems and highlight the properties of the memorization-to-generalization transition in diffusion models is the phenomenon of spurious patterns. Historically considered as detrimental to pattern recall [39, 40, 41], spurious patterns can be interpreted as combinations or interpolations of stored patterns, hinting at the network’s ability to synthesize new patterns from existing training data. This blending of fundamental memories resembles the generalization process in generative models, where learned representations are used to generate novel outputs. In this way, spurious patterns offer a fascinating framework for exploring the balance between memorization — where models store exact patterns from training data — and generalization, where they use underlying features to create genuinely new samples [42]. Studying the conditions under which spurious patterns emerge and how they contribute to the network’s behavior can thus shed light on generalization in both associative memory and contemporary generative models.
Previous works have explored memorization in generative models through various approaches. For instance, [43, 44] proposed general methods to measure memorization, while [33, 34, 45, 46] examined the memorization capacity in diffusion models as a function of the size of the training data. Other works have focused on understanding generalization, such as the theoretical estimates of the generalization gap [47], and the spectral analyses [48, 49, 50]. A recent study [51] proposed that generalization in diffusion models arises as a result of the combination of limited capacity and structural constraints neural network, which results in predictable patterns at the boundary between memorization and generalization. Other recent studies [52, 50, 49] explored memorization and generalization through the relationship between the dimensionalities of the ground-truth manifold and the manifold that is learned by the model. Meanwhile, [53] explored the evolution of the generated density as a function of the training time to provide an analysis on how mode collapse occurs and leads to the loss of diversity in the generated density. Finally, [54] and [55] characterized memorization in diffusion models as “condensation” phase transitions known in the statistical physics of spin glasses and random energy models. Although these studies shed some light on memorization and generalization, they do not fully explore the model’s behavior during the transition between these two phases.
Figure 1: A simple illustration depicting the change in the energy landscape as the size of the training dataset is increased. In the small data regime, the model memorizes the training data points as local minima of the energy. When the amount of training data exceeds the memory capacity of the model, spurious patterns are formed and training data points are no longer energy minima. Subsequent increase of the training set size leads to the generalization phase, which is defined by the formation of continuous manifold of the low energy states.
Contributions.   Most of the mentioned works on memorization-to-generalization transition tackled it in a generalization-centric fashion. Namely, the memorization phenomenon is typically viewed as a “small side effect” of the diffusion models. In our work we adopt a complimentary approach. From the beginning, we cast the diffusion modeling pipeline into the associative memory framework. The training phase of the diffusion model is conceptualized as writing into the memory operation. The generation phase is viewed as an attempt of memory recall. This recall can be successful, resulting in the retrieval of training samples (memorization), or unsuccessful, resulting in the generation of new previously unseen samples (generalization). Thus, according to this theory, generalization (creation of genuinely novel samples) is a failure of memory recall. This point of view allows us to apply theory developed in DenseAMs for analyzing the memorization-to-generalization transition.
In DenseAMs, it is well-known that successful memory recall deteriorates as the number of the stored data points increases [4]. Importantly, the transition from successful to unsuccessful memory recall is marked by the peculiar phenomenon — the emergence of spurious states, see Figure 1. Our theory provides a specific and empirically verifiable prediction — spurious states must exist in the conventional diffusion models, which are trained and ran using standard methods. In this paper we demonstrate that this is indeed the case. Spurious states emerge at the boundary between the memorization and generalization phases, see examples of spurious samples in Figure 2.
In accord with previous literature, we find that (1) diffusion models undergo a transition from memorization to generalization as the number of training points increases; (2) unlike previous works on diffusion models, at the onset of generalization we detect the emergence of spurious patterns, which have sizable basins of attraction, but are not memorized data points; and lastly, (3) we provide theoretical descriptions distinguishing these spurious states from memorized and generalized patterns in terms of energy landscapes. All of these findings suggest a close parallel between diffusion models and DenseAMs above the critical storage capacity.
Figure 2: Examples of memorized, spurious, and generalized samples in their respective columns for four datasets. For each target image (shown on the left), its top-4 nearest neighbors from the training set (top row) and the synthetic set (bottom row) are shown to highlight the novelty or the commonality of the target image with respect to training and synthetic sets. Memorized samples are duplicates of the training points. Spurious samples are copies of the synthetic set that do not appear in the training set. Generalized samples do not belong to either of the two sets, indicating that they are genuinely novel samples that are not generated frequently.
2 Diffusion Models and Dense Associative Memories
Given a dataset of i.i.d. samples ${\mathbf{y}}\in\mathbb{R}^{N}$ drawn from a target data distribution $p({\mathbf{y}})$, diffusion models are a class of generative models, which aims to approximate $p({\mathbf{y}})$ by placing a reversible process that maps data to noise and back. The mapping to noise (or forward process) is described by the following Itô Stochastic Differential Equation (SDE) [25]:
$\mathrm{d}{\mathbf{x}}_{t}=\mathbf{f}({\mathbf{x}}_{t},t)\mathrm{d}t+{g}(t)\mathrm{d}{\mathbf{w}}_{t}$ (1)
which transforms the given data distribution (${\mathbf{x}}_{0}={\mathbf{y}}$) into a simpler one, such as an isotropic Gaussian distribution. Here, ${\mathbf{w}}_{t}$ is the standard Wiener process (or Brownian motion) and $\mathbf{f}({\mathbf{x}}_{t},t)$ denotes the drift term that guides the diffusion process, which we will assume to be zero for the most part of this paper. Meanwhile, $g(t)$ represents the diffusion coefficient that controls the noise at each time step $t\rightarrow T$. The reverse process removes the injected noise and is described as
$\mathrm{d}{\mathbf{x}}_{t}=[\mathbf{f}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})]\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (2)
where $\bar{{\mathbf{w}}}_{t}$ is the standard Wiener process. To effectively solve this equation, one must reliably estimate the score $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$ via parameterization of the score as a neural network $s_{\theta}({\mathbf{x}},t)$, where $\theta$ denotes neural network’s trainable parameters. This network is trained using methods for denoising score matching across multiple times steps [23, 25, 56, 57, 58].
Our theoretical arguments pertain to the idealistic setting in which this neural network is capable of learning the true empirical distribution obtained by adding noise to the training data. This limiting case is often adopted in theoretical studies of diffusion models, see [54, 59, 60, 55, 61]. From the practical perspective, this is achieved if the neural network is large (over-parametrized) and trained for a long time, an assumption that may deviate from the practical use cases. This simplification allows us to make a theoretical prediction that diffusion models must generate spurious states. It is highly non-trivial that this theoretical prediction can be extrapolated to the regime in which the size of the score-modeling neural network is finite and the amount of training data is large. The main contribution of our work is the empirical demonstration that this extrapolation is valid, and spurious states indeed exist in practically used settings.
Consider the training data distribution in the variance-exploding setting of $f({\mathbf{x}}_{t},t)=0$ and $g(t)=\sigma$. In this case, the marginal probability distribution of new samples can be computed exactly as
$p({\mathbf{x}}_{t},t)=\underset{{\mathbf{y}}\sim\text{data}}{\mathbb{E}}\Bigg{[}\frac{1}{(2\pi\sigma^{2}t)^{\frac{N}{2}}}\exp\Big(-\frac{\lVert{\mathbf{x}}_{t}-\mathbf{y}\rVert^{2}_{2}}{2\sigma^{2}t}\Big{missing})\Bigg{]}$ (3)
Assuming the empirical distribution of the data $p(\mathbf{y})=\frac{1}{K}\sum\limits_{\mu=1}^{K}\delta^{(N)}(\mathbf{y}-\bm{\xi}^{\mu})$, where $\bm{\xi}^{\mu}$ represents an individual data point (with data size $K$), this marginal distribution can be written as
[…]
where we also defined the energy $E^{\mathrm{DM}}$ of diffusion model, which up to state– or ${\mathbf{x}}$–independent terms is equal to
$E^{\text{DM}}({\mathbf{x}}_{t},t)=-2\sigma^{2}t\log[\sum\limits_{\mu=1}^{K}\exp\Big(-\frac{\lVert{\mathbf{x}}_{t}-\bm{\xi}^{\mu}\rVert^{2}_{2}}{2\sigma^{2}t}\Big{missing})\bigg{]}$ (5)
As already observed in [37], this energy function (5) is closely related to the class of models called DenseAMs, or Modern Hopfield Networks, which are large memory storage variants of classical Hopfield networks [4, 62].
The core idea behind DenseAMs is to design an energy function that sharply peaks around the intended memory patterns. These networks can store and retrieve a much larger number of memories, compared to the classical Hopfield networks, which scales super-linearly (and possibly exponentially) to the size of the network and decouples information storage capacity from the dimensionality of the data [4, 6]. Of particular interest here is the DenseAM model studied in [12] (see also [11]):
$E^{\text{AM}}({\mathbf{x}})=-\beta^{-1}\log[\sum\limits_{\mu=1}^{K}\exp\Big(-\beta\lVert{\mathbf{x}}-\bm{\xi}^{\mu}\rVert^{2}_{2}\Big{missing})\bigg{]}$ (6)
where $\beta$ is the inverse “temperature”, which controls the steepness of the energy landscape around memories $\bm{\xi}^{\mu}$. A direct comparison of the energy formulas (5) and (6) highlights a close connection between these two frameworks. Specifically, the data points in the diffusion model framework play the role of memories in the associative memory formulation. The variance of the noise added during the forward diffusion process (1) plays the role of the effective temperature $\beta^{-1}$. Finally, the reverse denoising process (2) corresponds to memory retrieval dynamics [36, 37].
3 Memorization-to-Generalization in 2D
To better illustrate this connection, it is instructive to investigate a simple 2-dimensional toy model, see Figure 3, which exhibits many aspects of the memorization-generalization transition. Imagine that the training data lies on a unit circle. We are interested in exploring how the shape of the energy function (6) changes as the number of training data points increases.
$\bm{[K=1]}$ For a single data point $\bm{\xi}^{1}$, the energy function (6) has a single local minimum at that stored pattern.
$\begin{split}E^{\text{AM}}({\mathbf{x}})&=-\beta^{-1}\log[\exp\Big(-\beta\lVert{\mathbf{x}}-\bm{\xi}^{1}\rVert^{2}_{2}\Big{missing})\Big{]}=\lVert{\mathbf{x}}-\bm{\xi}^{1}\rVert^{2}_{2}\end{split}$ (7)
The shape of the energy landscape is independent of the value of $\beta$.
Figure 3: Energy landscape evolution for the 2D toy model as training data size $K$ increases. Models trained at $K\in\{2,9,1000\}$ use standard VE-SDE based diffusion pipeline with training data sampled from the unit circle, shown in black for $K\in\{2,9\}$. Generated samples are shown alongside the learned score field $s_{\theta}({\mathbf{x}}_{t},t)$ done via a neural network, aligned with the negative gradient of the energy (5). Hierarchical clustering identifies structure within the generations, with cluster centroid energies visualized by color and numerical value. The rightmost panel shows the exact solution as $K\rightarrow\infty$ derived in Equation 12. As $K$ grows, the model initially memorizes individual data points, forming isolated basins. Around $K=9$, spurious patterns emerge — distinct low-energy attractors not present in the data — which mark the onset of generalization. At large $K$, the model enters a fully generalized regime, where low-energy states lie on a flat, continuous manifold shown in Figure 1.
$\bm{[K=2]}$ For two patterns, the energy is given by
$\begin{split}E^{\text{AM}}({\mathbf{x}})=-\beta^{-1}&\log[\exp\Big(-\beta\lVert{\mathbf{x}}-\bm{\xi}^{1}\rVert^{2}_{2}\Big{missing})+\exp\Big(-\beta\lVert{\mathbf{x}}-\bm{\xi}^{2}\rVert^{2}_{2}\Big{missing})\Big{]}\end{split}$ (8)
For $\beta\rightarrow\infty$, this energy has two local minima corresponding to the stored patterns. Meanwhile, for finite values of $\beta$, there exist configurations of patterns such that the energy has only one local minimum:
$\bm{\eta}=\operatorname*{arg\,min}_{{\mathbf{x}}}E^{\text{AM}}({\mathbf{x}})$ (9)
such that this local minimum is different from either of the two stored patterns: $\bm{\eta}\neq\bm{\xi}^{1}$ and $\bm{\eta}\neq\bm{\xi}^{2}$. This “emergent” local minimum is the spurious state, see also the cartoon illustration in Figure 1.
$\bm{[K\rightarrow\infty]}$ Finally, consider the case when the number of training data points is infinite. In this case they can be described by a continuous density of states:
$p({\mathbf{y}})=\frac{1}{\pi}\delta\big{(}y_{1}^{2}+y_{2}^{2}-1\big{)}$ (10)
The probability of the generated data is proportional (up to terms independent of the state ${\mathbf{x}}$) to
[…]
where $I_{0}(\cdot)$ is a modified Bessel function of the first kind, see Appendix A for more details on the derivation of Equation 11. Thus, the energy of the model is given by
$\begin{split}E^{\text{AM}}(R,\phi)&=R^{2}+1-\frac{1}{\beta}\log[I_{0}(2\beta R)\big{]}\underset{\beta\rightarrow\infty}{\approx}(R-1)^{2}\end{split}$ (12)
where $R$ is the radius of the unit circle and $\phi$ is the polar angle.
Notice, the dependence on the polar angle $\phi$ in Equation 12 completely disappears from the final result. The local minima of the resulting energy function form a continuous manifold, corresponding to $R=1$. Data samples from the model occupy the vicinity of that manifold. This behavior describes the fully generalized phase. In the limit of large $\beta$ the energy landscape is described by a parabola centered around $R=1$.
For a single training data point, this simple model exhibits memorization. With a few training data points, the model can either be in the memorization or spurious phase, depending on the location of those points and the value of $\beta$. For a very large number of training data points, the model is in the generalization phase. All effects perfectly align with the cartoon illustration in Figure 1.
Overall, this calculation entirely ignores the neural network, which is used to model the score and approximate the empirical energy. There exists a logical possibility that the introduction of the neural network would entirely change the progression of distinct behaviors: memorization, spurious, generalization. To test this, we trained a set of diffusion models in this precise setting. Each model is trained on a dataset of size $K$. The resulting energy landscapes are shown in Figure 3. For small number of data points, e.g., $K=2$, the diffusion model exhibits memorization. The local minima of the energy correspond to the training data points. Importantly, at $K=9$, we are able to observe the first signs of spurious states. At this stage, the model begins to learn emergent (different from the training data) local minima of the energy. Subsequent increase of the size of the training set leads to fully generalized behavior, which is illustrated for $K=1000$ data points. At that stage, all of the samples from the model live in close proximity of the exact data manifold. The right panel shows the analytical expression for the energy landscape, defined by Equation 12. Thus, the conventional diffusion modeling pipeline agrees very well with the theoretical prediction of the empirical energy. Please refer to Appendix B for full details on the toy model experiment.
4 Detection Metrics
Figure 4: Different sample types across the memorization-to-generalization transition for CIFAR10 [63]. The grey histogram shows the distances between synthetic samples and their nearest neighbors from the synthetic set $\mathsf{S}^{\prime}$. The threshold $\delta_{s}$ defines a boundary between the two peaks. The olive histogram depicts the distances from the synthetic samples to their closest neighbor from the training set $\mathsf{S}$, with threshold $\delta_{m}$ separating the two peaks. Memorized samples are located in the left peak of the olive histogram, below $\delta_{m}$. In contrast, generalized and spurious samples appear to the right of $\delta_{m}$ in the olive histogram. Examples of the generated samples forming each of the four peaks of the histograms are shown in the inset frames. For each generated sample, top-4 nearest neighbors from the training set are shown in the top row, and those from the synthetic set are shown in the bottom row. Training set size $K=7310$ was used in this figure (peak of the frequency of spurious states), but the discussed phenomena are general and largely independent of this specific value. The fraction of the memorized, spurious, and generalized samples in the pool of all generated samples is shown in the bottom left panel as a function of the training set size. The inset shows amplified spurious fraction (green curve).
The main goal of our work is to establish the existence of spurious states in diffusion models trained on natural high-dimensional datasets, not just toy model problems. With this goal in mind we design detection metrics that can classify any generated sample into one of the three categories: memorized, spurious, and generalized samples. Fundamentally, these metrics defined below rely on two datasets: $\mathsf{S}$ — the training dataset used to train the diffusion model, and $\mathsf{S}^{\prime}$ — the synthetic dataset generated from the fully trained diffusion model. The size of the synthetic set $\mathsf{S}^{\prime}$ is assumed to be much bigger than the training set size.
The core intuition comes from the typical energy landscape around these three distinct kinds of generated samples, see Figure 1. Memorized samples are elements of the synthetic set $\mathsf{S}^{\prime}$, which have duplicates in the training set $\mathsf{S}$. Spurious samples are elements of the synthetic set $\mathsf{S}^{\prime}$, which have duplicates in the synthetic set $\mathsf{S}^{\prime}$, but do not have duplicates in $\mathsf{S}$. In other words, the same spurious sample appears in the synthetic set more than once. Finally, generalized samples are unique elements of $\mathsf{S}^{\prime}$, which do not have duplicates in either the synthetic set $\mathsf{S}^{\prime}$ nor the training set $\mathsf{S}$.
With this intuition in mind we examine two histograms, see Figure 4. First, for every element of the synthetic set $\mathsf{S}^{\prime}$ the distance to the closest nearest neighbor from the training set $\mathsf{S}$ is shown by the olive histogram. Second, for every element of the synthetic set $\mathsf{S^{\prime}}$ the distance to the closest nearest neighbor in the synthetic set $\mathsf{S}^{\prime}$ is shown by the gray histogram. Both histograms exhibit a clear bimodal shape, indicating that at least two groups of samples are present in the synthetic set. This bimodal structure makes it possible to select the memorized $\delta_{m}$ and spurious $\delta_{s}$ thresholds to separate the samples forming the two peaks in each histogram. The left peak of the olive histogram is composed of synthetic samples that are at a small distance away from the training set $\mathsf{S}$, which can be considered as memorized samples. The left peak of the gray histogram contains synthetic samples that have duplicates in the synthetic set $\mathsf{S}^{\prime}$, indicating they can be memorized if the same sample appears in the training set, or otherwise, spurious. The right peak of the gray histogram contains generalized samples, and the right peak of the olive histogram contains generalized and spurious samples. These observations make it possible to formally define the following detection metrics.
Metric 4.1 (Memorization Detection).
Following [45], we define the memorization detection metric $\mathcal{M}$. Given a target pattern $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}$ and its first nearest neighbor ${\mathbf{x}}_{1}$ extracted from the training set $\mathsf{S}$ according to a distance measure $d(\cdot,\cdot)$, where a small distance corresponds to high similarity, metric $\mathcal{M}$ detects sample $\hat{{\mathbf{x}}}$ as memorized if $d(\hat{{\mathbf{x}}},{\mathbf{x}}_{1})$ is small.
$\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S})=\mathbb{I}\bigg{(}d(\hat{{\mathbf{x}}},{\mathbf{x}}_{1})\leq\delta_{m}\bigg{)}$ (13)
where $\mathbb{I}$ represents the indicator function and $\delta_{m}\in\mathbb{R}$ is a threshold defined by the distance histograms, see Figure 4.
Metric 4.2 (Spurious Detection).
Given a synthetic sample $\hat{{\mathbf{x}}}$ and its first nearest neighbor from the synthetic set ${\mathbf{x}}^{\prime}_{1}\in\mathsf{S}^{\prime}$, the spurious pattern detection metric $\mathcal{S}$ identifies instances where the model generates outputs that do not belong to the training set but have high similarity with samples from the synthetic set (left peak of the gray histogram in Figure 4 after the exclusion of memorized samples).
$\mathcal{S}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime})=\mathbb{I}\bigg{(}d(\hat{{\mathbf{x}}},{\mathbf{x}}^{\prime}_{1})\leq\delta_{s}\bigg{)}\land\neg\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S})$ (14)
where $\delta_{s}\in\mathbb{R}$ is a threshold value chosen based on the gray distance histogram, see Figure 4.
Metric 4.3 (Generalization Detection).
In contrast, the sample $\hat{{\mathbf{x}}}$ is generalized if it is neither memorized nor spurious.
$\mathcal{G}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime})=\neg\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S})\land\neg\mathcal{S}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime})$ (15)
5 Memorization - Spurious - Generalization Transition
Figure 5: Fractions of memorized, spurious, and generalized samples in synthetic sets across training sizes and datasets. As the training data size $K$ increases, memorization decreases and the fraction of generalized samples steadily increases (see top row). The fraction of spurious patterns rises and decreases at the boundary between the memorization and generalization phases (see bottom row).
Using our detection metrics, we computed the fractions of memorized, spurious, and generalized samples for various sizes of training data for different datasets, see Figure 5. These datasets include MNIST [64], FASHION-MNIST [65], CIFAR10 [63], and LSUN-CHURCH [66] scaled down to $64\times 64$ resolution using center-crop and down-scale. For each dataset, we trained DDPM-based diffusion models [23] for $M=38$ different training set sizes. For each model $\alpha=1,...,M$, trained on the training set $\mathsf{S}_{\alpha}$, a synthetic set $\mathsf{S}^{\prime}_{\alpha}$ was generated. To account for duplication, we ensured that each $\mathsf{S}^{\prime}_{\alpha}$ is four times the size of its corresponding $\mathsf{S}_{\alpha}$, and checked that further increase of the synthetic set size does not qualitatively change our results, see Figure 10 in Appendix D. Each sample $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}_{\alpha}$ is classified as either memorized, spurious, or generalized using the above metrics (13), (14), and (15). The fractions of these three pattern types with respect to the size of $\mathsf{S}^{\prime}_{\alpha}$ are plotted in Figure 5. For each dataset, our smallest model was trained on the training set of $|\mathsf{S}_{1}|=2$ data points while the largest model was trained on the entire original training set $\mathsf{S}_{M}=\mathsf{S}$. Please refer to Appendix D for additional details on training, selection of data sizes, and detection metrics’ hyperparameters for each dataset.
The results in Figure 5 clearly demonstrate the transition from memorization to generalization as the dataset size increases. Meanwhile, the collected samples also show distinct characteristics in each of the considered phases, see Figure 2. Specifically, in the small data regime, we see in Figure 5 that the diffusion model predominantly replicates the training data (see the memorized panel of Figure 2, or additional samples in Figure 14 of the Appendix). As the data size surpasses the memorization capacity, e.g., at $K=3931$ for CIFAR10 in Figure 5, we observe a critical transition where the memorization fraction declines and spurious patterns become prominent. Such patterns exhibit strong duplication in the synthetic set (see the spurious panel of Figure 2 and additional samples in Figure 15 of the Appendix). The emergence of such patterns aligns with the onset of generalization, as the model moves away from strictly reproducing the training set and starts generating novel combinations of learned features. Subsequent increase of the training set size leads to the decrease of the fractions of memorized and spurious samples, signaling the transition to the full generalization regime. At this stage, the model completely loses its replication ability and no duplicate samples are detected in either training or synthetic sets (see the generalized panel of Figure 2 and additional samples in Figure 16 of the Appendix). Overall, this particular trend is also observed as the size of the diffusion model’s neural network is varied, see the results in Figure 9 of Appendix D.
6 The Basin of Attraction Across the Transition
Figure 6: Average log-volume of the hyperspheric basin of attraction of different sample types as the training data size $K$ grows, in CIFAR10 [63], computed with up to 512 samples per type. When fewer than 512 samples are available, all points are used. The log-volume of training samples and the transition plot from Figure 5 are included to confirm the shrinking basins of attraction in diffusion models. Shaded regions indicate standard deviation of the log-volume values.
Diffusion models provide us with an opportunity to define the notion of basins of attraction around generated samples to further strengthen their connection to associative memories. Specifically, as the training data size $K$ varies, the sizes of those basins for memorized, spurious, and generalized samples can exhibit interesting behaviors. This section studies these phenomena empirically.
Inspired by [38, 67, 54], we perturb a pattern ${\mathbf{x}}_{0}$ through the forward diffusion process for $T$ steps, and attempt to reverse it back through the reverse diffusion process. The maximum or critical time $t_{c}$, where the reverse process can still recover the pattern ${\mathbf{x}}_{0}$ from its perturbation ${\mathbf{x}}_{t_{c}}$ within an error margin $\delta\in\mathbb{R}$, defines the basin of attraction for pattern ${\mathbf{x}}_{0}$. The radius of this basin $R({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}})=\lVert{\mathbf{x}}_{0}-{\mathbf{x}}_{t_{c}}\rVert$ is the Euclidean norm of the difference between ${\mathbf{x}}_{0}$ and its recoverable perturbation ${\mathbf{x}}_{t_{c}}$. This radius can be converted to volume assuming that the basin of attraction is a sphere of radius $R({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}})$. Please also refer to Appendix G for more details on the experimentation and computation of $t_{c}$ and the volume.
The results for log-volume versus training size are shown in Figure 6. The basins of attraction for training data decrease as $K$ increases. This trend is expected, since packing larger amount of training data in a given volume of the configuration space results in a smaller volume per data point, similarly to the ideal gas in physics. Memorized samples follow a similar trend. These results suggest that as $K$ grows, the diffusion model becomes less effective at denoising the perturbed data points, and instead gets much better at synthesizing. The most significant finding is that each of the three sample types exhibits a distinct log-volume trend: memorized samples have the largest log-volume, generalized samples have the smallest log-volume (this is expected as they should be unlikely recoverable by definition, resulting in close to zero volume), and spurious samples fall in between the two. This finding suggests a clear separation in how different sample types behave as the model’s training data size progresses. Please refer to Figs. (17) and (18) in the Appendix for the results of average log-volume across the transition for the other datasets, which exhibit the same behavior as Figure 6.
To further illustrate the shrinkage of basins of attraction, we visualize the degradation in the denoising process as the training data size $K$ increases, as shown in Figure 7. For each training image, we apply the forward diffusion process using a timestep $t\approx t_{c}$, chosen near the end of the memorization phase (see Figure 19 in the Appendix). The perturbed image is then denoised using the deterministic DDIM sampler [24], starting from that same timestep $t$, across diffusion models trained with varying $K$. When $K$ is small, the diffusion model behaves like a DenseAM, effectively denoising and recovering the original image. However, as $K$ increases, this retrieval ability weakens, and the model shifts toward generating novel samples. In the intermediate regime, before full generalization, the outputs of the denoising process visually resemble spurious patterns: incomplete or noise-like novel generations which indicate the onset of generalization. Please refer to Appendix G for more details on this visual demonstration.
Figure 7: An illustration of the shrinkage of basins of attraction as the training data size increases. Two CIFAR10 [63] and two LSUN-CHURCH [66] training images are independently perturbed using different timesteps $t\approx t_{c}$ (shown on the right), each chosen near the end of the memorization phase (see Figure 19 in the Appendix), and then denoised using the deterministic DDIM sampler [24]. At small training data sizes, the model accurately recovers the original image. As the training data size increases, denoising performance deteriorates: the model shifts from retrieval of data points to generating novel outputs. In intermediate regimes, the outputs resemble spurious patterns: partial and noise-like generations indicative of this transition.
7 The Relative Energy Gap Across the Transition
Figure 8: Average relative energy gap of each sample type and the training samples in CIFAR10 [63] as training data size $K$ grows. The gaps are measured relative to training samples, whose energy is set to zero (energy is defined up to an additive constant). Shaded regions show standard deviation of the gap values.
Lastly, we study how the energy of the three sample types changes as the training data size increases. Following the efficient relative energy computation from [38], we measure the relative energy gap of the three sample types with respect to that of the training data samples. Please refer to Appendix H for full details on the experiment and computation of the relative energy gap for different sample types.
The results are recorded in Figure 8. For small training data sizes, the energy of the memorized samples is low and of the generalized samples is high. As $K$ increases, the energy of the memorized samples increases, while the energy of the generalized samples decreases. The energy of the spurious states is the lowest among these three types of sample at the memorization-generalization transition. Please refer to Figure 22 in the Appendix for the results of the average relative energy gap across the transition for the other datasets, which show a similar behavior as Figure 8.
8 Conclusion
In this work, we established a novel connection between diffusion models and DenseAMs. Specifically, we identified the emergence of creativity (generalization) in diffusion models with the failure of successful memory recalls, via an energy-based theoretical description of diffusion models. Using this theory, we identified a novel phase – spurious states – previously overlooked in the memorization-generalization literature. Finally, we established that diffusion models trained on natural datasets follow the same transition (memorization to spurious phase to generalization) as predicted by our theory. The peak frequency of spurious states among generated samples marks the onset of generalization as the training set size increases. The volume measurements and relative energy gap analysis of different sample types, across multiple datasets consistently support this three distinct phases of computation, providing a comprehensive understanding of how diffusion models evolve from memorization to generalization as the amount of training data increases.
Aknowledgments
M.N. acknowledges the support of PNRR MUR project PE0000013-FAIR.
B.P. acknowledges the support of IBM through the IBM-Rensselaer Future of Computing Research Collaboration (FCRC).
Appendix A Additional Details on the Energy Formulation of the 2D Toy Example
In order to obtain Equation 11, it is easiest to introduce polar coordinates for both the state vector ${\mathbf{x}}$ and the training data ${\mathbf{y}}$:
$\begin{cases}x_{1}=R\cos(\phi)\\ x_{2}=R\sin(\phi)\end{cases}\ \ \ \ \ \ \ \ \ \ \begin{cases}y_{1}=r\cos(\varphi)\\ y_{2}=r\sin(\varphi)\end{cases}$
The integral (11) can then be written as
[…]
and explicitly computed using the definition of the modified Bessel functions [68].
Appendix B Experimental Details of the 2D Toy Example
This section describes the two-dimensional toy example used in Section 3, including details about the neural network architecture, training, generation, and computation of the model’s energy. All of the corresponding code will be made publicly available.
Dataset. To construct the dataset, we uniformly sample points on the unit circle. Concretely, we take ${\mathbf{y}}\sim p({\mathbf{y}})=\frac{1}{2\pi}\delta(r-1)$, where $r=\sqrt{y_{1}^{2}+y_{2}^{2}}$ is the radius and $\delta(r-1)$ is the delta function ensures that all probability mass lies on radius $r=1$. The factor $\frac{1}{2\pi}$ guarantees that the angles are sampled uniformly, relying on the identity $\delta(\sqrt{y_{1}^{2}+y_{2}^{2}}-1)=2\delta(y_{1}^{2}+y_{2}^{2}-1)$ in $p({\mathbf{y}})=\frac{1}{\pi}\delta(y^{2}_{1}+y_{2}^{2}-1)$. In practice, we sample polar coordinates $(r,\varphi)$ with $r=1$ and uniformly sample the angular coordinate $\varphi$ from $[0,2\pi]$. We then convert $(r,\varphi)$ to Cartesian coordinates ($y_{1},y_{2}$) using $y_{1}=r\cos\varphi$ and $y_{2}=r\sin\varphi$.
We first created a data set of $60000$ points using a fixed random seed, then built smaller subsets (e.g., $K\in\{2,4,9,1000,10000\}$) by progressively adding distinct samples without replacement, ensuring that each subset is a strict subset of the next. This approach allows us to systematically examine how the model behaves, as we vary the size of the training set.
Exact Score. Following Equation 12 specified in Section 3, energy can be written in closed form, whose local minima form the continuous data manifold at $R=1$ (throughout this paper we use $r$ for the polar radius of the training data and $R$ for the polar radius of the generated samples). Furthermore, in the case of many data points or $K\rightarrow\infty$, we have $p({\mathbf{y}})=\frac{1}{\pi}\delta(y_{1}^{2}+y_{2}^{2}-1)$. Since $p(R)\propto\exp^{-E^{\text{AM}}(R)}$, the score can be described as
$\begin{split}-\nabla_{R}\log p(R)&=\nabla_{R}E^{\text{AM}}(R)\\ &=2R-\frac{2\beta I_{1}(2\beta R)}{\beta I_{0}(2\beta R)}\\ &=2R-2\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)}\end{split}$ (16)
where $E^{\text{AM}}(R)$ is Equation 12, and $I_{1}$ and $I_{0}$ are the first-order and zero-order modified Bessel functions of the first kind, respectively. With further expansion, the score becomes
$\nabla_{R}\log p(R)=2\left(\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)}-R\right)\qquad\Rightarrow\qquad\nabla_{x}\log p(x)=2\left(\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)}-R\right)\frac{x}{R}$ (17)
We implemented Equation 12 on a $20\times 20$ grid to obtain the exact energy and score of the model, overlaying 1000 samples from the training set. For visualization purposes, we set $\beta=20$, ensuring a well-defined energy landscape. Additionally, we normalized the energy by subtracting its minimum value to ensure that the lowest energy point remains at zero.
VE-SDE. To align with the theoretical framework, we employ a variance exploding (VE) SDE of the form
$\mathrm{d}{\mathbf{x}}_{t}=\sigma\,\mathrm{d}\mathbf{w}_{t}$ (18)
where the diffusion coefficient $g(t)=\sigma$. Assuming $t\in(0,1]$, the variance of the diffusion kernel is given by $\int_{0}^{t}g^{2}(s)\,\mathrm{d}s=\sigma^{2}t$, which allows us to construct the corresponding Gaussian kernel
$p({\mathbf{x}}_{t}|{\mathbf{y}})=\mathcal{N}({\mathbf{x}}_{t};{\mathbf{y}},\sigma^{2}t\mathbf{I})$
However, in practice we used $t\in[\epsilon,1]$ with $\epsilon=10^{-5}$ for numerical stability. The generative dynamics are given by
$\mathrm{d}\mathbf{x}_{t}=\big{[}-\sigma^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})\big{]}\mathrm{d}t+\sigma^{2}\,\mathrm{d}\mathbf{w}_{t}.$ (19)
where this runs backwards in time to match the description of [25]. The deterministic Probability Flow ODE is
$\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=-\frac{1}{2}\sigma^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ (20)
which shares the same marginal distributions as the SDE in Equation 19 and is useful for likelihood estimation.
Model and Training Details. For each training size $K$ in Figure 3, we use a U-Net inspired MLP-based network to estimate the score $s_{\theta}({\mathbf{x}}_{t},t)$, without using any down- or up- sampling blocks. The general architecture includes (1) a Fourier random feature (FRF) timestep embedding layer and a linear layer which projects the concatenation of FRF timestep embedding and input into a latent dimension of 256; (2) an encoder which consists of four non-convolutional residual blocks (using the same latent dimension) with Swish [69] activation in between; (3) a decoder which consists of the same number of residual blocks; and (4) a linear block which projects the latent variable back to the 2D space.
The network was trained in continuous time (with 1000 discretized steps) with $\sigma=1$ using the objective function from [23, 25]. Optimization is performed using the Adam [70] optimizer with a learning rate $lr=10^{-4}$. The batch size is set to $\textbf{min}(K,500)$. All models are trained for $800000$ iterations with the maximum batch size of $500$. Meanwhile, we use Euler-Maruyama discretization method [71] to solve Equation 19 of the reverse SDE for the generation of samples.
Empirical distribution. We represent the data as an empirical distribution
$p(\mathbf{y})=\frac{1}{K}\sum_{\mu=1}^{K}\delta^{(N)}(\mathbf{y}-\bm{\xi}^{\mu})$ (21)
where $N$ is the dimensionality of the data point ${\mathbf{y}}$. Since ${\mathbf{x}}_{t}$ is drawn from the forward process distribution $p({\mathbf{x}}_{t}|{\mathbf{y}})$ that is conditioned on the data point ${\mathbf{y}}$, which can be expressed as a Gaussian kernel $\mathcal{N}({\mathbf{x}}_{t};\mathbf{y},\sigma^{2}t\mathbf{I})$. We can obtain the distribution $p(\mathbf{x}_{t},t)$ as
$\begin{split}p(\mathbf{x}_{t},t)=\int p(\mathbf{x}_{t}\mid\mathbf{y})\,p(\mathbf{y})\,\mathrm{d}\mathbf{y}&=\frac{1}{K}\sum_{\mu=1}^{K}\int\mathcal{N}\!\bigl{(}\mathbf{x}_{t};\mathbf{y},\sigma^{2}t\mathbf{I}\bigr{)}\,\delta^{(N)}\!\bigl{(}\mathbf{y}-\bm{\xi}^{\mu}\bigr{)}\,\mathrm{d}\mathbf{y}\\ &=\frac{1}{K}\sum_{\mu=1}^{K}\mathcal{N}\!\bigl{(}\mathbf{x}_{t};\,\bm{\xi}^{\mu},\,\sigma^{2}t\,\mathbf{I}\bigr{)}\\ &=\frac{1}{K}\sum_{\mu=1}^{K}\frac{1}{(2\pi\,\sigma^{2}t)^{\frac{N}{2}}}\exp\!\Bigl{(}-\,\frac{\|\mathbf{x}_{t}-\bm{\xi}^{\mu}\|_{2}^{2}}{2\,\sigma^{2}t}\Bigr{)}\end{split}$ (22)
which is Equation 4, when the variance of the Gaussian kernel is $\sigma^{2}t$.
Likelihood Computation. Following [25] we can compute the log-likelihood $\log p_{\theta}({\mathbf{x}}_{0})$ given by a diffusion model, with the instantaneous change of variable formula [72], where ${\mathbf{x}}_{0}$ are the generated samples from the model. Replacing Equation 19 into the log-likelihood equation,
$\log p_{0}({\mathbf{x}}_{0};\mathbf{\theta})=\log p_{T}({\mathbf{x}}_{T})+\int_{0}^{T}\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)\mathrm{d}t,\qquad\text{ with }\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)=-\frac{1}{2}\sigma^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ (23)
The function $\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ came from the above Probability Flow ODE (Equation 20), and $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ denotes the divergence of the function (or the trace of its Jacobian).
To estimate the likelihood of the model, we integrate $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ from a small time ${\epsilon}=10^{-5}$ to $T=1$ using a numerical integrator and add the prior logarithmic likelihood to it following Equation 23. The divergence term $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ is computed using the Laplacian, which is computationally feasible for this toy example, instead of using the Hutchinson trace estimator [73] done in [25]. We use the RK45 method [74] implemented in the ODE solver scipy.integrate.solve_ivp from Scipy [75] to solve the above integral.
Energy Computation. To compute the energy, we use Equation 5, which can be also computed using Equation 23. From the Boltzmann distribution $p({\mathbf{x}})=\frac{1}{Z}\exp[-\beta E(\mathbf{x})\bigl{]}$, we obtained $\log p({\mathbf{x}})\propto-\beta E(\mathbf{x})$. In our case the inverse temperature is $\beta=\frac{1}{2\sigma^{2}t}$. The equation for our VE-SDE thus becomes
$E^{\text{DM}}({\mathbf{x}}_{t},t)=-2\sigma^{2}t\log p({\mathbf{x}}_{t},t)$ (24)
To visualize the energy landscape, we observe the energy at $t=0.15$ which corresponds to $\beta=3.3$.
Sample Clustering. To cluster the generated samples, we use the AgglomerativeClustering algorithm from scikit-learn, which is a part of SciPy [75].
Appendix C Additional Details on the Likelihood of Diffusion Models
For concreteness, we further explain the likelihood computation in Equation 23 within this section. Specifically, we want to highlight the formulation done by [25, 72] which resulted in Equation 23.
Probability Flow ODE. Suppose we have the following forward process
$\mathrm{d}{\mathbf{x}}_{t}=\mathbf{f}({\mathbf{x}}_{t},t)\mathrm{d}t+\mathbf{g}({\mathbf{x}}_{t},t)\mathrm{d}\mathbf{w}_{t}$ (25)
where $\mathbf{f}(\cdot,t):\mathbb{R}^{N}\rightarrow\mathbb{R}^{N}$, $\mathbf{g}(\cdot,t):\mathbb{R}^{N}\rightarrow\mathbb{R}^{N\times N}$ and $N$ denotes the dimensionality of ${\mathbf{x}}_{t}$. Using the derivations done by [25] in their Appendix D, the evolution of the marginal probability density $p_{t}({\mathbf{x}}_{t})$ is
$\frac{\partial p_{t}({\mathbf{x}}_{t})}{\partial t}=-\sum^{N}_{i=1}\frac{\partial}{\partial x_{i}}\big{[}f_{i}({\mathbf{x}}_{t},t)p_{t}({\mathbf{x}}_{t})\big{]}+\frac{1}{2}\sum^{N}_{i=1}\sum^{N}_{j=1}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}\bigg{[}\sum^{N}_{k=1}g_{ij}({\mathbf{x}}_{t},t)g_{jk}({\mathbf{x}}_{t},t)p_{t}({\mathbf{x}}_{t})\bigg{]}$ (26)
which corresponds to Fokker-Planck equation [76]:
$\begin{split}\frac{\partial p_{t}({\mathbf{x}}_{t})}{\partial t}&=-\sum^{N}_{i=1}\frac{\partial}{\partial x_{i}}\big{[}f_{i}({\mathbf{x}}_{t},t)p_{t}({\mathbf{x}}_{t})\big{]}+\frac{1}{2}\sum^{N}_{i=1}\frac{\partial}{\partial x_{i}}\bigg{[}\sum^{N}_{j=1}\frac{\partial}{\partial x_{j}}\bigg{[}\sum^{N}_{k=1}g_{ij}({\mathbf{x}}_{t},t)g_{jk}({\mathbf{x}}_{t},t)p_{t}({\mathbf{x}}_{t})\bigg{]}\bigg{]}\\ &=-\sum^{N}_{i=1}\frac{\partial}{\partial x_{i}}\big{[}f_{i}({\mathbf{x}}_{t},t)p_{t}({\mathbf{x}}_{t})\big{]}\\ &\quad\,\,+\frac{1}{2}\sum^{N}_{i=1}\frac{\partial}{\partial x_{i}}\bigg{[}p_{t}({\mathbf{x}}_{t})\,\nabla\cdot\big{[}\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big{]}+p_{t}({\mathbf{x}}_{t})\big{[}\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big{]}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg{]}\\ &=-\sum^{N}_{i=1}\frac{\partial}{\partial x_{i}}\bigg{\{}f_{i}({\mathbf{x}}_{t},t)p_{t}({\mathbf{x}}_{t})\\ &\quad\,\,-\frac{1}{2}\bigg{[}\nabla\cdot\big{[}\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big{]}+\big{[}\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big{]}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg{]}p_{t}({\mathbf{x}}_{t})\bigg{\}}\\ &=-\sum^{N}_{i=1}\frac{\partial}{\partial x_{i}}\big{[}\tilde{f}_{i}({\mathbf{x}}_{t},t)\big{]}p_{t}({\mathbf{x}}_{t})\end{split}$ (27)
where $\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)={\mathbf{f}}({\mathbf{x}}_{t},t)-\frac{1}{2}\nabla\cdot\big{[}\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big{]}-\frac{1}{2}\big{[}\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big{]}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$. With careful inspection of Equation 27, it is equal to the Liouville equation if the diffusion term $\tilde{\mathbf{g}}({\mathbf{x}},t)=0$ and essentially, it is the probability flow ODE where
$\begin{split}\mathrm{d}{\mathbf{x}}_{t}&=\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)\mathrm{d}t+\tilde{\mathbf{g}}({\mathbf{x}}_{t},t)\mathrm{d}{\mathbf{w}}_{t}\\ &=\bigg{\{}{\mathbf{f}}({\mathbf{x}}_{t},t)-\frac{1}{2}\nabla\cdot\big{[}\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big{]}-\frac{1}{2}\big{[}\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big{]}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg{\}}\mathrm{d}t\end{split}$ (28)
Using Equation 28, we can derive an appropriate probability flow ODE from the forward process (1) introduced in the main text. For example, we have the following equation
$\mathrm{d}{\mathbf{x}}_{t}=\underset{:=\,\tilde{\mathbf{f}}({\mathbf{x}}_{t},t)}{\underbrace{\bigg{\{}\mathbf{f}({\mathbf{x}}_{t},t)-\frac{1}{2}g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg{\}}}}\mathrm{d}t$ (29)
for the toy model where $\mathbf{f}({\mathbf{x}}_{t},t)=0$.
Log-likelihood. Furthermore, if we first take the logarithm of Equation 27 where
$\begin{split}\frac{\partial\log p_{t}({\mathbf{x}}_{t})}{\partial t}=\frac{1}{p_{t}({\mathbf{x}}_{t})}\frac{\partial p_{t}({\mathbf{x}}_{t})}{\partial t}=-\sum^{N}_{i=1}\frac{\partial\tilde{f}_{i}({\mathbf{x}}_{t},t)}{\partial x_{i}}=\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)\end{split}$ (30)
we can compute the log-likelihood of $p_{0}({\mathbf{x}}_{0})$ using the following equation
$\log p_{0}({\mathbf{x}}_{0})=\log p_{T}({\mathbf{x}}_{T})+\int_{0}^{T}\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)\mathrm{d}t$ (31)
where $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ is parameterized as $\nabla\cdot\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)$ since $s_{\theta}({\mathbf{x}}_{t},t)=\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t};\theta)$. Moreover, in general the term $\nabla\cdot\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)$ is computed via the Hutchinson trace estimator [73],
$\nabla\cdot\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)=\mathbb{E}_{p(\epsilon)}\big{[}\epsilon^{\top}\nabla_{{\mathbf{x}}}\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)\epsilon\big{]}$ (32)
where $\nabla_{{\mathbf{x}}}\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)$ denotes the Jacobian of $\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)$ and the random variable $\epsilon$ satisfies $\mathbb{E}_{p(\epsilon)}[\epsilon]=0$ and $\text{Cov}_{p(\epsilon)}[\epsilon]=\mathbf{I}$.
However, for the toy model in Section 3, due to its low dimensionality we find that it is possible to not utilize the Hutchinson trace estimator at all, and instead opt to compute the trace of the Jacobian without using the noise estimators — to get the actual trace of the Hessian of the energy.
Appendix D Experimental Details on the Transition Mapping
Dataset Point A Point B Total
(Data Size) (Data Size) (Data Size)
MNIST 4,000 32,000 60,000
CIFAR10 2,000 16,000 50,000
LSUN-CHURCH 2,000 16,000 126,227
FASHION-MNIST 4,000 16,000 60,000
Table 1: Table showing the critical points $A$ and $B$ of the memorization-generalization transition for each dataset. We use 30 linearly spaced points between $A$ and $B$ to finely characterize the transition plots shown in Figure 5.
Points Selection. For the computation of Figure 5, we follow the experiment setup of [45] and conduct a sparse search starting at data size $500$ and doubling it all the way to the total data size $|\mathsf{S}|$, i.e., $K=500,1000,2000,\dots,|\mathsf{S}|$. We then identified two transitional critical points, $A$ and $B$, see Table 1, to conduct a more thorough search to elucidate the memorization-generalization transition. Point $A$ indicates the initial drop in memorization while $B$ signals the plateauing of memorization. To capture the finer details of this transition, we performed a search of $30$ inclusive linearly spaced points, from point $A$ to point $B$. For regions outside of the transition, using linear spacing, we sample $5$ points from $|\mathsf{S}|=2$ to point $A$, and another $5$ points from $B$ to the total dataset size $|\mathsf{S}|$, inclusively. In other words, we train a separate model for each selected data size and generate the corresponding evaluation and synthetic sets — keeping the same configuration for all models, see Table 3. Then, we computed the memorization, spurious, and generalization fractions using equations (13), (14), and (15). The selection process for the values $\delta_{m}$ and $\delta_{s}$ is explained below.
Dataset $\delta_{m}$ $\delta_{s}$ $d(\cdot)$
CIFAR10 0.03 0.018 LPIPS
LSUN-CHURCH 0.12 0.05 LPIPS
MNIST 4.5 3.5 $L_{2}-$distance
FASHION-MNIST 4.5 2.5 $L_{2}-$distance
Table 2: Table displaying memorized and spurious threshold values in the computation of the transition plots in Figure 5.
Selection of the Distance Metric. For high-dimensional datasets, CIFAR10 [63] and LSUN-CHURCH [66], we utilized LPIPS [77] with the AlexNet [78] backbone, as the function $d(\cdot,\cdot)$ for both memorization (13) and spurious (14) detection metrics. We selected this approach since it is a commonly used perceptual metric that compares the similarity between two images based on their feature representations. Moreover, it has been shown to better align with human judgment of visual similarity, making it ideal for assessing the quality and diversity of generated samples in these high-dimensional image datasets [77]. For simpler datasets like MNIST [64] and FASHION-MNIST [65], where images are single-channel and less complex, we found that $L_{2}$-distance suffices for both memorization and spurious detection metrics.
Selection of Threshold Values $\delta_{m}$ and $\delta_{s}$. The detection thresholds, $\delta_{m}$ and $\delta_{s}$, were set based on the chosen distance metric $d(\cdot,\cdot)$ and visual inspection of the distance histograms for bimodality as mentioned in Section 4 of the main text. Moreover, these thresholds were chosen to reflect the varying visual complexity and feature richness of the datasets, and their values for each dataset are shown in Table 2. For CIFAR10, we utilized the histograms obtained at $K=7310$, which demonstrate strong bimodality shape, to tune our threshold values to minimize the number of spurious samples classified as memorized samples and vice versa, via manual inspection. For manual inspection, a general rule we followed — it is better to have more false positives in the generalized set than the memorized set since spurious samples can be seen as early-generalized samples. Thus, we observe the least-$5$ spurious samples and ensure that they do not resemble memorized samples. We performed the same process for memorized samples. The entire process is then repeated for the other three datasets. We utilized $K=4896$ for LSUN-CHURCH, $K=21379$ for MNIST, and $K=7724$ for FASHION-MNIST to select $\delta_{m}$ and $\delta_{s}$. For more details on the selected histograms, please refer to Figs. (4), (13), (11), and (12).
Dataset Initial Latent Channel Multipliers Number of Parameters Batch Size Training Iterations
MNIST 128 (1, 2, 2) 24.5M 128 400,000
CIFAR10 128 (1, 2, 2, 2) 35.7M 128 500,000
CIFAR10 96 (1, 2, 2, 2) 20.1M 128 500,000
CIFAR10 64 (1, 2, 2, 2) 8.9M 128 500,000
LSUN-CHURCH 96 (1, 1, 2, 2, 4, 4) 61.7M 64 800,000
FASHION-MNIST 128 (1, 2, 2) 24.5M 128 400,000
Table 3: Table displaying both model and training configurations for each dataset.
Model and Training Details. For each point in our transition plots in Figure 5, we train a DDPM-based diffusion model, where the score model is a PixelCNN++ based U-Net [79, 80]. We keep the variances, $\beta_{\text{min}}=10^{-4}$ and $\beta_{\text{max}}$ = $2\times 10^{-2}$, timesteps $T=1000$, and learning rate $lr=$ $2\times 10^{-4}$ for all models and datasets. Each model has $2$ residual blocks [81] for each down- and up- sampling layer, while an attention block is placed at 16×16$\times$16 × resolution. We only modified the channel multipliers for each model based on the complexity of the dataset, see Table 3. If the dataset size is smaller than the specified batch size, we take the batch size to be equal to that small dataset size. For generation or inference, we use the exponential moving average (EMA) of each trained model, as delineated in [23], which was obtained with the decay value set as $0.9999$ during training. Please note that we did not use random flipping in the training of our models, since we want our measurements to reflect the training data size at best as random flipping implicitly increases the number of patterns that the models see during training. However, we did use dropout (of value $0.1$) for the training of CIFAR10, MNIST, and FASHION-MNIST models. Lastly, for each of the training set $\mathsf{S}_{\alpha}$, where $\alpha=1,\dots,M$, they are split from the original dataset given a specific size, using the same random seed value of $3407$.
Figure 9: The fractions of memorized, spurious, and generalized samples in synthetic sets across different training sizes for CIFAR10 and various model’s widths. As the training data size $K$ increases, memorization decreases while the emergence of spurious patterns follows and the fraction of generalized patterns steadily increase (see top row). The fraction of spurious patterns rises and decreases at the boundary between the memorization and generalization phases (see bottom row). Additionally, with a smaller number of parameters, the memorization-generalization transition happens at an earlier stage (or smaller $K$), since the model has a memorization capacity. In contrast, as the model’s size increases, the transition happens at an later stage (or larger $K$), where it is able to memorize more training data. The numbering, e.g., $64$, $96$, and $128$, denotes the initial latent dimension of the U-Net, see Table 3.
Additional Experiments. Inspired by [45, 46], we extend our investigation on the memorization-generalization transition by studying the effects of varying the parameter size of the diffusion model for CIFAR10. The results of this experiment are recorded in Figure 9. Specifically, we trained two additional sets of 38 diffusion models, with the initial latent dimensions of 64 and 96 respectively, and repeat the same experimental setup delineated above, see also Table 3. For U-Net (64), we identified the critical points, $A$ and $B$, to be $500$ and $8000$ respectively. Meanwhile, for U-Net (96), these critical points are $2000$ and $16000$. To select $\delta_{m}$ and $\delta_{s}$, we followed the same process detailed above by observing the distance histograms to see where bimodality occurs with respect to the distance value, and manually inspecting the images to ensure there is a minimal amount of false positives in memorized and spurious samples. For U-Net (64), the threshold values $\delta_{m}$ and $\delta_{s}$ are both selected to be $0.02$ — while for U-Net (96), the threshold values $\delta_{m}$ and $\delta_{s}$ are chosen as $0.03$ and $0.02$, respectively.
With the smaller U-Net (64) model, we observe the memorization-generalization transition happens at an earlier stage with respect to the larger U-Net (128) model, where the peak of spurious samples appears at $K=2568$ instead of $K=5862$. Similarly, for U-Net (96), the peak of spurious samples happens at an early stage of $K=3931$ instead. Overall, these results indicate that as the parameter size of the diffusion model increases, the memorization capacity of the model is increased and thus, generalization is delayed since the model can memorize more of its training data denoted by the appearance of spurious patterns happening at a larger training data size $K$. In contrast, when the memorization capacity is reduced, generalization is initiated at an earlier stage (or smaller training data size) with respect to when the parameter size of the model is larger. Most importantly, the emergence of spurious patterns remains consistent despite the variations in the parameter size of the diffusion model.
Figure 10: Memorization-Generalization transition plots for the CIFAR10 dataset. The left figure depicts the memorization-generalization transition computed with 4×4$\times$4 × synthetic sets for various training data sizes. Meanwhile, right figure shows the transition computed using 8×8$\times$8 × synthetic sets for the same training data sizes. With the larger synthetic sets, more spurious patterns are detected at the peak ($K=5862$). However, the overall trend of these patterns quickly rise, during the co-existence of memorization and generalization, and fall as generalization becomes the dominant phase, is maintained. Both figures use the same memorized and spurious threshold values detailed in Table 2.
Additionally, to verify the consistency of the memorization-generalization transition, we explored the influence of larger synthetic sets and recorded our results in Figure 10. Specifically, we repeated our experiment done in Figure 5 for CIFAR10 using 8×8$\times$8 × synthetic sets — which was originally done using 4×4$\times$4 × synthetic sets. Based on Figure 10, we observe a slight increase in the fraction of spurious samples when using the larger 8×8$\times$8 × synthetic sets. Specifically, at $K=5862$, the fraction increases from $0.04$ to roughly $0.06$ with respect to the synthetic set size. Nonetheless, the trend of memorization-generalization, with the emergence of spurious samples, is still maintained regardless of the multiplier for the synthetic size — highlighting the consistency of the spurious phenomenon in diffusion models in relation to the training data size.
Hardware Details. Training of the models detailed in Table 3 were done using many NVIDIA Tesla V100 GPUs (a single GPU was used to train each model). Each GPU has 32GB of memory and is linked with IBM Power9 processors which clocked at 3.15 GHz. Each CIFAR10 model took roughly 18, 20, and 22 hours to train for different initial latent dimensions, e.g., 64, 96, 128, respectively. Meanwhile, each LSUN-CHURCH model requires roughly 56 hours of training. In contrast, FASHION-MNIST and MNIST models require approximately 17 hours to be fully trained (per model). Please note that these durations are with respect to the required training iterations for each model and their corresponding dataset detailed in Table 3. Please note that we utilized the same hardware for our experiments as well.
Appendix E Additional Visualizations of Different Sample Types Across the Transition
E.1 LSUN-CHURCH
Figure 11: Different sample types across the memorization-to-generalization transition for LSUN-CHURCH [66], defined by the spurious and memorized thresholds, $\delta_{s}$ and $\delta_{m}$. The grey histogram shows the distances between synthetic samples and their nearest neighbors from the synthetic set $\mathsf{S}^{\prime}$. The threshold $\delta_{s}$ is defined as a boundary between the two peaks. The olive histogram depicts the distances from the synthetic samples to their closest neighbor from the training set $\mathsf{S}$, with threshold $\delta_{m}$ separating the two peaks. The threshold $\delta_{m}$ is chosen much stricter here such that it works well for all training dataset sizes via visual inspection. Memorized samples are located in the left peak of the olive histogram, below $\delta_{m}$. In contrast, generalized and spurious samples appear to the right of $\delta_{m}$ (olive histogram). Examples of the generated samples forming each of the four peaks of the histograms are shown in the inset frames. For each generated sample top-4 nearest neighbor images from the training set are shown in the top row, and top-4 nearest neighbors from the synthetic set are shown in the bottom row. Training set size $K=4896$ was used in this figure (at the peak of the frequency of spurious states), but phenomena discussed are generic and largely independent of this specific value. The fraction of the memorized, spurious, and generalized samples in the pool of all generated samples is shown in the bottom left panel as a function of the training set size. The inset shows amplified spurious fraction (green curve).
E.2 FASHION-MNIST
Figure 12: Different sample types across the memorization-to-generalization transition for FASHION-MNIST [65], defined by the spurious and memorized thresholds, $\delta_{s}$ and $\delta_{m}$. The grey histogram shows the distances between synthetic samples and their nearest neighbors from the synthetic set $\mathsf{S}^{\prime}$. The threshold $\delta_{s}$ is defined as a boundary between the two peaks. The olive histogram depicts the distances from the synthetic samples to their closest neighbor from the training set $\mathsf{S}$, with threshold $\delta_{m}$ separating the two peaks. Memorized samples are located in the left peak of the olive histogram, below $\delta_{m}$. In contrast, generalized and spurious samples appear to the right of $\delta_{m}$ (olive histogram). Examples of the generated samples forming each of the four peaks of the histograms are shown in the inset frames. For each generated sample top-4 nearest neighbor images from the training set are shown in the top row, and top-4 nearest neighbors from the synthetic set are shown in the bottom row. Training set size $K=7724$ was used in this figure (at the peak of the frequency of spurious states), but phenomena discussed are generic and largely independent of this specific value. The fraction of the memorized, spurious, and generalized samples in the pool of all generated samples is shown in the bottom left panel as a function of the training set size. The inset shows amplified spurious fraction (green curve).
E.3 MNIST
Figure 13: Different sample types across the memorization-to-generalization transition for MNIST [64], defined by the spurious and memorized thresholds, $\delta_{s}$ and $\delta_{m}$. The grey histogram shows the distances between synthetic samples and their nearest neighbors from the synthetic set $\mathsf{S}^{\prime}$. The threshold $\delta_{s}$ is defined as a boundary between the two peaks. The olive histogram depicts the distances from the synthetic samples to their closest neighbor from the training set $\mathsf{S}$, with threshold $\delta_{m}$ separating the two peaks. Memorized samples are located in the left peak of the olive histogram, below $\delta_{m}$. In contrast, generalized and spurious samples appear to the right of $\delta_{m}$ (olive histogram). Examples of the generated samples forming each of the four peaks of the histograms are shown in the inset frames. For each generated sample top-4 nearest neighbor images from the training set are shown in the top row, and top-4 nearest neighbors from the synthetic set are shown in the bottom row. Training set size $K=21379$ was used in this figure, which is slightly past the peak of the frequency of spurious states. The phenomena discussed are generic and largely independent of this specific value. The fraction of the memorized, spurious, and generalized samples in the pool of all generated samples is shown in the bottom left panel as a function of the training set size. The inset shows amplified spurious fraction (green curve).
Appendix F Additional Visualizations of the Three Sample Types
F.1 Memorized Samples
Figure 14: Visualizations of additional memorized patterns and their top-4 nearest neighbors for different datasets. The top row illustrates nearest neighbors from the training set while the bottom row depicts those from the synthetic set. Memorized samples are duplicates of the training set. During the strong memorization phase, duplicates are also found within the synthetic set. Note, even though our memorized detection metric (Eq. 13) does not utilize the synthetic set, we are showing the nearest neighbors obtained from it, for consistency.
F.2 Spurious Samples
Figure 15: Visualizations of additional spurious patterns and their top-4 nearest neighbors for different datasets. The top row depicts nearest neighbors from the training set, while the bottom row shows those from the synthetic set. Spurious patterns are demonstrated to arise from the onset of generalization where the mixing of training data points begins. Since the model’s generalization is at its infancy, duplicates of spurious patterns sometimes appear several times in the synthetic set $S^{\prime}$, much like memorized patterns. Additionally, these samples lack the uniqueness to be considered as generalized samples as the model has yet to fully learn the underlying data distribution.
F.3 Generalized Samples
Figure 16: Visualization of generalized patterns and their top-4 nearest neighbors for different datasets. The top row illustrates nearest neighbors from the training set while the bottom row depicts those from the synthetic set. Generalized samples are novel samples, which have little to no resemblance to their nearest neighbors in training and synthetic sets.
Appendix G Experimental Details on the Basin of Attraction Analysis Across the Transition
Algorithm Details. To obtain the results recorded in the above Figs. (17) and (18), we first computed the critical time $t_{c}$ given an image ${\mathbf{x}}_{0}={\mathbf{y}}$. Specifically, this is done by using Algorithm (1), where $t_{c}$ is obtained using the deterministic DDIM [24] sampler to denoise the perturbed version ${\mathbf{x}}_{t}$ of ${\mathbf{x}}_{0}$ for each step in $t=1,\dots,T$, given $N^{\prime}$ number of denoising steps starting backward from $t$ to 00. To save computational time, the algorithm utilizes a stride constant $s$ to skip over some of the timesteps $t=1,1+s,1+2s,\dots,T-s-1$. For each timestep $t$, the variance preserving forward process is applied onto the targeted image ${\mathbf{x}}_{0}$ yielding ${\mathbf{x}}_{t}$. To account for the variations in noise, the image ${\mathbf{x}}_{0}$ is perturbed for $M$ number of times at each step $t$.
To check whether if the critical time $t_{c}$ is reached, the algorithm utilizes a distance metric $d$ — which is LPIPS [77] equipped with the vgg-16 backbone [82] for all of the datasets — checking if a recovered image $\hat{{\mathbf{x}}}_{0}$ obtained from the DDIM sampler [24] with $N^{\prime}$ denoising steps is similar to the original pattern ${\mathbf{x}}_{0}$ within an error threshold $\delta_{d}\in(0,1)$. Unlike the experiment on the memorization-generalization transition detailed in Section 5, we can afford to trade computational time for better accuracy in this experiment by using a larger backbone for LPIPS. Please note that the metric LPIPS is bounded between 0 (most similar) and 1 (least similar). To keep track of the critical time $t_{c}$ for all perturbations, we allow it to be a vector of integers $t_{c}\in\mathbb{N}^{M}$, where $M$ is the number of perturbations or trials for the image ${\mathbf{x}}_{0}$. To ensure that $t_{c}$ is correct, we have a binary stopping vector $\hat{S}\in\{0,1\}^{M}$, which keeps track of whether each perturbation of ${\mathbf{x}}_{0}$ is no longer recoverable, such that we only update an entry of the vector $t_{c}$ if and only if the corresponding entry in $\hat{S}$ is a zero (not stopped).
Since the entire process of finding $t_{c}$ is a rough search over a set of timesteps, we require a “breaking condition” for the search. In Algorithm 1, this condition is determined by the value $p\in[0,1]$, which is a probability corresponding to the recovery probability of the image ${\mathbf{x}}_{0}$ with respect to its $M$ perturbations. To compute $p$, we calculate the distance vector $\mathbf{d}=d(\hat{{\mathbf{x}}}_{0},{\mathbf{x}}_{0})$ where $\hat{{\mathbf{x}}}_{0}$ is the predicted version of ${\mathbf{x}}_{0}$ obtained from denoising ${\mathbf{x}}_{t}$. Then, we compute the recovery binary vector $\mathbf{r}\in[0,1]^{M}$ by applying the indicator function onto $\mathbf{d}$ using the threshold $\delta_{d}$. Lastly, we calculate the probability value $p=\frac{1}{M}\sum^{M}_{i=1}\mathbf{r}_{i}$. If this value $p$ is below the probability threshold $\delta_{p}$, we stop the search for critical time across all trials.
Experimentation Details. For each training data size $K$, we used a maximum of the top 512 samples belonging to each of the sample types — e.g., training data, memorized, spurious, and generalized samples — to find their critical time $t_{c}$ using Algorithm 1. Note, with the exception of the training data points, we sorted samples belonging to each set — identified via the detection metrics (13), (14) and (15) — by their distance from least to greatest. We computed the log volume using the following equation
$V({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}}=\frac{\pi^{\frac{N}{2}}}{\Gamma(\frac{N}{2}+1)}R({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}})^{N}$ (33)
where $\Gamma(\cdot)$ is the Euler’s gamma function [83] and $N$ is the dimensionality of the image. The radius $R({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}})$ is the Euclidean norm of the difference between the original image ${\mathbf{x}}_{0}$ and its recoverable perturbation ${\mathbf{x}}_{t_{c}}$. For Figs. (6), (17), and (18), we plotted the average log-volume value per each training data size $K$. Meanwhile, for Algorithm 1, we fixed the stopping probability $\delta_{p}$ as $0.8$, the number of perturbations $M$ as $20$, the DDIM steps $N^{\prime}$ as $10$ and the stride constant $s$ as $10$ for all datasets. At some training data size $K$, in which there are very few samples in either memorized, spurious or generalized set, we opted to not include the statistics of such a set, partly due to its large variance. We remove the statistics of such a set if the number of its samples is less than $0.1\%$ of its corresponding synthetic set for a training data size $K$.
For CIFAR10 and LSUN-CHURCH, we utilized the threshold $\delta_{d}$ as $0.03$ and $0.12$ respectively. We kept these threshold values similar to the memorized threshold values $\delta_{m}$ detailed in Table 2, since they have proven to be effective at separating the memorized samples from non-memorized ones in the previous experiment. However, since it is difficult to select the threshold value for the $L_{2}$-distance metric — because the range of the metric is between 00 and $\infty$ — we opted to use LPIPS to experiment with MNIST and FASHION-MNIST using the same threshold $\delta_{d}=0.1$ for both. To find this optimal threshold value, we selected three data sizes, where each belongs to one of the three phases: memorization, spurious, or generalization. We then observed which of the several threshold values, starting from $0.01$ to $0.15$ with the increment of $0.01$, reflects best of the performance given by LPIPS using vgg-16 [82] as its backbone. Specifically, we evaluated these thresholds using $K\in\{4000,11724,60000\}$ for MNIST, and $K\in\{4000,8965,60000\}$ for FASHION-MNIST. We simply perform Algorithm 1 for those three data sizes on a small batch of $32$ samples for the three sample types, and manually check whether the recovered samples (using the obtained critical time $t_{c}$ to denoise from) are similar to the original images and if this observation is consistent across those three data sizes.
G.1 Algorithm for Computing the Critical Time
1
2
3Inputs
4        Pattern ${\mathbf{y}}$
5        Score model $s_{\theta}$
6        DDIM sampler $\Phi$
7        Error threshold $\delta_{d}$
8        Distance metric $d$
9        Denoising steps $N^{\prime}$
10        Total diffusion time $T$
11        Time stride $s$
12        Number of trials $M$
13        Probability threshold $\delta_{p}$
14        Cumulative variances $\bar{\alpha}$
15        Variances $\beta$
16Compute $t_{c}$
        ${\mathbf{x}}_{0}\leftarrow\text{duplicate}({\mathbf{y}},M)$
         // Duplicate pattern ${\mathbf{y}}$ for $M$ times
        $t_{c}\leftarrow\mathbf{0}$
         // Initialize critical time tracker
        $\hat{S}\leftarrow\mathbf{0}$
         // Track trials that have stopped
17        for $t=1,1+s,1+2s,\dots,T-s-1$ do
               ${\mathbf{z}}_{t}\sim\mathcal{N}(0,\mathbf{I})$
                // Sample noise
18
              ${\mathbf{x}}_{t}\leftarrow\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\mathbf{z}}_{t}$
                // Forward diffusion
19
              $\hat{{\mathbf{x}}}_{0}\leftarrow\Phi(s_{\theta},{\mathbf{x}}_{t},t,\bar{\alpha},\beta,N^{\prime})$
                // DDIM denoising
20
              $\mathbf{d}\leftarrow d(\hat{{\mathbf{x}}}_{0},{\mathbf{x}}_{0})$
                // Compute distance per trial
21
              $\mathbf{r}_{i}\leftarrow\begin{cases}1&\text{if }\mathbf{d}_{i}\leq\delta_{d}\\ 0&\text{otherwise}\end{cases}$ for $i=1,\dots,M$
                // Binary recovery vector
22
              $p\leftarrow\frac{1}{M}\sum_{i=1}^{M}\mathbf{r}_{i}$
                // Recovery probability
23
24              if $p<\delta_{p}$ then
25                      return $t_{c}$
26              else
27                      $\hat{S}\leftarrow\text{update-stop}(\hat{S},\mathbf{r})$
28                      $t_{c}\leftarrow\text{update-time}(t,t_{c},\hat{S})$
29
30
31Output
32        Critical time $t_{c}\in\mathbb{N}^{M}$
Algorithm 1 Compute Critical Time $t_{c}$ for a single pattern ${\mathbf{y}}$
G.2 The Basin of Attraction Across the Transition for All Datasets
Figure 17: Average log-volume of the hyperspheric basin of attraction of different sample types as the training data size $K$ grows, in CIFAR10 and LSUN-CHURCH computed using Equation 33 with up to 512 samples per type. When fewer than 512 samples are available, all points are used. The log-volume of training samples and the transition plot from Figure 5 are included to confirm the shrinking basins of attraction in diffusion models. Variations in the critical time $t_{c}$ during the transition, which affect the radius $R$, lead to distinct log-volume trends across sample types, see the recorded critical times in Figure 19. Shaded regions indicate standard deviation of the log-volume.
Figure 18: Average log-volume of the hyperspheric basin of attraction of different sample types as the training data size $K$ grows, in FASHION-MNIST and MNIST, computed using Equation 33 with up to 512 samples per type. When fewer than 512 samples are available, all points are used. The log-volume of training samples and the transition plot from Figure 5 are included to confirm the shrinking basins of attraction in diffusion models. Variations in critical time $t_{c}$ and radius $R$ lead to distinct log-volume trends across sample types, see the recorded critical times in Figure 20. Shaded regions indicate standard deviation of the log-volume.
G.3 Average Critical Time Across The Transition
Figure 19: Average critical time, obtained using Algorithm 1, with a maximum of 512 samples across the memorization-generalization transition for each sample type in CIFAR10 and LSUN-CHURCH. The average critical times for the training data samples at each training data size $K$ are included to illustrate the overall decrease in the time and also the amount of perturbation in which the diffusion model can effectively recover the original sample within some error. Each sample type exhibits a distinctive critical time at each training data size $K$. The shaded region (or the error bars) shown here is the standard deviation computed from all of the critical times at each $K$ with respect to each type of samples. The transition plot of each dataset is shown here to visually guide where the memorization-generalization transition occurs.
Figure 20: Average critical time, obtained using Algorithm 1, with a maximum of 512 samples across the memorization-generalization transition for each sample type in FASHION-MNIST and MNIST. The average critical times for the training data samples at each training data size $K$ are included to illustrate the overall decrease in the time and also the amount of perturbation in which the diffusion model can effectively recover the original sample within some error. Each sample type exhibits a distinctive critical time at each training data size $K$. The shaded region (or the error bars) shown here is the standard deviation computed from all of the critical times at each $K$ with respect to each type of samples. The transition plot of each dataset is shown here to visually guide where the memorization-generalization transition occurs.
G.4 Visual Demonstration of the Shrinkage of the Basins of Attraction
Experiment Details. For Figure 7, we randomly selected a subset of training images that are shared across all 38 diffusion models, each trained with a different training set size. To determine an appropriate timestep $t\approx t_{c}$, we referred to Figure 19, which plots the average critical time $t_{c}$ as a function of training data size $K$. From this plot, we identified a plateau point in the memorization fraction curve, occurring around $K=3931$, and selected $t_{c}\approx 450$ for CIFAR10. Using the timestep starting at $450$, each image was perturbed via the forward process and then denoised using the DDIM sampler [24] with 25 steps across all models (starting backward from the indicated timestep). We manually examined the denoised outputs across training sizes to assess whether the resulting generations effectively illustrate the shrinkage of the basin of attraction. If the output appeared overly memorized (or too close to the original), we increased the timestep by increments of 10 and repeated the process to obtain a more illustrative transition. Lastly, in the case of LSUN-CHURCH, we selected the initial timestep to be $550$ accordingly to Figure 19.
Appendix H Additional Details on the Relative Energy Computation and its Experimentation
In this section, we provide details on the formulation of the relative energy done by [38] and an overview on the experiment done in Section 7 of the main text.
Relative Energy Formulation. Consider the SDE for the reverse process in Equation 2 of the main text
$\mathrm{d}{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})]\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (34)
We can re-express the above equation with the potential function $u(x,t)$
$-\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)=\mathbf{f}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$ (35)
by using the drift term. By moving the minus sign and taking the integration of Eq. (35) with respect to ${\mathbf{x}}_{t}$, we have
[…]
which is the potential function $u({\mathbf{x}}_{t},t)$ of the generative process. Substituting this potential function (36) into the reverse SDE (2), we have the following generative process
$\mathrm{d}{\mathbf{x}}_{t}=-\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (37)
For our experimental setting, where we use the DDPM [23] setting, we have the following variance preserving forward SDE
$\mathrm{d}{\mathbf{x}}_{t}=-\frac{1}{2}\beta(t){\mathbf{x}}_{t}\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}{\mathbf{w}}_{t}$ (38)
and the corresponding reverse (or generative) process SDE
$\mathrm{d}{\mathbf{x}}_{t}=\bigg{[}-\frac{1}{2}\beta(t){\mathbf{x}}_{t}-\beta(t)\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg{]}\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (39)
With Equation 37, we have
$\mathrm{d}{\mathbf{x}}_{t}=-\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (40)
and by using Equation 35, the potential for our experimental setting equates to
[…]
Assume we have two images, ${\mathbf{x}}_{1}(t)$ and ${\mathbf{x}}_{2}(t)$, where ${\mathbf{x}}_{1}$ is a target image and ${\mathbf{x}}_{2}(t)$ is a reference image at time $t$. Since we are interested in the energy of unperturbed images, let ${\mathbf{x}}_{1}={\mathbf{x}}_{1}(0)$ and ${\mathbf{x}}_{2}={\mathbf{x}}_{2}(0)$. As discussed in [38], the $N$-dimensional potential of a diffusion model can be projected down into 1-dimension via a generative trajectory between ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{2}$. We can obtain such a trajectory via circular interpolation between ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{2}$ following
$\tilde{{\mathbf{x}}}(\alpha)=\cos(\alpha){\mathbf{x}}_{1}+\sin(\alpha){\mathbf{x}}_{2}$ (42)
and compute the relative potential (up to a constant) as a function of $\alpha$ by using
$\tilde{u}(\alpha)=u(\tilde{{\mathbf{x}}}(\alpha))=\int^{\pi/2}_{0}\nabla u(\tilde{{\mathbf{x}}}(\alpha))\cdot{\mathbf{v}}\,\mathrm{d}\alpha$ (43)
where ${\mathbf{v}}=-\sin(a){\mathbf{x}}_{1}+\cos(a){\mathbf{x}}_{2}$. See Figure 21 for a visualization of this circular interpolation between two samples. For discrete times, we instead estimate the relative energy as
$\tilde{u}(\bar{\alpha}_{L})\approx\sum^{L}_{i=1}\nabla u(\tilde{{\mathbf{x}}}(\bar{\alpha}_{i}))\cdot{\mathbf{v}}\Delta\alpha$ (44)
given that $\bar{\alpha}=\{0,\dots,\frac{\pi}{2}\}$ and the potential $u(\cdot)$ is obtained from the above Equation 41. Experimentation Details. In our experimentation of Section 7, we utilized a maximum of the top 2048 samples for each of three sample types (i.e., memorized, spurious, and generalized sets), and the training data set corresponding at each training data size $K$. Note, with the exception of the training data set, the samples in the three sets are sorted from least to greatest, based on their distance corresponding to their respective detection metric, e.g., Eq. (13), (14) or (15). We computed the relative energy of each set following Algorithm 2, which is strictly following the formulations detailed above. Note, we randomly selected and utilized an image from the entire training set as part of the relative energy calculation.
To define the average relative energy gap, let $\tilde{u}_{\text{target}}\in\mathbb{R}^{M}$ be a set of $M$ relative potentials computed from using either a memorized, spurious, or generalized set and following Equation 44. Please note that $M$ is capped at $2048$ samples. Since at some training data size $K$, certain sample type sets might not have less than $2048$ samples. For such cases, we use the entirety of that set to compute the relative energy of its samples. Meanwhile, let $\tilde{u}_{\text{data}}\in\mathbb{R}^{M^{\prime}}$ be the set of relative potentials computed from the training dataset at size $K$, where $M^{\prime}=2048$ for $K\geq 2048$. The average relative energy gap is defined as
$\Delta(\tilde{u}_{\text{target}},\tilde{u}_{\text{data}})=\frac{1}{M}\sum^{M}_{i=1}\big{(}\tilde{u}_{\text{target}}\big{)}_{i}-\frac{1}{M^{\prime}}\sum^{M^{\prime}}_{j=1}\big{(}\tilde{u}_{\text{data}}\big{)}_{j}=\mu(\tilde{u}_{\text{target}})-\mu(\tilde{u}_{\text{data}})$ (45)
Using Equation 45, we computed the average relative energy gap and recorded our results in Figs. (8) and (22). The shaded regions reported in these figures represent the standard deviations computed from using $\tilde{u}_{\text{target}}$, and the average relative energy gap as the mean value. Since at some training data size $K$, there might be very few samples in one of the three sets, we opted not to include the relative energy gap of such sets, partly due to their large variance. Similar to the basin of attraction experiment, we determine whether a set is too small if the number of its samples is less than $0.1\%$ of its corresponding synthetic set. Finally, for our interpolation values $\bar{\alpha}=\{0,\dots,\frac{\pi}{2}\}$ detailed in Algorithm 2 and used in Equation 44, we generate $20$ points starting from 0 to $\frac{\pi}{2}$ using linear spacing.
Moreover, since the relative energy computation requires a reference image, we studied the effects of having different reference images on the relative energy gap in Figs. (23), (24), (25), and (26). Overall, there is a slight difference in the values of the relative energy gap across the memorization-generalization transition for all datasets. However, the general trend is still maintained. Specifically, the gap for memorized samples is initially close to zero and quickly diverges as the training data size $K$ increases. Spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential of spurious states quickly diverges as the later stages of generalization are initiated and consequently, where the gap between generalized and training data samples’ relative energy significantly decreases.
Figure 21: Circular interpolation of a target image ${\mathbf{x}}_{1}$ and a reference image ${\mathbf{x}}_{2}$ over $\alpha\in\{0,\dots,\frac{\pi}{2}\}$ for different datasets.
H.1 Algorithm for Computing Relative Energy
1
2
3Inputs
4        Target pattern ${\mathbf{x}}_{1}\in\mathbb{R}^{N}$
5        Reference pattern ${\mathbf{x}}_{2}\in\mathbb{R}^{N}$
6        Score model $s_{\theta}$
7        Interpolation values $\bar{\alpha}$
8        Initial variance $\beta_{0}$
9Assumption: ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{2}$ are samples at $t=0$
10 Compute $\tilde{u}(\bar{\alpha}_{L})$
11        $\tilde{u}(\bar{\alpha}_{L})\leftarrow 0$
12
       $\bar{\alpha}\leftarrow\{0,\cdots,\frac{\pi}{2}\}$
         // Linear Spacing from 0 to $\frac{\pi}{2}$
13
14       $\Delta\alpha\leftarrow\bar{\alpha}_{1}-\bar{\alpha}_{0}$
15
16       for $\alpha\in\bar{\alpha}$ do
17               $\tilde{{\mathbf{x}}}\leftarrow\cos(\alpha){\mathbf{x}}_{1}+\sin(\alpha){\mathbf{x}}_{2}$
18               $\tilde{{\mathbf{v}}}\leftarrow-\sin(a){\mathbf{x}}_{1}+\cos(a){\mathbf{x}}_{2}$
19               $\nabla_{\tilde{{\mathbf{x}}}}\log p_{0}(\tilde{{\mathbf{x}}})\leftarrow-s_{\theta}(\tilde{{\mathbf{x}}},0)/\sqrt{\beta_{0}}$
               $\nabla_{\tilde{{\mathbf{x}}}}u(\tilde{{\mathbf{x}}})\cdot{\mathbf{v}}\,\Delta\alpha\leftarrow\big{(}-\frac{1}{2}\beta_{0}\tilde{{\mathbf{x}}}-\beta_{0}\nabla_{\tilde{{\mathbf{x}}}}\log p_{0}(\tilde{{\mathbf{x}}})\big{)}\cdot{\mathbf{v}}\,\Delta\alpha$
                // See Equation 39
20
               $\tilde{u}(\alpha)\leftarrow\sum^{N}_{j=1}\bigg{(}\nabla_{\tilde{{\mathbf{x}}}}u(\tilde{{\mathbf{x}}})\cdot{\mathbf{v}}\,\Delta\alpha\bigg{)}_{j}$
                // Sum over $N$ pixels
21
               $\tilde{u}(\bar{\alpha}_{L})\leftarrow\tilde{u}(\bar{\alpha}_{L})+\tilde{u}(\alpha)$
                // See Equation 44
22
23
24Output
25        Relative energy $\tilde{u}(\bar{\alpha}_{L})$
Algorithm 2 Compute the Discretized Relative Energy between a target ${\mathbf{x}}_{1}$ and a reference ${\mathbf{x}}_{2}$
H.2 Average Relative Energy Gap of Sample Types for All Datasets
Figure 22: Average energy gap of each sample type and the training samples for the four datasets as the training data size $K$ grows, computed using Equation 44. Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points. Shaded regions show standard deviation of the gap values. With the exception of LSUN-CHURCH, the relative energy of the memorized samples matches that of the training data samples at $K=2$, but quickly diverges as $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
H.3 Average Relative Energy Gap using Different Reference Images for All Datasets
Figure 23: Average energy gap of each sample type and the training samples of CIFAR10 as the training data size $K$ grows, computed using Equation 44 and four different reference images. Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points. Shaded regions show standard deviation of the gap values. The relative energy gap values are slightly different for distinct reference images but the general trend remains. The gap for memorized samples is initially close to zero and quickly diverges as the training data size $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
Figure 24: Average energy gap of each sample type and the training samples of LSUN-CHURCH as the training data size $K$ grows, computed using Equation 44 and four different reference images. Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points. Shaded regions show standard deviation of the gap values. The relative energy gap values are slightly different for distinct reference images but the general trend remains. The gap for memorized samples is initially close to zero and quickly diverges as the training data size $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
Figure 25: Average energy gap of each sample type and the training samples of FASHION-MNIST as the training data size $K$ grows, computed using Equation 44 and four different reference images. Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points. Shaded regions show standard deviation of the gap values. The relative energy gap values are slightly different for distinct reference images but the general trend remains. The gap for memorized samples is initially close to zero (for three of the reference images) and quickly diverges as the training data size $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
Figure 26: Average energy gap of each sample type and the training samples of MNIST as the training data size $K$ grows, computed using Equation 44 and four different reference images. Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself. A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points. Shaded regions show standard deviation of the gap values. The relative energy gap values are slightly different for distinct reference images but the general trend remains. The gap for memorized samples is initially close to zero (for three of the reference images) and quickly diverges as the training data size $K$ increases. Similarly, spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase. However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.

### v2

@@
 Spurious states appear at the boundary of this transition and correspond to emergent attractor states, which are absent in the training set, but, at the same time, have distinct basins of attraction around them.
-Our findings provide: a novel perspective on the memorization-generalization phenomenon in diffusion models via the lens of AMs, theoretical prediction of existence of spurious states, empirical validation of this prediction in commonly-used diffusion models.
-denotes equal contribution.
+Our findings provide: a novel perspective on the memorization-generalization phenomenon in diffusion models via the lens of AMs, theoretical prediction of existence of spurious states, empirical validation of this prediction in commonly-used diffusion models.The code is available at: https://github.com/Lemon-cmd/Diffusion-Models-and-Associative-Memory denotes equal contribution.
 1 Introduction
@@
 CIFAR10 64 (1, 2, 2, 2) 8.9M 128 500,000
+LSUN-CHURCH 64 (1, 1, 2, 2, 4, 4) 27.4M 64 800,000
 LSUN-CHURCH 96 (1, 1, 2, 2, 4, 4) 61.7M 64 800,000
+LSUN-CHURCH 128 (1, 1, 2, 2, 4, 4) 109.7M 64 800,000
 FASHION-MNIST 128 (1, 2, 2) 24.5M 128 400,000
@@
 Lastly, for each of the training set $\mathsf{S}_{\alpha}$, where $\alpha=1,\dots,M$, they are split from the original dataset given a specific size, using the same random seed value of $3407$.
-Figure 9: The fractions of memorized, spurious, and generalized samples in synthetic sets across different training sizes for CIFAR10 and various model’s widths.
+Figure 9: The fractions of memorized, spurious, and generalized samples in synthetic sets across different training sizes for CIFAR10 [63] and various model’s widths.
 As the training data size $K$ increases, memorization decreases while the emergence of spurious patterns follows and the fraction of generalized patterns steadily increase (see top row).
@@
 Additional Experiments.
-Inspired by [45, 46], we extend our investigation on the memorization-generalization transition by studying the effects of varying the parameter size of the diffusion model for CIFAR10.
-The results of this experiment are recorded in Figure 9.
+Inspired by [45, 46], we extend our investigation on the memorization-generalization transition by studying the effects of varying the parameter size of the diffusion model for CIFAR10 [63] and LSUN-CHURCH [66].
+The results of this experiment are recorded in Figs.
+(9) and (10), respectively.
 Specifically, we trained two additional sets of 38 diffusion models, with the initial latent dimensions of 64 and 96 respectively, and repeat the same experimental setup delineated above, see also Table 3.
-For U-Net (64), we identified the critical points, $A$ and $B$, to be $500$ and $8000$ respectively.
-Meanwhile, for U-Net (96), these critical points are $2000$ and $16000$.
+For CIFAR10 and the U-Net (64) models, we identified the critical points, $A$ and $B$, to be $500$ and $8000$ respectively.
+Meanwhile, for the U-Net (96) models, these critical points are $2000$ and $16000$.
 To select $\delta_{m}$ and $\delta_{s}$, we followed the same process detailed above by observing the distance histograms to see where bimodality occurs with respect to the distance value, and manually inspecting the images to ensure there is a minimal amount of false positives in memorized and spurious samples.
-For U-Net (64), the threshold values $\delta_{m}$ and $\delta_{s}$ are both selected to be $0.02$ — while for U-Net (96), the threshold values $\delta_{m}$ and $\delta_{s}$ are chosen as $0.03$ and $0.02$, respectively.
-With the smaller U-Net (64) model, we observe the memorization-generalization transition happens at an earlier stage with respect to the larger U-Net (128) model, where the peak of spurious samples appears at $K=2568$ instead of $K=5862$.
-Similarly, for U-Net (96), the peak of spurious samples happens at an early stage of $K=3931$ instead.
+For U-Net (64) models, the threshold values $\delta_{m}$ and $\delta_{s}$ are both selected to be $0.02$ based on the bimodality of the distance histograms observed at $K=2568$ — meanwhile, for U-Net (96), the threshold values $\delta_{m}$ and $\delta_{s}$ are chosen as $0.03$ and $0.02$ according to distance histograms observed at $K=3931$.
+Similarly, for the LSUN-CHURCH models, we identified the critical points, $A$ and $B$, to be $1000$ and $8000$ for the U-Net (64) models; and for the larger U-Net (128) models, $4000$ and $16000$ as the critical points.
+The threshold values, $\delta_{m}$ and $\delta_{s}$, are both selected to be $0.04$ after observing the bimodality of the distances at $K=3413$ — meanwhile, for U-Net (128), these values are selected to be $\delta_{s}=0.04$ and $\delta_{m}=0.1$ after observing the distance histograms at $K=5241$.
+Figure 10: The fractions of memorized, spurious, and generalized samples in synthetic sets across different training sizes for LSUN-CHURCH [66] and various model’s widths.
+As the training data size $K$ increases, memorization decreases while the emergence of spurious patterns follows and the fraction of generalized patterns steadily increase (see top row).
+The fraction of spurious patterns rises and decreases at the boundary between the memorization and generalization phases (see bottom row).
+Additionally, with a smaller number of parameters, the memorization-generalization transition happens at an earlier stage (or smaller $K$), since the model has a memorization capacity.
+In contrast, as the model’s size increases, the transition happens at an later stage (or larger $K$), where it is able to memorize more training data.
+The numbering, e.g., $64$, $96$, and $128$, denotes the initial latent dimension of the U-Net, see Table 3.
+With the smaller U-Net (64) models, we observe the memorization-generalization transition happens at an earlier stage with respect to the larger U-Net (128) models, where the peak of spurious samples appears at $K=2568$ instead of $K=5862$ for CIFAR10 and $K=1000$ instead of $K=4000$ for LSUN-CHURCH.
+Similarly, for CIFAR10 and its U-Net (96) models, the peak of spurious samples happens at an early stage of $K=3931$ instead.
 Overall, these results indicate that as the parameter size of the diffusion model increases, the memorization capacity of the model is increased and thus, generalization is delayed since the model can memorize more of its training data denoted by the appearance of spurious patterns happening at a larger training data size $K$.
@@
 Hardware Details.
-Training of the models detailed in Table 3 were done using many NVIDIA Tesla V100 GPUs (a single GPU was used to train each model).
+The training of the models detailed in Table 3 was done using NVIDIA Tesla V100 GPUs, where a single GPU was used to train each model.
 Each GPU has 32GB of memory and is linked with IBM Power9 processors which clocked at 3.15 GHz.
 Each CIFAR10 model took roughly 18, 20, and 22 hours to train for different initial latent dimensions, e.g., 64, 96, 128, respectively.
-Meanwhile, each LSUN-CHURCH model requires roughly 56 hours of training.
+Meanwhile, each LSUN-CHURCH model requires roughly 48, 56, and 103 hours of training with respect to the model’s initial latent dimension.
 In contrast, FASHION-MNIST and MNIST models require approximately 17 hours to be fully trained (per model).

### v3

@@
 Abstract
-Hopfield networks are associative memory (AM) systems, designed for storing and retrieving patterns as local minima of an energy landscape.
-In the classical Hopfield model, an interesting phenomenon occurs when the amount of training data reaches its critical memory load — spurious states, or unintended stable points, emerge at the end of the retrieval dynamics, leading to incorrect recall.
-In this work, we examine diffusion models, commonly used in generative modeling, from the perspective of AMs.
-The training phase of diffusion model is conceptualized as memory encoding (training data is stored in the memory).
-The generation phase is viewed as an attempt of memory retrieval.
-In the small data regime the diffusion model exhibits a strong memorization phase, where the network creates distinct basins of attraction around each sample in the training set, akin to the Hopfield model below the critical memory load.
-In the large data regime, a different phase appears where an increase in the size of the training set fosters the creation of new attractor states that correspond to manifolds of the generated samples.
-Spurious states appear at the boundary of this transition and correspond to emergent attractor states, which are absent in the training set, but, at the same time, have distinct basins of attraction around them.
-Our findings provide: a novel perspective on the memorization-generalization phenomenon in diffusion models via the lens of AMs, theoretical prediction of existence of spurious states, empirical validation of this prediction in commonly-used diffusion models.The code is available at: https://github.com/Lemon-cmd/Diffusion-Models-and-Associative-Memory denotes equal contribution.
+Dense Associative Memories (DenseAMs) are generalizations of Hopfield networks, which have superior information storage capacity and can store training data points (memories) at local minima of the energy landscape.
+When the amount of training data exceeds the critical memory storage capacity of these models, new local minima, which are different from the training data, emerge.
+In Associative Memory these emergent local minima are called spurious states, which hinder memory retrieval.
+In this work, we examine diffusion models (DMs) through the DenseAM lens, viewing their generative process as an attempt of a memory retrieval.
+In the small data regimes, DMs create distinct attractors for each training sample, akin to DenseAMs below the critical memory storage.
+As the training data size increases, they transition from memorization to generalization.
+We identify a critical intermediate phase, predicted by DenseAM theory – the spurious states.
+In generative modeling, these states are no longer negative artifacts but rather are the first signs of generative capabilities.
+We characterize the basins of attraction, energy landscape curvature, and computational properties of these previously overlooked states.
+Their existence is demonstrated across a wide range of architectures and datasets.
+Corresponding email: phamb@rpi.edu.
+The code is available from this link.
+denotes equal contribution.
 1  Introduction
-Hopfield networks are the simplest energy-based models that conceptualize memories as attractor states corresponding to local minima of their energy function, and the memory retrieval as dynamical convergence towards those attractor states [1, 2, 3].
-Recently, they have seen a resurgence of interest due to advances in their memorization capacity.
-Notably, Dense Associative Memories (DenseAMs), which are extensions of Hopfield networks with super-linear memory capacity [4, 5] have paved the way for more sophisticated associative memory systems [6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18], driven in part by their connection to the attention mechanism in transformers [19, 20, 13] and to neurobiology [20, 21].
-Simultaneously, generative diffusion models [22] have gained considerable popularity, due to their flexibility and accuracy in modeling high-dimensional distributions for a variety of domains —such as image generation [23, 24, 25], audio [26, 27, 28], video synthesis [29, 30, 31, 32], and other scientific applications.
-However, despite their effectiveness, diffusion models pose challenges related to privacy and security, as concerns grow about their tendency to replicate training data [33, 34, 35].
-Such matters consequently emphasize the need for further understanding of memorization and generalization behaviors in diffusion models.
-Recent works [36, 37, 38] have begun establishing theoretical connections between DenseAMs and generative diffusion models, offering a foundation for bridging the two fields.
-Specifically, it was shown that the logarithm of the probability of the generated samples in diffusion models can be interpreted as the energy function of a commonly used DenseAM model with the softmax activation function.
-This makes it possible to apply theoretical tools developed for associative memory to better understand the computational properties of diffusion models.
-A possible key that can further bridge these two systems and highlight the properties of the memorization-to-generalization transition in diffusion models is the phenomenon of spurious patterns.
-Historically considered as detrimental to pattern recall [39, 40, 41], spurious patterns can be interpreted as combinations or interpolations of stored patterns, hinting at the network’s ability to synthesize new patterns from existing training data.
-This blending of fundamental memories resembles the generalization process in generative models, where learned representations are used to generate novel outputs.
-In this way, spurious patterns offer a fascinating framework for exploring the balance between memorization — where models store exact patterns from training data — and generalization, where they use underlying features to create genuinely new samples [42].
-Studying the conditions under which spurious patterns emerge and how they contribute to the network’s behavior can thus shed light on generalization in both associative memory and contemporary generative models.
-Previous works have explored memorization in generative models through various approaches.
-For instance, [43, 44] proposed general methods to measure memorization, while [33, 34, 45, 46] examined the memorization capacity in diffusion models as a function of the size of the training data.
-Other works have focused on understanding generalization, such as the theoretical estimates of the generalization gap [47], and the spectral analyses [48, 49, 50].
-A recent study [51] proposed that generalization in diffusion models arises as a result of the combination of limited capacity and structural constraints neural network, which results in predictable patterns at the boundary between memorization and generalization.
-Other recent studies [52, 50, 49] explored memorization and generalization through the relationship between the dimensionalities of the ground-truth manifold and the manifold that is learned by the model.
-Meanwhile, [53] explored the evolution of the generated density as a function of the training time to provide an analysis on how mode collapse occurs and leads to the loss of diversity in the generated density.
-Finally, [54] and [55] characterized memorization in diffusion models as “condensation” phase transitions known in the statistical physics of spin glasses and random energy models.
-Although these studies shed some light on memorization and generalization, they do not fully explore the model’s behavior during the transition between these two phases.
-Figure 1: A simple illustration depicting the change in the energy landscape as the size of the training dataset is increased.
-In the small data regime, the model memorizes the training data points as local minima of the energy.
-When the amount of training data exceeds the memory capacity of the model, spurious patterns are formed and training data points are no longer energy minima.
-Subsequent increase of the training set size leads to the generalization phase, which is defined by the formation of continuous manifold of the low energy states.
-Contributions.
-Most of the mentioned works on memorization-to-generalization transition tackled it in a generalization-centric fashion.
-Namely, the memorization phenomenon is typically viewed as a “small side effect” of the diffusion models.
-In our work we adopt a complimentary approach.
-From the beginning, we cast the diffusion modeling pipeline into the associative memory framework.
-The training phase of the diffusion model is conceptualized as writing into the memory operation.
-The generation phase is viewed as an attempt of memory recall.
+Hopfield networks are energy-based Associative Memory (AM) models, which conceptualize memories as attractor states corresponding to local minima of their energy function, and the memory retrieval as dynamical convergence towards such attractors Hopfield (1982, 1984); Amari (1972).
+Modern variants of such networks, called Dense Associative Memories (DenseAMs) Krotov and Hopfield (2016, 2018), have helped revitalize scientific interest in these ideas and paved the way for more sophisticated AM systems Demircigil et al.
+(2017); Agliari et al.
+(2020); Agliari and De Marzo (2020); Albanese et al.
+(2022); Millidge et al.
+(2022); Saha et al.
+(2023); Krotov (2023); Hoover et al.
+(2023a), driven primarily by their connection to the attention mechanism in Transformers Ramsauer et al.
+(2021); Hoover et al.
+(2023a) and neurobiology Krotov and Hopfield (2021); Kozachkov et al.
+(2025).
+Simultaneously, diffusion models (DMs) Sohl-Dickstein et al.
+(2015) have gained popularity, due to their flexibility and accuracy in modeling various high-dimensional distributions Ho et al.
+(2020); Song and Ermon (2019); Song et al.
+(2021); Rombach et al.
+(2022).
+However, despite their effectiveness, DMs pose challenges related to privacy and security, as concerns grow about their tendency to replicate training data alongside a lack of understanding of how generalization arises in these models Somepalli et al.
+(2023a, b); Carlini et al.
+(2023); Wen et al.
+(2024); Jeon et al.
+(2024); Webster (2023).
+Hence, such matters emphasize the need for further understanding of memorization and generalization behaviors of DMs.
+Currently, much of the recent works on the memorization-to-generalization transition of DMs tackle it in a generalization-centric fashion Meehan et al.
+(2020); den Burg and Williams (2021); Somepalli et al.
+(2023a, b); Kadkhodaie et al.
+(2023); Ventura et al.
+(2024); Achilli et al.
+(2024); Kamb and Ganguli (2024); Ross et al.
+(2024); Biroli et al.
+(2024); Cui et al.
+(2025); Achilli et al.
+(2025).
+Namely, the memorization phenomenon is typically viewed as a “small side effect,” alleviated by drastically increasing the training data size Yoon et al.
+(2023); Gu et al.
+(2023).
+Although these studies shed some light on memorization and generalization, they do not fully explore the intermediate transition regime between these two phases.
+Our work adopts a complimentary approach which further ties memorization and generalization together.
+From the beginning, we cast the diffusion modeling pipeline into the AM framework.
+The training phase of diffusion modeling is conceptualized as the operation of writing the training data into the memory.
+The generation phase is viewed as the attempt of memory recall.
 This recall can be successful, resulting in the retrieval of training samples (memorization), or unsuccessful, resulting in the generation of new previously unseen samples (generalization).
-Thus, according to this theory, generalization (creation of genuinely novel samples) is a failure of memory recall.
-This point of view allows us to apply theory developed in DenseAMs for analyzing the memorization-to-generalization transition.
-In DenseAMs, it is well-known that successful memory recall deteriorates as the number of the stored data points increases [4].
-Importantly, the transition from successful to unsuccessful memory recall is marked by the peculiar phenomenon — the emergence of spurious states, see Figure 1.
-Our theory provides a specific and empirically verifiable prediction — spurious states must exist in the conventional diffusion models, which are trained and ran using standard methods.
-In this paper we demonstrate that this is indeed the case.
-Spurious states emerge at the boundary between the memorization and generalization phases, see examples of spurious samples in Figure 2.
-In accord with previous literature, we find that (1) diffusion models undergo a transition from memorization to generalization as the number of training points increases; (2) unlike previous works on diffusion models, at the onset of generalization we detect the emergence of spurious patterns, which have sizable basins of attraction, but are not memorized data points; and lastly, (3) we provide theoretical descriptions distinguishing these spurious states from memorized and generalized patterns in terms of energy landscapes.
-All of these findings suggest a close parallel between diffusion models and DenseAMs above the critical storage capacity.
-Figure 2: Examples of memorized, spurious, and generalized samples in their respective columns for four datasets.
-For each target image (shown on the left), its top-4 nearest neighbors from the training set (top row) and the synthetic set (bottom row) are shown to highlight the novelty or the commonality of the target image with respect to training and synthetic sets.
-Memorized samples are duplicates of the training points.
-Spurious samples are copies of the synthetic set that do not appear in the training set.
-Generalized samples do not belong to either of the two sets, indicating that they are genuinely novel samples that are not generated frequently.
+Consequently, this point of view allows us to apply the theory developed for AMs to the memorization-to-generalization transition in DMs.
+Recent works Hoover et al.
+(2023b); Ambrogioni (2024); Raya and Ambrogioni (2024) have also begun establishing theoretical connections between DenseAMs and DMs, illustrating that the score function of typical DMs can be interpreted as the gradient of a DenseAM’s energy function.
+These prior works, however, do not discuss a cornerstone phenomenon of AMs – the emergence of spurious states – which can further bridge both DenseAMs and DMs.
+Historically considered as detrimental to reliable memory recall Hopfield (1982); Hopfield et al.
+(1983); Hopfield (1984); Amit et al.
+(1985); Abu-Mostafa and St.
+Jacques (1985), spurious states were viewed as negative artifacts of AM networks.
+In contrast, for generative modelling these same spurious states serve as the first signatures of generalization capabilities, thus becoming a desirable emergent phenomenon Kalaj et al.
+(2025).
+In accord with previous literature, we empirically find that DMs undergo a transition from memorization to generalization as the number of training points increases on realistic datasets.
+Unlike previous works, at the onset of generalization we detect the emergence of spurious patterns, which have sizable basins of attraction, but do not correspond to memorized data points.
+No longer negative artifacts, spurious states are now observed to be positive indicators of early generalization in DMs and they may also exist even during the full generalization stage of DMs, including Stable Diffusion Rombach et al.
+(2022).
+Lastly, we provide theoretical descriptions distinguishing these peculiar states from memorized and generalized patterns, in terms of energy landscapes.
+Our findings suggest a close parallel between two seemingly different systems, DMs and DenseAMs, studied by two disjoint communities.
+Specifically, the DenseAM’s perspective on DMs can illuminate some of the computational properties of DMs at the boundary of the memorization-generalization transition.
+Figure 1: Panel (A) schematically illustrates the change in the energy landscape as the size of the training dataset is increased.
+In the small data regime, the model stores the training data points as local minima of the energy.
+When the amount of training data exceeds the model’s memory capacity, spurious patterns are formed and training data points are no longer energy minima.
+Subsequent increase of the training data size leads to the generalization phase, defined by the formation of continuous manifold of the low energy states.
+Examples of memorized, spurious, and generalized samples in their respective columns for four datasets (MNIST, FASHION-MNIST, CIFAR10, and LSUN-CHURCH) are provided in Panel (B), see Sec.
+(4.1) for our definitions of these three sample types.
+For each target image (shown on the left), its top-4 nearest neighbors from the training set (top row) and the synthetic set (bottom row) are shown to highlight the novelty and commonality of the target image with respect to its training and synthetic sets.
+To help highlight the novelty of spurious patterns, we provide rectangular markers guiding where the features differ in their corresponding nearest neighbors.
 2  Diffusion Models and Dense Associative Memories
-Given a dataset of i.i.d.
-samples ${\mathbf{y}}\in\mathbb{R}^{N}$ drawn from a target data distribution $p({\mathbf{y}})$, diffusion models are a class of generative models, which aims to approximate $p({\mathbf{y}})$ by placing a reversible process that maps data to noise and back.
-The mapping to noise (or forward process) is described by the following Itô Stochastic Differential Equation (SDE) [25]:
-$\mathrm{d}{\mathbf{x}}_{t}=\mathbf{f}({\mathbf{x}}_{t},t)\mathrm{d}t+{g}(t)\mathrm{d}{\mathbf{w}}_{t}$ (1)
-which transforms the given data distribution (${\mathbf{x}}_{0}={\mathbf{y}}$) into a simpler one, such as an isotropic Gaussian distribution.
-Here, ${\mathbf{w}}_{t}$ is the standard Wiener process (or Brownian motion) and $\mathbf{f}({\mathbf{x}}_{t},t)$ denotes the drift term that guides the diffusion process, which we will assume to be zero for the most part of this paper.
-Meanwhile, $g(t)$ represents the diffusion coefficient that controls the noise at each time step $t\rightarrow T$.
-The reverse process removes the injected noise and is described as
-$\mathrm{d}{\mathbf{x}}_{t}=[\mathbf{f}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})]\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (2)
+Motivated by Hoover et al.
+(2023b); Ambrogioni (2024); Raya and Ambrogioni (2024), we first provide a theoretical link between DMs and DenseAMs through the perspective of energy.
+Given a dataset of samples ${\mathbf{y}}\in\mathbb{R}^{N}$ drawn from a target data distribution ${\mathbf{y}}\overset{\text{i.i.d}}{\sim}p({\mathbf{y}})$, we rely on the following stochastic differential equation (SDE) which describes the forward process Song et al.
+(2021):
+$\mathrm{d}{\mathbf{x}}_{t}=\mathbf{f}({\mathbf{x}}_{t},t)\mathrm{d}t+{g}(t)\mathrm{d}{\mathbf{w}}_{t},$ (1)
+which transforms $p({\mathbf{y}})$, where ${\mathbf{x}}_{0}={\mathbf{y}}$, into a simpler distribution, e.g., isotropic Gaussian.
+Here, ${\mathbf{w}}_{t}$ is the standard Wiener process and $\mathbf{f}({\mathbf{x}}_{t},t)$ denotes the drift term that guides the diffusion process, which we assume to be zero for most of the paper.
+Meanwhile, $g(t)$ denotes the diffusion coefficient controlling the noise at each time step $t\rightarrow T$.
+The reverse, denoising, or generative process is described as
+$\mathrm{d}{\mathbf{x}}_{t}=[\mathbf{f}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})]\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t},$ (2)
 where $\bar{{\mathbf{w}}}_{t}$ is the standard Wiener process.
-To effectively solve this equation, one must reliably estimate the score $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$ via parameterization of the score as a neural network $s_{\theta}({\mathbf{x}},t)$, where $\theta$ denotes neural network’s trainable parameters.
-This network is trained using methods for denoising score matching across multiple times steps [23, 25, 56, 57, 58].
-Our theoretical arguments pertain to the idealistic setting in which this neural network is capable of learning the true empirical distribution obtained by adding noise to the training data.
-This limiting case is often adopted in theoretical studies of diffusion models, see [54, 59, 60, 55, 61].
-From the practical perspective, this is achieved if the neural network is large (over-parametrized) and trained for a long time, an assumption that may deviate from the practical use cases.
-This simplification allows us to make a theoretical prediction that diffusion models must generate spurious states.
-It is highly non-trivial that this theoretical prediction can be extrapolated to the regime in which the size of the score-modeling neural network is finite and the amount of training data is large.
-The main contribution of our work is the empirical demonstration that this extrapolation is valid, and spurious states indeed exist in practically used settings.
+To effectively solve this equation, one must reliably estimate the score $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})=-\nabla_{{\mathbf{x}}_{t}}E({\mathbf{x}}_{t},t)$ via training a neural network $s_{\theta}({\mathbf{x}}_{t},t)$, where $\theta$ denotes its set of trainable parameters, using methods for denoising score matching across multiple time steps Ho et al.
+(2020); Song et al.
+(2021); Hyvärinen (2005); Vincent (2011).
 Consider the training data distribution in the variance-exploding setting of $f({\mathbf{x}}_{t},t)=0$ and $g(t)=\sigma$.
-In this case, the marginal probability distribution of new samples can be computed exactly as
-$p({\mathbf{x}}_{t},t)=\underset{{\mathbf{y}}\sim\text{data}}{\mathbb{E}}\Bigg{[}\frac{1}{(2\pi\sigma^{2}t)^{\frac{N}{2}}}\exp\Big(-\frac{\lVert{\mathbf{x}}_{t}-\mathbf{y}\rVert^{2}_{2}}{2\sigma^{2}t}\Big{missing})\Bigg{]}$ (3)
-Assuming the empirical distribution of the data $p(\mathbf{y})=\frac{1}{K}\sum\limits_{\mu=1}^{K}\delta^{(N)}(\mathbf{y}-\bm{\xi}^{\mu})$, where $\bm{\xi}^{\mu}$ represents an individual data point (with data size $K$), this marginal distribution can be written as
+In this case, the marginal probability distribution of noisy samples at time $t$ can be computed exactly as
+$p({\mathbf{x}}_{t},t)=\underset{{\mathbf{y}}\sim\text{data}}{\mathbb{E}}\Bigg[\frac{1}{(2\pi\sigma^{2}t)^{\frac{N}{2}}}\exp\Big(-\frac{\lVert{\mathbf{x}}_{t}-\mathbf{y}\rVert^{2}_{2}}{2\sigma^{2}t}\Big)\Bigg].$ (3)
+Assuming the empirical distribution of the data $p(\mathbf{y})=\frac{1}{K}\sum\limits_{\mu=1}^{K}\delta^{(N)}(\mathbf{y}-\bm{\xi}^{\mu})$, where $\bm{\xi}^{\mu}$ represents an individual data point (belonging to a dataset with training data size $K$), this marginal distribution can be written as
+$p({\mathbf{x}}_{t},t)=\frac{1}{K}\sum\limits_{\mu=1}^{K}\frac{1}{(2\pi\sigma^{2}t)^{\frac{N}{2}}}\exp\Big({-\frac{\lVert{\mathbf{x}}_{t}-\bm{\xi}^{\mu}\rVert^{2}_{2}}{2\sigma^{2}t}}\Big)\overset{\text{def}}{\equiv}\exp\Bigg(-{\frac{E^{\text{DM}}({\mathbf{x}}_{t},t)}{2\sigma^{2}t}}\Bigg),$ (4)
+where we also defined the energy $E^{\mathrm{DM}}$ of the DM, up to $\mathbf{x}$}–independent terms, to be
+$E^{\text{DM}}({\mathbf{x}}_{t},t)=-2\sigma^{2}t\log\bigg[\sum\limits_{\mu=1}^{K}\exp\Big(-\frac{\lVert{\mathbf{x}}_{t}-\bm{\xi}^{\mu}\rVert^{2}_{2}}{2\sigma^{2}t}\Big)\bigg].$ (5)
+As already observed in Ambrogioni (2024), Eq. (5) is closely related to a commonly studied DenseAM’s model, see for example Saha et al. (2023):
+$E^{\text{AM}}({\mathbf{x}})=-\beta^{-1}\log\bigg[\sum\limits_{\mu=1}^{K}\exp\Big(-\beta\lVert{\mathbf{x}}-\bm{\xi}^{\mu}\rVert^{2}_{2}\Big)\bigg].$ (6)
+A direct comparison of Eqs.
+(5) and (6) highlights a close connection between these two seemingly different frameworks.
+Specifically, the data points $\bm{\xi}^{\mu}$ in the DM framework play the role of memories in the AM formulation.
+The variance of the noise added during the forward process (1) plays the role of the effective temperature $\beta^{-1}$ (which controls the separation of memories), while the reverse process (2) corresponds to memory retrieval dynamics Hoover et al.
+(2023b); Ambrogioni (2024).
+For a single data point $\bm{\xi}^{1}$ or $K=1$, Eq.
+(6) has a single local minimum at that stored pattern, where the shape of this energy landscape is independent of the value of $\beta$.
+This is an example of a memorized state.
+In contrast, for $K=2$, there are several possibilities.
+The energy landscape has two local minima when $\beta\rightarrow\infty$ is large.
+But, for finite values of $\beta$, there exist configurations of patterns such that Eq.
+(6) has only one local minimum:
+$\bm{\eta}=\operatorname*{arg\,min}_{{\mathbf{x}}}E^{\text{AM}}({\mathbf{x}}),$ (7)
+such that $\bm{\eta}\neq\bm{\xi}^{1}$ and $\bm{\eta}\neq\bm{\xi}^{2}$.
+This is an example of a spurious state.
+Such states are ubiquitous in AM models, when the amount of training data (memories) crosses the critical memory storage capacity of the model (which depends on the correlations between the training data points).
+Thus, the following question can be posed: Is it possible to observe spurious states in DMs trained using score matching objective when the energy landscape is modeled by the score-predicting neural network?
+3  Discussion
+We argue below that the answer to the above question is positive, and the spurious states indeed emerge, as DMs progress from memorization to generalization as the training data size $K$ increases.
+As we observed numerically in Fig.
+(4A), when $K$ is small, DMs initially memorize their training data points and store them as local minima of their energy function.
+In the large data regime, a different phase appears where a sufficient increase of $K$ fosters the creation of new attractor states corresponding to the manifold of the generated samples.
+Spurious states appear at the boundary of this transition and correspond to emergent attractor states of the energy function, absent in the training set, but at the same time, having distinct basins of attraction around them.
+Subsequent increases of $K$ leads to a decrease of the number of memorized and spurious states, and the creation of generalized states, points that occupy the continuous manifold.
+This sequence of transitions is illustrated in Fig.
+(1A).
+In order to empirically establish this sequence of transitions, we have trained a set of conventional diffusion models on datasets of gradually increasing sizes.
+Each of the trained models is characterized by its training set $\mathsf{S}$, and a synthetic set of generated samples drawn from this model $\mathsf{S}^{\prime}$ (the synthetic set is assumed to be very large).
+A sample $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}$ is memorized if it has duplicates in both training and synthetic sets.
+This is clear from the three universality classes shown in Fig.
+(1A).
+Since memorized samples have a basin of attraction around them, the same sample should appear in the synthetic set more than once, and that generated sample should be a copy of the training data point.
+A sample $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}$ is spurious if it has duplicates in the synthetic set, but not in the training set.
+Again, this is clear from Fig.
+(1A): spurious samples have a well-defined basin of attraction around them, but the bottom of that basin corresponds to a newly generated sample – not the training example.
+Finally, the generalized sample $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}$ has no duplicates in either the training or synthetic sets.
+Generalized samples occupy flat manifolds in the configuration space.
+For this reason, every generation produces a slightly different sample resulting in absence of duplicates in $\mathsf{S}^{\prime}$.
+Additional details pertaining to these detection metrics are given in section 4.
+Examples of generated samples belonging to these three categories are shown in Fig.
+(1B).
+For each sample, the top-4 nearest neighbors from the training set (top row) and synthetic set (bottom row) are shown.
+It is instructive to visually examine some of the samples classified as spurious.
+For instance, the first spurious sample in Fig.
+(1B) (the model was trained on CIFAR-10 dataset Krizhevsky et al.
+(2014)) is a plane that has a large wing (highlighted by the cyan frame), which is absent in the training set, but has an exact duplicate in the synthetic set.
+Another interesting example is shown in the 4-th row (the model was trained on FASHION-MNIST dataset Xiao et al.
+(2017)).
+There are many examples of shirts in the training set with either two sleeves, or no sleeves.
+The generated sample shows only one sleeve.
+This sample has exact duplicates in the synthetic set, indicating the existence of a well-defined basin of attraction around it, but does not appear in the training set.
+These examples illustrate the general phenomenon, which is predicted by AM perspective and is clearly observed empirically for conventional diffusion models trained with the score matching objective.
+The abundance of memorized, spurious, and generalized samples is quantified as a function of training data set size in Fig.
+(4A).
+As training set gets larger, the amount of memorized samples decreases, while the amount of generalized samples increases.
+The spurious states appear as a peak at the boundary of this transition.
+The distinct nature of these three different states is further substantiated by the hierarchy of the log-volume of their basins of attraction in Fig.
+(4B) and their energy curvature spectra in Fig.
+(5A).
+Regarding the basin volume, memorized samples consistently exhibit the largest log-volume, followed by spurious samples, while generalized samples have near-zero volume.
+This is in accord with the intuitive expectation for the energy profiles in Fig.
+(1A).
+Memorized and spurious samples should have substantial basins of attraction around them, while generalized samples should have almost vanishing basins, since they occupy an almost flat energy manifold.
+The energy curvature spectral analysis reveals a similar finding: as the training data size $K$ increases: the memorized and spurious samples tend to have higher curvature (indicated by their fewer near-zero singular values) compared to the generalized samples.
+Initially, our studies focused on relatively small scale datasets: MNIST Deng (2012), FASHION-MNIST Xiao et al.
+(2017), CIFAR10 Krizhevsky et al.
+(2014), and LSUN-CHURCH Yu et al.
+(2015).
+Focusing on these settings allowed us to train a dense grid of models with varying size of the training set and carefully examine the existence of duplicates in the training and synthetic sets.
+Such a careful analysis would be intractable for large-scale real world scenarios such as Stable Diffusion Rombach et al.
+(2022): training these models requires substantial compute resources, the training set is large, and the synthetic set needs to be even larger in order to reliably perform the analysis similar to what was described above.
+At the same time, the analysis based on the curvature of the energy landscape around generated samples is more feasible.
+Thus, we have identified several candidate samples from Stable Diffusion that fall into three categories: sharp energy curvature, intermediate energy curvature, and almost flat energy.
+Examples of these samples together with their singular values of the score function Jacobian are shown in Fig.
+(5B).
+For each sample a language prompt was given to the model and the generated samples are shown together with the training image corresponding to that prompt.
+High curvature samples (candidate memorized states) are exact duplicates of the training image.
+Intermediate curvature samples (candidate spurious states) are somewhat similar to the training image, but deviate from it in certain aspects (color and pattern of the rug, color of the curtains).
+Finally, small curvature samples (candidate generalized states) demonstrate a greater degree of variability and larger deviations from the training image.
+Qualitatively, these results align with our primary findings on the carefully studied small dataset models.
 […]
-where we also defined the energy $E^{\mathrm{DM}}$ of diffusion model, which up to state– or ${\mathbf{x}}$–independent terms is equal to
-$E^{\text{DM}}({\mathbf{x}}_{t},t)=-2\sigma^{2}t\log[\sum\limits_{\mu=1}^{K}\exp\Big(-\frac{\lVert{\mathbf{x}}_{t}-\bm{\xi}^{\mu}\rVert^{2}_{2}}{2\sigma^{2}t}\Big{missing})\bigg{]}$ (5)
-As already observed in [37], this energy function (5) is closely related to the class of models called DenseAMs, or Modern Hopfield Networks, which are large memory storage variants of classical Hopfield networks [4, 62].
-The core idea behind DenseAMs is to design an energy function that sharply peaks around the intended memory patterns.
-These networks can store and retrieve a much larger number of memories, compared to the classical Hopfield networks, which scales super-linearly (and possibly exponentially) to the size of the network and decouples information storage capacity from the dimensionality of the data [4, 6].
-Of particular interest here is the DenseAM model studied in [12] (see also [11]):
-$E^{\text{AM}}({\mathbf{x}})=-\beta^{-1}\log[\sum\limits_{\mu=1}^{K}\exp\Big(-\beta\lVert{\mathbf{x}}-\bm{\xi}^{\mu}\rVert^{2}_{2}\Big{missing})\bigg{]}$ (6)
-where $\beta$ is the inverse “temperature”, which controls the steepness of the energy landscape around memories $\bm{\xi}^{\mu}$.
-A direct comparison of the energy formulas (5) and (6) highlights a close connection between these two frameworks.
-Specifically, the data points in the diffusion model framework play the role of memories in the associative memory formulation.
-The variance of the noise added during the forward diffusion process (1) plays the role of the effective temperature $\beta^{-1}$.
-Finally, the reverse denoising process (2) corresponds to memory retrieval dynamics [36, 37].
-3 Memorization-to-Generalization in 2D
-To better illustrate this connection, it is instructive to investigate a simple 2-dimensional toy model, see Figure 3, which exhibits many aspects of the memorization-generalization transition.
-Imagine that the training data lies on a unit circle.
-We are interested in exploring how the shape of the energy function (6) changes as the number of training data points increases.
-$\bm{[K=1]}$ For a single data point $\bm{\xi}^{1}$, the energy function (6) has a single local minimum at that stored pattern.
-$\begin{split}E^{\text{AM}}({\mathbf{x}})&=-\beta^{-1}\log[\exp\Big(-\beta\lVert{\mathbf{x}}-\bm{\xi}^{1}\rVert^{2}_{2}\Big{missing})\Big{]}=\lVert{\mathbf{x}}-\bm{\xi}^{1}\rVert^{2}_{2}\end{split}$ (7)
-The shape of the energy landscape is independent of the value of $\beta$.
-Figure 3: Energy landscape evolution for the 2D toy model as training data size $K$ increases.
-Models trained at $K\in\{2,9,1000\}$ use standard VE-SDE based diffusion pipeline with training data sampled from the unit circle, shown in black for $K\in\{2,9\}$.
-Generated samples are shown alongside the learned score field $s_{\theta}({\mathbf{x}}_{t},t)$ done via a neural network, aligned with the negative gradient of the energy (5).
-Hierarchical clustering identifies structure within the generations, with cluster centroid energies visualized by color and numerical value.
-The rightmost panel shows the exact solution as $K\rightarrow\infty$ derived in Equation 12.
-As $K$ grows, the model initially memorizes individual data points, forming isolated basins.
-Around $K=9$, spurious patterns emerge — distinct low-energy attractors not present in the data — which mark the onset of generalization.
-At large $K$, the model enters a fully generalized regime, where low-energy states lie on a flat, continuous manifold shown in Figure 1.
-$\bm{[K=2]}$ For two patterns, the energy is given by
-$\begin{split}E^{\text{AM}}({\mathbf{x}})=-\beta^{-1}&\log[\exp\Big(-\beta\lVert{\mathbf{x}}-\bm{\xi}^{1}\rVert^{2}_{2}\Big{missing})+\exp\Big(-\beta\lVert{\mathbf{x}}-\bm{\xi}^{2}\rVert^{2}_{2}\Big{missing})\Big{]}\end{split}$ (8)
-For $\beta\rightarrow\infty$, this energy has two local minima corresponding to the stored patterns.
-Meanwhile, for finite values of $\beta$, there exist configurations of patterns such that the energy has only one local minimum:
-$\bm{\eta}=\operatorname*{arg\,min}_{{\mathbf{x}}}E^{\text{AM}}({\mathbf{x}})$ (9)
-such that this local minimum is different from either of the two stored patterns: $\bm{\eta}\neq\bm{\xi}^{1}$ and $\bm{\eta}\neq\bm{\xi}^{2}$.
-This “emergent” local minimum is the spurious state, see also the cartoon illustration in Figure 1.
-$\bm{[K\rightarrow\infty]}$ Finally, consider the case when the number of training data points is infinite. In this case they can be described by a continuous density of states:
-$p({\mathbf{y}})=\frac{1}{\pi}\delta\big{(}y_{1}^{2}+y_{2}^{2}-1\big{)}$ (10)
-The probability of the generated data is proportional (up to terms independent of the state ${\mathbf{x}}$) to
-[…]
-where $I_{0}(\cdot)$ is a modified Bessel function of the first kind, see Appendix A for more details on the derivation of Equation 11. Thus, the energy of the model is given by
-$\begin{split}E^{\text{AM}}(R,\phi)&=R^{2}+1-\frac{1}{\beta}\log[I_{0}(2\beta R)\big{]}\underset{\beta\rightarrow\infty}{\approx}(R-1)^{2}\end{split}$ (12)
-where $R$ is the radius of the unit circle and $\phi$ is the polar angle.
-Notice, the dependence on the polar angle $\phi$ in Equation 12 completely disappears from the final result.
-The local minima of the resulting energy function form a continuous manifold, corresponding to $R=1$.
-Data samples from the model occupy the vicinity of that manifold.
-This behavior describes the fully generalized phase.
-In the limit of large $\beta$ the energy landscape is described by a parabola centered around $R=1$.
-For a single training data point, this simple model exhibits memorization.
-With a few training data points, the model can either be in the memorization or spurious phase, depending on the location of those points and the value of $\beta$.
-For a very large number of training data points, the model is in the generalization phase.
-All effects perfectly align with the cartoon illustration in Figure 1.
-Overall, this calculation entirely ignores the neural network, which is used to model the score and approximate the empirical energy.
-There exists a logical possibility that the introduction of the neural network would entirely change the progression of distinct behaviors: memorization, spurious, generalization.
-To test this, we trained a set of diffusion models in this precise setting.
+4  Results
+The main result of our work is the empirical demonstration that DMs trained with standard methods generate spurious samples at the onset of memorization-to-generalization transition.
+The existence of these samples is a cornerstone of energy-based AM models.
+By casting DMs’ generation step into DenseAM setting, one can theoretically expect the emergence of these states at the transition boundary.
+We have characterized computational properties of memorized, spurious, and generalized samples by studying their abundance, volume of the basins of attraction, and curvature of the energy landscape as a function of the DM training set size.
+Spurious samples in DMs have well-defined basins of attraction around them and have not been discussed in prior literature.
+Our theoretical arguments pertain to the idealistic setting in which the neural network $s_{\theta}({\mathbf{x}}_{t},t)$ is capable of learning the true empirical distribution obtained by adding noise to the training data.
+This limiting case is often adopted in theoretical studies of DMs, see Biroli et al.
+(2024); Biroli and Mézard (2023); Achilli et al.
+(2025); Vastola (2025).
+From a practical perspective, this is achieved if $s_{\theta}({\mathbf{x}}_{t},t)$ is over-parametrized and trained for a long time.
+This simplification allows us to make a theoretical prediction that DMs must generate spurious states.
+In practical settings, the size of the score modeling neural network is finite.
+It is non-trivial that the theoretical prediction obtained in the regime of an infinitely large neural network can be extrapolated to the regime in which the size of $s_{\theta}({\mathbf{x}}_{t},t)$ is finite and the amount of training data is large.
+A possible perspective on our contribution is the empirical demonstration that this extrapolation is valid, and spurious states indeed exist in practically used settings.
+In the AM literature, specifically on Hopfield networks, the concept of spurious states is often entangled with the phenomenon of mixture states, which are linear combinations of a certain number of memory vectors.
+While in certain kinds of AM models these two notions can be closely related, it is important to emphasize that they are different concepts.
+Spurious states are defined by the existence of a basin of attraction around them.
+They may or may not be mixture states.
+The main message of our work is to establish that spurious states exist in DMs.
+We do not claim that these spurious states are necessarily mixture states.
+It is up to the model, defined by the score modeling neural network, how the spurious states are constructed from the training data.
+Careful characterization of the memorization-generalization transition and the computational aspects of spurious states that define the onset of this transition may offer future insights for guiding the development of methods that can mitigate DMs’ memorization.
+4.1  Memorization-to-Generalization Transition
+Toy Model.
+An instructive example, which can reinforce the energy connection between Eqs.
+(5) and (6), can be shown via a simple 2-dimensional toy model exhibiting many aspects of the memorization-to-generalization transition.
+Imagine that the training data lies on a unit circle, see Fig.
+(2).
+We are interested in exploring how the shape of $E^{\text{DM}}$ (6) changes as the number of training data points increases.
+Consider the case when the number of training data points is infinite or $K\rightarrow\infty$.
+Using Eq.
+(6), the energy of this toy model can be derived as
+$\begin{split}E^{\text{AM}}(R,\phi)&=R^{2}+1-\frac{1}{\beta}\log\big[I_{0}(2\beta R)\big]\underset{\beta\rightarrow\infty}{\approx}(R-1)^{2},\end{split}$ (8)
+where $R$ is the radius of the unit circle, $\phi$ is the polar angle, and $I_{0}(\cdot)$ is a modified Bessel function of the first kind.
+The dependence on $\phi$ in Eq.
+(15) completely disappears from the final result, since the local minima of the resulting energy function form a continuous manifold, described by a parabola centered around $R=1$ as the inverse temperature $\beta\rightarrow\infty$.
+This behavior describes the fully generalized phase of this toy model.
+Please refer to the Supplementary for Eq.
+(15) derivation.
+To test the resemblance of $E^{\text{DM}}$ and $E^{\text{AM}}$, we trained a set of DMs in this precise setting.
 Each model is trained on a dataset of size $K$.
-The resulting energy landscapes are shown in Figure 3.
-For small number of data points, e.g., $K=2$, the diffusion model exhibits memorization.
+The resulting energy landscapes are shown in Fig.
+(2).
+For a small data size of $K=2$, the DM exhibits memorization.
 The local minima of the energy correspond to the training data points.
-Importantly, at $K=9$, we are able to observe the first signs of spurious states.
+Importantly, at $K=9$ we can observe the first signs of spurious states.
 At this stage, the model begins to learn emergent (different from the training data) local minima of the energy.
-Subsequent increase of the size of the training set leads to fully generalized behavior, which is illustrated for $K=1000$ data points.
-At that stage, all of the samples from the model live in close proximity of the exact data manifold.
-The right panel shows the analytical expression for the energy landscape, defined by Equation 12.
+Subsequent increase of the size of the training set leads to fully generalized behavior, illustrated at $K=1000$.
+At this stage, all of the samples from the model live in close proximity of the exact data manifold.
+The right panel shows the analytical expression of the energy landscape defined by Eq.
+(15).
 Thus, the conventional diffusion modeling pipeline agrees very well with the theoretical prediction of the empirical energy.
-Please refer to Appendix B for full details on the toy model experiment.
-4 Detection Metrics
-Figure 4: Different sample types across the memorization-to-generalization transition for CIFAR10 [63].
+Figure 3: Different sample types across the memorization-to-generalization transition for CIFAR10.
 The grey histogram shows the distances between synthetic samples and their nearest neighbors from the synthetic set $\mathsf{S}^{\prime}$.
@@
 For each generated sample, top-4 nearest neighbors from the training set are shown in the top row, and those from the synthetic set are shown in the bottom row.
-Training set size $K=7310$ was used in this figure (peak of the frequency of spurious states), but the discussed phenomena are general and largely independent of this specific value.
+Training set size $K=7310$ is used in this figure, but the discussed phenomena are general and largely independent of this specific value.
 The fraction of the memorized, spurious, and generalized samples in the pool of all generated samples is shown in the bottom left panel as a function of the training set size.
 The inset shows amplified spurious fraction (green curve).
-The main goal of our work is to establish the existence of spurious states in diffusion models trained on natural high-dimensional datasets, not just toy model problems.
-With this goal in mind we design detection metrics that can classify any generated sample into one of the three categories: memorized, spurious, and generalized samples.
-Fundamentally, these metrics defined below rely on two datasets: $\mathsf{S}$ — the training dataset used to train the diffusion model, and $\mathsf{S}^{\prime}$ — the synthetic dataset generated from the fully trained diffusion model.
-The size of the synthetic set $\mathsf{S}^{\prime}$ is assumed to be much bigger than the training set size.
-The core intuition comes from the typical energy landscape around these three distinct kinds of generated samples, see Figure 1.
-Memorized samples are elements of the synthetic set $\mathsf{S}^{\prime}$, which have duplicates in the training set $\mathsf{S}$.
-Spurious samples are elements of the synthetic set $\mathsf{S}^{\prime}$, which have duplicates in the synthetic set $\mathsf{S}^{\prime}$, but do not have duplicates in $\mathsf{S}$.
-In other words, the same spurious sample appears in the synthetic set more than once.
-Finally, generalized samples are unique elements of $\mathsf{S}^{\prime}$, which do not have duplicates in either the synthetic set $\mathsf{S}^{\prime}$ nor the training set $\mathsf{S}$.
-With this intuition in mind we examine two histograms, see Figure 4.
-First, for every element of the synthetic set $\mathsf{S}^{\prime}$ the distance to the closest nearest neighbor from the training set $\mathsf{S}$ is shown by the olive histogram.
-Second, for every element of the synthetic set $\mathsf{S^{\prime}}$ the distance to the closest nearest neighbor in the synthetic set $\mathsf{S}^{\prime}$ is shown by the gray histogram.
-Both histograms exhibit a clear bimodal shape, indicating that at least two groups of samples are present in the synthetic set.
-This bimodal structure makes it possible to select the memorized $\delta_{m}$ and spurious $\delta_{s}$ thresholds to separate the samples forming the two peaks in each histogram.
-The left peak of the olive histogram is composed of synthetic samples that are at a small distance away from the training set $\mathsf{S}$, which can be considered as memorized samples.
-The left peak of the gray histogram contains synthetic samples that have duplicates in the synthetic set $\mathsf{S}^{\prime}$, indicating they can be memorized if the same sample appears in the training set, or otherwise, spurious.
-The right peak of the gray histogram contains generalized samples, and the right peak of the olive histogram contains generalized and spurious samples.
-These observations make it possible to formally define the following detection metrics.
-Metric 4.1 (Memorization Detection).
-Following [45], we define the memorization detection metric $\mathcal{M}$.
-Given a target pattern $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}$ and its first nearest neighbor ${\mathbf{x}}_{1}$ extracted from the training set $\mathsf{S}$ according to a distance measure $d(\cdot,\cdot)$, where a small distance corresponds to high similarity, metric $\mathcal{M}$ detects sample $\hat{{\mathbf{x}}}$ as memorized if $d(\hat{{\mathbf{x}}},{\mathbf{x}}_{1})$ is small.
-$\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S})=\mathbb{I}\bigg{(}d(\hat{{\mathbf{x}}},{\mathbf{x}}_{1})\leq\delta_{m}\bigg{)}$ (13)
-where $\mathbb{I}$ represents the indicator function and $\delta_{m}\in\mathbb{R}$ is a threshold defined by the distance histograms, see Figure 4.
-Metric 4.2 (Spurious Detection).
-Given a synthetic sample $\hat{{\mathbf{x}}}$ and its first nearest neighbor from the synthetic set ${\mathbf{x}}^{\prime}_{1}\in\mathsf{S}^{\prime}$, the spurious pattern detection metric $\mathcal{S}$ identifies instances where the model generates outputs that do not belong to the training set but have high similarity with samples from the synthetic set (left peak of the gray histogram in Figure 4 after the exclusion of memorized samples).
-$\mathcal{S}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime})=\mathbb{I}\bigg{(}d(\hat{{\mathbf{x}}},{\mathbf{x}}^{\prime}_{1})\leq\delta_{s}\bigg{)}\land\neg\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S})$ (14)
-where $\delta_{s}\in\mathbb{R}$ is a threshold value chosen based on the gray distance histogram, see Figure 4.
-Metric 4.3 (Generalization Detection).
-In contrast, the sample $\hat{{\mathbf{x}}}$ is generalized if it is neither memorized nor spurious.
-$\mathcal{G}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime})=\neg\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S})\land\neg\mathcal{S}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime})$ (15)
-5 Memorization - Spurious - Generalization Transition
+Natural Datasets.
+However, the main goal of our work is to also establish the existence of spurious states in DMs trained on natural high-dimensional datasets.
+With this goal in mind, we design detection metrics that can classify any generated sample into one of three categories – memorized, spurious, and generalized samples – to aid in our analyses.
+Fundamentally, these metrics defined below rely on two datasets: the training dataset $\mathsf{S}$ used to train the DM, and the synthetic dataset $\mathsf{S}^{\prime}$ generated from the fully trained DM.
+The size of $\mathsf{S}^{\prime}$ is assumed to be much bigger than $\mathsf{S}$.
+Our core intuition is rooted in the energy landscape, where its topography dictates the nature of the generated samples, as shown in Fig.
+(1A).
+Each sample type corresponds to a distinct feature of this landscape.
+Memorized samples originate from the basins of attraction the model creates around individual data points from the training set $\mathsf{S}$.
+Because these are strong (having large basins), the generation process frequently converges to exact memorized samples at small data sizes, producing outputs in the synthetic set $\mathsf{S}^{\prime}$ that are perfect duplicates of training data.
+In contrast, spurious samples arise from new emergent basins of attraction, formed during training, which do not correspond to any specific sample in $\mathsf{S}$.
+These emergent minima are also stable attractors, which explains why they appear as duplicates within $\mathsf{S}^{\prime}$ but are absent from $\mathsf{S}$.
+Finally, generalized samples are drawn from flatter regions of the energy landscape.
+Since these areas support a continuum of low-energy states rather than a single sharp minimum, the probability of landing on the exact same point twice is negligible, resulting in unique novel creations.
+With this intuition in mind we examine two histograms, see Fig.
+(3).
+First, for every element of $\mathsf{S}^{\prime}$, their distance to their closest nearest neighbor from $\mathsf{S}$ is shown by the olive histogram.
+Second, for every element of $\mathsf{S^{\prime}}$, their distance to their closest nearest neighbor in $\mathsf{S}^{\prime}$ (excluding itself) is shown by the gray histogram.
+Both histograms exhibit a clear bimodal shape, indicating that at least two groups of samples are present in $\mathsf{S}^{\prime}$.
+This bimodality makes it possible to select the memorized $\delta_{m}$ and spurious $\delta_{s}$ thresholds to separate the samples forming the two peaks in each histogram.
+The left peak of the olive histogram identifies memorized samples by their small distance to the training set $\mathsf{S}$, leaving spurious and generalized samples in the right peak.
+The gray histogram then distinguishes truly novel patterns: its left peak contains high-frequency duplicates (both memorized and spurious), while its right peak consists of unique generalized samples.
+This bimodality provides an empirical basis for defining our detection metrics below.
+Following Yoon et al.
+(2023), we define the memorization detection metric $\mathcal{M}$.
+A sample $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}$ is considered memorized if its distance to its nearest neighbor ${\mathbf{x}}_{1}$, from the training set $\mathsf{S}$, falls below a threshold $\delta_{m}\in\mathbb{R}$:
+$\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S})=\mathbb{I}\bigg(d(\hat{{\mathbf{x}}},{\mathbf{x}}_{1})\leq\delta_{m}\bigg),$ (9)
+where $\mathbb{I}$ represents the indicator function and $\delta_{m}$ is a threshold derived from the olive distance histograms, see Fig.
+(3).
+Second, a sample $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}$ is spurious if it is not a memorized state but is a duplicate of the synthetic set $\mathsf{S}^{\prime}$.
+This is detected by the metric $\mathcal{S}$, which identifies non-memorized samples that are close to their respective nearest neighbor ${\mathbf{x}}^{\prime}_{1}$ within the synthetic set $\mathsf{S}^{\prime}$:
+$\mathcal{S}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime})=\mathbb{I}\bigg(d(\hat{{\mathbf{x}}},{\mathbf{x}}^{\prime}_{1})\leq\delta_{s}\bigg)\land\neg\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S}),$ (10)
+where $\delta_{s}\in\mathbb{R}$ is a threshold value chosen based on the gray distance histogram. Finally, a sample $\hat{{\mathbf{x}}}$ is generalized if it is neither memorized nor spurious:
+$\mathcal{G}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime})=\neg\mathcal{M}(\hat{{\mathbf{x}}},\mathsf{S})\land\neg\mathcal{S}(\hat{{\mathbf{x}}},\mathsf{S},\mathsf{S}^{\prime}).$ (11)
 Figure 4: Fractions of memorized, spurious, and generalized samples in synthetic sets across training sizes and datasets.
-As the training data size $K$ increases, memorization decreases and the fraction of generalized samples steadily increases (see top row).
-The fraction of spurious patterns rises and decreases at the boundary between the memorization and generalization phases (see bottom row).
-Using our detection metrics, we computed the fractions of memorized, spurious, and generalized samples for various sizes of training data for different datasets, see Figure 5.
-These datasets include MNIST [64], FASHION-MNIST [65], CIFAR10 [63], and LSUN-CHURCH [66] scaled down to $64\times 64$ resolution using center-crop and down-scale.
-For each dataset, we trained DDPM-based diffusion models [23] for $M=38$ different training set sizes.
-For each model $\alpha=1,...,M$, trained on the training set $\mathsf{S}_{\alpha}$, a synthetic set $\mathsf{S}^{\prime}_{\alpha}$ was generated.
-To account for duplication, we ensured that each $\mathsf{S}^{\prime}_{\alpha}$ is four times the size of its corresponding $\mathsf{S}_{\alpha}$, and checked that further increase of the synthetic set size does not qualitatively change our results, see Figure 11 in Appendix D.
-Each sample $\hat{{\mathbf{x}}}\in\mathsf{S}^{\prime}_{\alpha}$ is classified as either memorized, spurious, or generalized using the above metrics (13), (14), and (15).
-The fractions of these three pattern types with respect to the size of $\mathsf{S}^{\prime}_{\alpha}$ are plotted in Figure 5.
-For each dataset, our smallest model was trained on the training set of $|\mathsf{S}_{1}|=2$ data points while the largest model was trained on the entire original training set $\mathsf{S}_{M}=\mathsf{S}$.
-Please refer to Appendix D for additional details on training, selection of data sizes, and detection metrics’ hyperparameters for each dataset.
-The results in Figure 5 clearly demonstrate the transition from memorization to generalization as the dataset size increases.
-Meanwhile, the collected samples also show distinct characteristics in each of the considered phases, see Figure 2.
-Specifically, in the small data regime, we see in Figure 5 that the diffusion model predominantly replicates the training data (see the memorized panel of Figure 2, or additional samples in Figure 15 of the Appendix).
-As the data size surpasses the memorization capacity, e.g., at $K=3931$ for CIFAR10 in Figure 5, we observe a critical transition where the memorization fraction declines and spurious patterns become prominent.
-Such patterns exhibit strong duplication in the synthetic set (see the spurious panel of Figure 2 and additional samples in Figure 16 of the Appendix).
-The emergence of such patterns aligns with the onset of generalization, as the model moves away from strictly reproducing the training set and starts generating novel combinations of learned features.
-Subsequent increase of the training set size leads to the decrease of the fractions of memorized and spurious samples, signaling the transition to the full generalization regime.
-At this stage, the model completely loses its replication ability and no duplicate samples are detected in either training or synthetic sets (see the generalized panel of Figure 2 and additional samples in Figure 17 of the Appendix).
-Overall, this particular trend is also observed as the size of the diffusion model’s neural network is varied, see the results in Figure 9 of Appendix D.
-6 The Basin of Attraction Across the Transition
-Figure 6: Average log-volume of the hyperspheric basin of attraction of different sample types as the training data size $K$ grows, in CIFAR10 [63], computed with up to 512 samples per type.
-When fewer than 512 samples are available, all points are used.
-The log-volume of training samples and the transition plot from Figure 5 are included to confirm the shrinking basins of attraction in diffusion models.
-Shaded regions indicate standard deviation of the log-volume values.
-Diffusion models provide us with an opportunity to define the notion of basins of attraction around generated samples to further strengthen their connection to associative memories.
-Specifically, as the training data size $K$ varies, the sizes of those basins for memorized, spurious, and generalized samples can exhibit interesting behaviors.
-This section studies these phenomena empirically.
-Inspired by [38, 67, 54], we perturb a pattern ${\mathbf{x}}_{0}$ through the forward diffusion process for $T$ steps, and attempt to reverse it back through the reverse diffusion process.
-The maximum or critical time $t_{c}$, where the reverse process can still recover the pattern ${\mathbf{x}}_{0}$ from its perturbation ${\mathbf{x}}_{t_{c}}$ within an error margin $\delta\in\mathbb{R}$, defines the basin of attraction for pattern ${\mathbf{x}}_{0}$.
-The radius of this basin $R({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}})=\lVert{\mathbf{x}}_{0}-{\mathbf{x}}_{t_{c}}\rVert$ is the Euclidean norm of the difference between ${\mathbf{x}}_{0}$ and its recoverable perturbation ${\mathbf{x}}_{t_{c}}$.
-This radius can be converted to volume assuming that the basin of attraction is a sphere of radius $R({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}})$.
-Please also refer to Appendix G for more details on the experimentation and computation of $t_{c}$ and the volume.
-The results for log-volume versus training size are shown in Figure 6.
-The basins of attraction for training data decrease as $K$ increases.
-This trend is expected, since packing larger amount of training data in a given volume of the configuration space results in a smaller volume per data point, similarly to the ideal gas in physics.
-Memorized samples follow a similar trend.
-These results suggest that as $K$ grows, the diffusion model becomes less effective at denoising the perturbed data points, and instead gets much better at synthesizing.
-The most significant finding is that each of the three sample types exhibits a distinct log-volume trend: memorized samples have the largest log-volume, generalized samples have the smallest log-volume (this is expected as they should be unlikely recoverable by definition, resulting in close to zero volume), and spurious samples fall in between the two.
-This finding suggests a clear separation in how different sample types behave as the model’s training data size progresses.
-Please refer to Figs.
-(18) and (19) in the Appendix for the results of average log-volume across the transition for the other datasets, which exhibit the same behavior as Figure 6.
-To further illustrate the shrinkage of basins of attraction, we visualize the degradation in the denoising process as the training data size $K$ increases, as shown in Figure 7.
-For each training image, we apply the forward diffusion process using a timestep $t\approx t_{c}$, chosen near the end of the memorization phase (see Figure 20 in the Appendix).
-The perturbed image is then denoised using the deterministic DDIM sampler [24], starting from that same timestep $t$, across diffusion models trained with varying $K$.
-When $K$ is small, the diffusion model behaves like a DenseAM, effectively denoising and recovering the original image.
-However, as $K$ increases, this retrieval ability weakens, and the model shifts toward generating novel samples.
-In the intermediate regime, before full generalization, the outputs of the denoising process visually resemble spurious patterns: incomplete or noise-like novel generations which indicate the onset of generalization.
-Please refer to Appendix G for more details on this visual demonstration.
-Figure 7: An illustration of the shrinkage of basins of attraction as the training data size increases.
-Two CIFAR10 [63] and two LSUN-CHURCH [66] training images are independently perturbed using different timesteps $t\approx t_{c}$ (shown on the right), each chosen near the end of the memorization phase (see Figure 20 in the Appendix), and then denoised using the deterministic DDIM sampler [24].
-At small training data sizes, the model accurately recovers the original image.
-As the training data size increases, denoising performance deteriorates: the model shifts from retrieval of data points to generating novel outputs.
-In intermediate regimes, the outputs resemble spurious patterns: partial and noise-like generations indicative of this transition.
-7 The Relative Energy Gap Across the Transition
-Figure 8: Average relative energy gap of each sample type and the training samples in CIFAR10 [63] as training data size $K$ grows.
+As the training data size $K$ increases, memorization decreases and the fraction of generalized samples steadily increases, see top row of Panel (A).
+The fraction of spurious patterns rises and decreases at the boundary between the memorization and generalization phases, see bottom row of Panel (A).
+Panel (B) illustrates the average log-volume of the basins of attraction for memorized and spurious samples being statistically larger than that of generalized samples across all datasets.
+The shaded regions indicate standard deviation of the log-volume.
+[…]
+Furthermore, motivated by Biroli et al.
+(2024); Raya and Ambrogioni (2024); Li and Chen (2024), for these three distinct states we measure the critical time $t_{c}$, where the reverse process (2) can still recover a target pattern $\hat{{\mathbf{x}}}$ from its perturbation ${\mathbf{x}}_{t_{c}}$ within an error margin.
+The time $t_{c}$ helps define the radius $R(\hat{{\mathbf{x}}},{\mathbf{x}}_{t_{c}})=\lVert\hat{{\mathbf{x}}}-{\mathbf{x}}_{t_{c}}\rVert$ of the basin of the attractions of a sample $\hat{{\mathbf{x}}}$, which is the Euclidean norm of the difference between $\hat{{\mathbf{x}}}$ and its recoverable perturbation ${\mathbf{x}}_{t_{c}}$.
+The volume of this basin is simply defined by the $N$-ball volume:
+$V(\hat{{\mathbf{x}}},{\mathbf{x}}_{t_{c}})=\frac{\pi^{\frac{N}{2}}}{\Gamma(\frac{N}{2}+1)}R(\hat{{\mathbf{x}}},{\mathbf{x}}_{t_{c}})^{N},$ (12)
+where $\Gamma(\cdot)$ is the Gamma function. For numerical stability, we show the results of the average logarithmic of Eq. (34) in Fig. (4B) computed from samples belonging to each distinct state.
+Confirming the phenomenon observed in our toy model and core intuition, our experiments on these datasets demonstrate a clear three-phase transition from memorization to generalization.
+As shown in Fig.
+(4A), models trained in the small data regime predominantly replicate training samples.
+As the training data size $K$ surpasses a critical capacity, memorization sharply declines, and spurious samples emerge, signaling the onset of generalization.
+Further increase of $K$ diminishes the fraction of both memorized and spurious samples, and the model enters the full generalization regime, frequently producing unique novel patterns, see Fig.
+(1B) for the visualization of these three distinct patterns.
+Additionally, our results reveal a distinct hierarchy in the basins of attraction for the three sample types.
+Memorized samples consistently exhibit the largest log-volume, followed by spurious samples, while generalized samples have a much smaller volume, consistent with theoretical expectations.
+Please refer to Supplementary for more details of this section and additional results.
+Figure 5: Panel (A) depicts the average singular values of the energy curvature for different sample types and datasets and the shaded region is the standard deviation of the singular values (a particular training set size $K$ is shown).
+Memorized and spurious samples generally exhibit higher energy curvature, characterized by fewer near-zero singular values than generalized samples.
+Bottom row of (A) illustrates the average spectra computed for the training data points, where the changing color scheme denotes small to large data sizes, suggesting a drop in curvature (indicated by the decrease in singular values) as $K$ increases.
+Panel (B) shows candidate examples of memorized, generalized, and spurious samples from a Stable Diffusion model Rombach et al.
+(2022) trained on the LAION dataset Schuhmann et al.
+(2022), each corresponding to a distinct curvature signature.
+The candidate memorized sample has much larger singular values, while the candidate generalized sample has much smaller ones.
+The selected possible spurious samples have larger singular values than the candidate generalized sample, but smaller than the candidate memorized sample.
+The common trait between candidate memorized and spurious samples is that both are stable attractors, as demonstrated by their repeated similar generations given different initial points (or noise vectors) and the conditioning on text-prompt.
+The y-axis is clipped at the value of 1500 to better contrast the shown examples’ spectra.
+4.2  Energy Curvature
+While the changing fractions of memorized, spurious, and generalized samples constitute a good signature of the memorization-to-generalization transition, the underlying mechanism is also rooted in the geometry of the model’s learned energy landscape, as shown in Figs.
+(1A) and (2).
+To elucidate this mechanism, we quantify the energy curvature at the minima corresponding to each sample type obtained at a small time $t_{0}$.
+To measure this curvature, we rely on the relationship $\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})=-\nabla_{{\mathbf{x}}_{t}}E_{t}({\mathbf{x}}_{t})$ to analyze the Hessian of the energy function or the Jacobian of the score, evaluated at the location of each generated sample $\hat{{\mathbf{x}}}\in\mathbb{R}^{N}$.
+Specifically, following Stanczuk et al.
+(2022); Achilli et al.
+(2024); Ventura et al.
+(2024), we numerically approximate the spectra of the energy Hessian by exploiting the tangent space of the model’s learned manifold at a given pattern $\hat{{\mathbf{x}}}$.
+First, we generate a set of $4N$ perturbations around $\hat{{\mathbf{x}}}$ using a small time $t_{0}$ in Eq.
+(1).
+The learned score $s_{\theta}(\hat{{\mathbf{x}}}_{t_{0}},t_{0})$ is then evaluated at each perturbed point $\hat{{\mathbf{x}}}_{t_{0}}$ and the resulting $4N$ score vectors form a matrix $\mathsf{M}\in\mathbb{R}^{N\times 4N}$.
+Singular value decomposition is then performed on $\mathsf{M}$ to extract its singular values, serving as our measure of the local curvature of $\hat{{\mathbf{x}}}$ in all directions.
+The curvature spectra for the three sample types for different datasets are illustrated in the top row of Fig.
+(5A).
+Please refer to the Supplementary for further experimental details.
+While a consistent separation in the curvature spectra of memorized, spurious, and generalized samples is not readily apparent for all training data sizes $K$, likely due to the geometric complexity of the transitional phases, we are still able to observe very well the separation of the curvature spectra between memorized and generalized samples at many sizes of the training dataset (prominently at or after the peak of spurious states).
+In addition, we repeated a similar analysis for training data points.
+There is a consistent decrease in the singular values measured at the locations of the training data points as their number $K$ increases, see bottom row of Fig.
+(5A).
+This provides a direct geometric evidence of the mechanism behind generalization: the model transitions from creating sharp, high-curvature minima for individual data points in the memorization regime to learning a smoother, lower-curvature manifold that encompasses the entire dataset in the generalization regime.
+To validate these geometric findings on a large-scale model, we performed a case study using Stable Diffusion Rombach et al.
+(2022), trained on the LAION dataset Schuhmann et al.
+(2022), shown in Fig.
+(5B).
+The analysis of handpicked candidate examples of memorized, spurious, and generalized samples from LAION reveals a similar hierarchy of energy curvatures.
+As expected, a heavily memorized sample occupies a high-curvature minimum, evidenced by large singular values.
+In contrast, a more generalized sample exhibits a significantly lower-curvature signature, while spurious samples occupy an intermediate geometric space.
+Notably, the spectra for all generated samples remain relatively sharp, with few near-zero singular values, as similarly found in Jeon et al.
+(2024).
+We argue this is a consequence of text conditioning, which constrains the generative manifold, making Stable Diffusion more prone to operate in high-curvature modes, increasing the risk of memorization.
+4.3  Relative Energy Across the Transition
+Figure 6: Average relative energy gap of each sample type and the training samples in CIFAR10 Krizhevsky et al.
+(2014) as training data size $K$ grows.
 The gaps are measured relative to training samples, whose energy is set to zero (energy is defined up to an additive constant).
 Shaded regions show standard deviation of the gap values.
-Lastly, we study how the energy of the three sample types changes as the training data size increases.
-Following the efficient relative energy computation from [38], we measure the relative energy gap of the three sample types with respect to that of the training data samples.
-Please refer to Appendix H for full details on the experiment and computation of the relative energy gap for different sample types.
-The results are recorded in Figure 8.
+Alongside our energy curvature analyses, we study how the energy of these three distinct sample types changes as the training data size increases.
+Following the efficient relative energy computation from Raya and Ambrogioni (2024), we measure the relative energy gap of the three sample types with respect to that of the training data samples.
+The results are recorded in Figure˜6.
+Please refer to Appendix˜E for full details on the experiment and computation of the relative energy gap for different sample types, alongside additional results.
 For small training data sizes, the energy of the memorized samples is low and of the generalized samples is high.
 As $K$ increases, the energy of the memorized samples increases, while the energy of the generalized samples decreases.
-The energy of the spurious states is the lowest among these three types of sample at the memorization-generalization transition.
-Please refer to Figure 23 in the Appendix for the results of the average relative energy gap across the transition for the other datasets, which show a similar behavior as Figure 8.
+The energy of the spurious states is the lowest among these three distinct sample types during the memorization-generalization transition.
 5  Conclusion
-In this work, we established a novel connection between diffusion models and DenseAMs.
-Specifically, we identified the emergence of creativity (generalization) in diffusion models with the failure of successful memory recalls, via an energy-based theoretical description of diffusion models.
-Using this theory, we identified a novel phase – spurious states – previously overlooked in the memorization-generalization literature.
-Finally, we established that diffusion models trained on natural datasets follow the same transition (memorization to spurious phase to generalization) as predicted by our theory.
-The peak frequency of spurious states among generated samples marks the onset of generalization as the training set size increases.
-The volume measurements and relative energy gap analysis of different sample types, across multiple datasets consistently support this three distinct phases of computation, providing a comprehensive understanding of how diffusion models evolve from memorization to generalization as the amount of training data increases.
-Aknowledgments
-M.N. acknowledges the support of PNRR MUR project PE0000013-FAIR.
-B.P. acknowledges the support of IBM through the IBM-Rensselaer Future of Computing Research Collaboration (FCRC).
-Appendix A Additional Details on the Energy Formulation of the 2D Toy Example
-In order to obtain Equation 11, it is easiest to introduce polar coordinates for both the state vector ${\mathbf{x}}$ and the training data ${\mathbf{y}}$:
+In conclusion, we established a novel connection between DMs and DenseAMs by identifying the emergence of creativity (generalization) in DMs with the failure of successful memory recalls via an energy-based associative memory.
+Using this theory, we empirically identified a novel intermediate phase, spurious states, previously overlooked in related DM literature, which marks the onset of generalization in these models as the training data size increases.
+Additional analyses of the volume of the basins of attraction on samples generated from DMs and various aspects about their energy help validate our hypothesis delineated in Fig.
+(1).
+Our findings on Stable Diffusion suggest that these novel, stable attractors persist even in large models trained on massive datasets.
+Thus, our framework provides a new tool for understanding the computational mechanisms of DMs, opening new avenues for characterizing their behaviors.
+Acknowledgments
+The results presented here were obtained while Dmitry Krotov was employed by IBM Research.
+At the time of the submission Dmitry Krotov is no longer employed by IBM Research.
+Matteo Negri acknowledges the support of PNRR MUR project PE0000013-FAIR.
+Bao Pham and Mohammed Zaki acknowledge the support of RPI-IBM Future of Computing Research Collaboration program (FCRC).
+Appendix
+Appendix A Toy Example
+A.1  Overview
+For the 2D toy example, our training data points lie on a unit circle.
+In the case of infinite training data size or $K\rightarrow\infty$, the distribution of these data points can be described by a continuous density of states:
+$p({\mathbf{y}})=\frac{1}{\pi}\delta\big(y_{1}^{2}+y_{2}^{2}-1\big).$ (13)
+The probability of the generated data is proportional (up to terms independent of the state $\mathbf{x}$}) to
+$p({\mathbf{x}})\sim\int\limits_{-\infty}^{+\infty}dy_{1}dy_{2}\ p({\mathbf{y}})\ e^{-\beta\lVert{\mathbf{x}}-{\mathbf{y}}\rVert^{2}_{2}}=e^{-\beta(R^{2}+1)}I_{0}(2\beta R),$ (14)
+where $I_{0}(\cdot)$ is a modified Bessel function of the first kind. Thus, the energy of the model is given by
+$\begin{split}E^{\text{AM}}(R,\phi)&=R^{2}+1-\frac{1}{\beta}\log\big[I_{0}(2\beta R)\big]\underset{\beta\rightarrow\infty}{\approx}(R-1)^{2},\end{split}$ (15)
+where $R$ is the radius of the unit circle and $\phi$ is the polar angle.
+The dependence on the polar angle $\phi$ in Eq.
+(15) completely disappears from the final result, since the local minima of the resulting energy function form a continuous manifold, described by a parabola centered around $R=1$ as $\beta\rightarrow\infty$.
+This behavior describes the fully generalized phase.
+In order to obtain Eq. (14), it is easiest to introduce polar coordinates for both the state vector $\mathbf{x}$} and the training data $\mathbf{y}$}:
 $\begin{cases}x_{1}=R\cos(\phi)\\ x_{2}=R\sin(\phi)\end{cases}\ \ \ \ \ \ \ \ \ \ \begin{cases}y_{1}=r\cos(\varphi)\\ y_{2}=r\sin(\varphi)\end{cases}$
 The integral (14) can then be written as
-[…]
+$p({\mathbf{x}})\sim\int\limits_{0}^{2\pi}d\varphi\int\limits_{0}^{\infty}rdr\frac{1}{\pi}\delta(r^{2}-1)e^{-\beta[R^{2}+r^{2}-2Rr\cos(\varphi-\phi)]}=e^{-\beta(R^{2}+1)}I_{0}(2\beta R)$
 and explicitly computed using the definition of the modified Bessel functions [19].
-Appendix B Experimental Details of the 2D Toy Example
-This section describes the two-dimensional toy example used in Section 3, including details about the neural network architecture, training, generation, and computation of the model’s energy.
-All of the corresponding code will be made publicly available.
+A.2  Experimental Details
 Dataset.
-To construct the dataset, we uniformly sample points on the unit circle.
-Concretely, we take ${\mathbf{y}}\sim p({\mathbf{y}})=\frac{1}{2\pi}\delta(r-1)$, where $r=\sqrt{y_{1}^{2}+y_{2}^{2}}$ is the radius and $\delta(r-1)$ is the delta function ensures that all probability mass lies on radius $r=1$.
-The factor $\frac{1}{2\pi}$ guarantees that the angles are sampled uniformly, relying on the identity $\delta(\sqrt{y_{1}^{2}+y_{2}^{2}}-1)=2\delta(y_{1}^{2}+y_{2}^{2}-1)$ in $p({\mathbf{y}})=\frac{1}{\pi}\delta(y^{2}_{1}+y_{2}^{2}-1)$.
-In practice, we sample polar coordinates $(r,\varphi)$ with $r=1$ and uniformly sample the angular coordinate $\varphi$ from $[0,2\pi]$.
-We then convert $(r,\varphi)$ to Cartesian coordinates ($y_{1},y_{2}$) using $y_{1}=r\cos\varphi$ and $y_{2}=r\sin\varphi$.
-We first created a data set of $60000$ points using a fixed random seed, then built smaller subsets (e.g., $K\in\{2,4,9,1000,10000\}$) by progressively adding distinct samples without replacement, ensuring that each subset is a strict subset of the next.
-This approach allows us to systematically examine how the model behaves, as we vary the size of the training set.
+To construct the dataset, we uniformly sampled points on the unit circle.
+Concretely, we had ${\mathbf{y}}\sim p({\mathbf{y}})=\frac{1}{2\pi}\delta(r-1)$, where $r=\sqrt{y_{1}^{2}+y_{2}^{2}}$ is the radius and $\delta(r-1)$ is the delta function ensures that all probability mass lies on radius $r=1$.
+The factor $\frac{1}{2\pi}$ guarantees that the angles are sampled uniformly, relying on the identity:
+$\delta(\sqrt{y_{1}^{2}+y_{2}^{2}}-1)=2\delta(y_{1}^{2}+y_{2}^{2}-1),$ (16)
+where $p({\mathbf{y}})=\frac{1}{\pi}\delta(y^{2}_{1}+y_{2}^{2}-1)$.
+In practice, we sampled polar coordinates $(r,\varphi)$ with $r=1$ and uniformly sampled the angular coordinate $\varphi\in[0,2\pi]$.
+We then converted $(r,\varphi)$ to Cartesian coordinates ($y_{1},y_{2}$) using $y_{1}=r\cos\varphi$ and $y_{2}=r\sin\varphi$.
+Then, We created a data set of $60000$ points using a fixed random seed, then built smaller subsets by progressively adding distinct samples without replacement, ensuring that each subset is a strict subset of the next.
+This approach allowed us to systematically examine how the model behaves as we vary the size of the training set.
 Exact Score.
-Following Equation 12 specified in Section 3, energy can be written in closed form, whose local minima form the continuous data manifold at $R=1$ (throughout this paper we use $r$ for the polar radius of the training data and $R$ for the polar radius of the generated samples).
+Following Eq.
+(15), the energy can be written in closed form, whose local minima form the continuous data manifold at R=1R=1111Please note throughout this paper we use $r$ for the polar radius of the training data and $R$ for the polar radius of the generated samples..
 Furthermore, in the case of many data points or $K\rightarrow\infty$, we have $p({\mathbf{y}})=\frac{1}{\pi}\delta(y_{1}^{2}+y_{2}^{2}-1)$.
 Since $p(R)\propto\exp^{-E^{\text{AM}}(R)}$, the score can be described as
-$\begin{split}-\nabla_{R}\log p(R)&=\nabla_{R}E^{\text{AM}}(R)\\ &=2R-\frac{2\beta I_{1}(2\beta R)}{\beta I_{0}(2\beta R)}\\ &=2R-2\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)}\end{split}$ (16)
-where $E^{\text{AM}}(R)$ is Equation 12, and $I_{1}$ and $I_{0}$ are the first-order and zero-order modified Bessel functions of the first kind, respectively. With further expansion, the score becomes
-$\nabla_{R}\log p(R)=2\left(\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)}-R\right)\qquad\Rightarrow\qquad\nabla_{x}\log p(x)=2\left(\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)}-R\right)\frac{x}{R}$ (17)
-We implemented Equation 12 on a $20\times 20$ grid to obtain the exact energy and score of the model, overlaying 1000 samples from the training set.
+$\begin{split}-\nabla_{R}\log p(R)&=\nabla_{R}E^{\text{AM}}(R)\\ &=2R-\frac{2\beta I_{1}(2\beta R)}{\beta I_{0}(2\beta R)}\\ &=2R-2\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)},\end{split}$ (17)
+where $E^{\text{AM}}(R)$ is Eq. (15), and $I_{1}$ and $I_{0}$ are the first-order and zero-order modified Bessel functions of the first kind, respectively. With further expansion, the score becomes
+$\nabla_{R}\log p(R)=2\left(\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)}-R\right)\quad\Rightarrow\quad\nabla_{x}\log p(x)=2\left(\frac{I_{1}(2\beta R)}{I_{0}(2\beta R)}-R\right)\frac{x}{R}.$ (18)
+We implemented Eq.
+(15) on a $20\times 20$ grid to obtain the exact energy and score of the model, overlaying 1000 samples from the training set.
 For visualization purposes, we set $\beta=20$, ensuring a well-defined energy landscape.
 Additionally, we normalized the energy by subtracting its minimum value to ensure that the lowest energy point remains at zero.
-VE-SDE. To align with the theoretical framework, we employ a variance exploding (VE) SDE of the form
+To align with the theoretical framework, we employed the variance-exploding (VE) SDE of the form:
 $\mathrm{d}{\mathbf{x}}_{t}=\sigma\,\mathrm{d}\mathbf{w}_{t}$ (19)
 where the diffusion coefficient $g(t)=\sigma$.
-Assuming $t\in(0,1]$, the variance of the diffusion kernel is given by $\int_{0}^{t}g^{2}(s)\,\mathrm{d}s=\sigma^{2}t$, which allows us to construct the corresponding Gaussian kernel
-$p({\mathbf{x}}_{t}|{\mathbf{y}})=\mathcal{N}({\mathbf{x}}_{t};{\mathbf{y}},\sigma^{2}t\mathbf{I})$
+Assuming $t\in(0,1]$, the variance of the diffusion kernel is given by $\int_{0}^{t}g^{2}(s)\,\mathrm{d}s=\sigma^{2}t$, enabling us to construct the corresponding Gaussian kernel:
+$p({\mathbf{x}}_{t}|{\mathbf{y}})=\mathcal{N}({\mathbf{x}}_{t};{\mathbf{y}},\sigma^{2}tI)$
 However, in practice we used $t\in[\epsilon,1]$ with $\epsilon=10^{-5}$ for numerical stability. The generative dynamics are given by
@@
 $\frac{\mathrm{d}\mathbf{x}_{t}}{\mathrm{d}t}=-\frac{1}{2}\sigma^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ (21)
-which shares the same marginal distributions as the SDE in Equation 19 and is useful for likelihood estimation.
+which shares the same marginal distributions as the SDE (20) and is useful for likelihood estimation.
+Empirical distribution. We represented the data as an empirical distribution
+$p(\mathbf{y})=\frac{1}{K}\sum_{\mu=1}^{K}\delta^{(N)}(\mathbf{y}-\bm{\xi}^{\mu})$ (22)
+where $N$ is the dimensionality of the data point $\mathbf{y}$}.
+Since ${\mathbf{x}}_{t}$ is drawn from the forward process distribution $p({\mathbf{x}}_{t}|{\mathbf{y}})$ that is conditioned on the data point $\mathbf{y}$}, which can be expressed as a Gaussian kernel $\mathcal{N}({\mathbf{x}}_{t};\mathbf{y},\sigma^{2}tI)$.
+We obtained the distribution $p(\mathbf{x}_{t},t)$ as
+$\begin{split}p(\mathbf{x}_{t},t)&=\int p(\mathbf{x}_{t}\mid\mathbf{y})\,p(\mathbf{y})\,\mathrm{d}\mathbf{y}\\ &=\frac{1}{K}\sum_{\mu=1}^{K}\int\mathcal{N}\!\bigl(\mathbf{x}_{t};\mathbf{y},\sigma^{2}tI\bigr)\,\delta^{(N)}\!\bigl(\mathbf{y}-\bm{\xi}^{\mu}\bigr)\,\mathrm{d}\mathbf{y}\\ &=\frac{1}{K}\sum_{\mu=1}^{K}\mathcal{N}\!\bigl(\mathbf{x}_{t};\,\bm{\xi}^{\mu},\,\sigma^{2}t\,I\bigr)\\ &=\frac{1}{K}\sum_{\mu=1}^{K}\frac{1}{(2\pi\,\sigma^{2}t)^{\frac{N}{2}}}\exp\!\Bigl(-\,\frac{\|\mathbf{x}_{t}-\bm{\xi}^{\mu}\|_{2}^{2}}{2\,\sigma^{2}t}\Bigr)\end{split}$ (23)
+which is the derived $E^{\text{DM}}$ in the main text, when the variance of the Gaussian kernel is $\sigma^{2}t$.
+Energy Computation.
+Following [62] we computed the log-likelihood $\log p_{\theta}({\mathbf{x}}_{0})$ given by a diffusion model, with the instantaneous change of variable formula [13], where ${\mathbf{x}}_{0}$ are the generated samples from the model.
+Replacing Eq.
+(20) into the log-likelihood equation,
+$\log p_{0}({\mathbf{x}}_{0};\mathbf{\theta})=\log p_{T}({\mathbf{x}}_{T})+\int_{0}^{T}\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)\mathrm{d}t$ (24)
+with $\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)=-\frac{1}{2}\sigma^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$.
+The function $\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ comes from the above Probability Flow ODE (21), and $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ denotes the divergence of the function (or the trace of its Jacobian).
+To estimate the likelihood of the model, we integrated $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$, from a small time ${\epsilon}=10^{-5}$ to $T=1$, using a numerical integrator and added the prior logarithmic likelihood to it following Eq.
+(24).
+The divergence term $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ is computed using the Laplacian, which is computationally feasible for this toy example, instead of using the Hutchinson trace estimator [28] done in [62].
+We use the RK45 method [18] implemented in the ODE solver scipy.integrate.solve_ivp from Scipy [68] to solve the above integral.
+To compute the energy, we used the equation of $E^{\text{DM}}$ in the main text, which can be also computed using Eq.
+(24).
+From the Boltzmann distribution $p({\mathbf{x}})=\frac{1}{Z}\exp\bigl[-\beta E(\mathbf{x})\bigl]$, we obtained $\log p({\mathbf{x}})\propto-\beta E(\mathbf{x})$.
+In our case, the inverse temperature is $\beta=\frac{1}{2\sigma^{2}t}$.
+The equation for our VE-SDE thus becomes
+$E^{\text{DM}}({\mathbf{x}}_{t},t)=-2\sigma^{2}t\log p({\mathbf{x}}_{t},t)$ (25)
+To visualize the energy landscape, we observed the energy at $t=0.15$ which corresponds to $\beta=3.3$.
 Model and Training Details.
-For each training size $K$ in Figure 3, we use a U-Net inspired MLP-based network to estimate the score $s_{\theta}({\mathbf{x}}_{t},t)$, without using any down- or up- sampling blocks.
+For each training data size $K$, we used an MLP-based network to estimate the score $s_{\theta}({\mathbf{x}}_{t},t)$.
 The general architecture includes (1) a Fourier random feature (FRF) timestep embedding layer and a linear layer which projects the concatenation of FRF timestep embedding and input into a latent dimension of 256; (2) an encoder which consists of four non-convolutional residual blocks (using the same latent dimension) with Swish [47] activation in between; (3) a decoder which consists of the same number of residual blocks; and (4) a linear block which projects the latent variable back to the 2D space.
-The network was trained in continuous time (with 1000 discretized steps) with $\sigma=1$ using the objective function from [23, 25].
+The network was trained in continuous time with $\sigma=1$ using the objective function from [22, 62].
 Optimization is performed using the Adam [34] optimizer with a learning rate $lr=10^{-4}$.
 The batch size is set to $\textbf{min}(K,500)$.
-All models are trained for $800000$ iterations with the maximum batch size of $500$.
-Meanwhile, we use Euler-Maruyama discretization method [71] to solve Equation 19 of the reverse SDE for the generation of samples.
-Empirical distribution. We represent the data as an empirical distribution
-$p(\mathbf{y})=\frac{1}{K}\sum_{\mu=1}^{K}\delta^{(N)}(\mathbf{y}-\bm{\xi}^{\mu})$ (21)
-where $N$ is the dimensionality of the data point ${\mathbf{y}}$.
-Since ${\mathbf{x}}_{t}$ is drawn from the forward process distribution $p({\mathbf{x}}_{t}|{\mathbf{y}})$ that is conditioned on the data point ${\mathbf{y}}$, which can be expressed as a Gaussian kernel $\mathcal{N}({\mathbf{x}}_{t};\mathbf{y},\sigma^{2}t\mathbf{I})$.
-We can obtain the distribution $p(\mathbf{x}_{t},t)$ as
-$\begin{split}p(\mathbf{x}_{t},t)=\int p(\mathbf{x}_{t}\mid\mathbf{y})\,p(\mathbf{y})\,\mathrm{d}\mathbf{y}&=\frac{1}{K}\sum_{\mu=1}^{K}\int\mathcal{N}\!\bigl{(}\mathbf{x}_{t};\mathbf{y},\sigma^{2}t\mathbf{I}\bigr{)}\,\delta^{(N)}\!\bigl{(}\mathbf{y}-\bm{\xi}^{\mu}\bigr{)}\,\mathrm{d}\mathbf{y}\\ &=\frac{1}{K}\sum_{\mu=1}^{K}\mathcal{N}\!\bigl{(}\mathbf{x}_{t};\,\bm{\xi}^{\mu},\,\sigma^{2}t\,\mathbf{I}\bigr{)}\\ &=\frac{1}{K}\sum_{\mu=1}^{K}\frac{1}{(2\pi\,\sigma^{2}t)^{\frac{N}{2}}}\exp\!\Bigl{(}-\,\frac{\|\mathbf{x}_{t}-\bm{\xi}^{\mu}\|_{2}^{2}}{2\,\sigma^{2}t}\Bigr{)}\end{split}$ (22)
-which is Equation 4, when the variance of the Gaussian kernel is $\sigma^{2}t$.
-Likelihood Computation.
-Following [25] we can compute the log-likelihood $\log p_{\theta}({\mathbf{x}}_{0})$ given by a diffusion model, with the instantaneous change of variable formula [72], where ${\mathbf{x}}_{0}$ are the generated samples from the model.
-Replacing Equation 19 into the log-likelihood equation,
-$\log p_{0}({\mathbf{x}}_{0};\mathbf{\theta})=\log p_{T}({\mathbf{x}}_{T})+\int_{0}^{T}\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)\mathrm{d}t,\qquad\text{ with }\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)=-\frac{1}{2}\sigma^{2}\nabla_{\mathbf{x}_{t}}\log p_{t}(\mathbf{x}_{t})$ (23)
-The function $\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ came from the above Probability Flow ODE (Equation 20), and $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ denotes the divergence of the function (or the trace of its Jacobian).
-To estimate the likelihood of the model, we integrate $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ from a small time ${\epsilon}=10^{-5}$ to $T=1$ using a numerical integrator and add the prior logarithmic likelihood to it following Equation 23.
-The divergence term $\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)$ is computed using the Laplacian, which is computationally feasible for this toy example, instead of using the Hutchinson trace estimator [73] done in [25].
-We use the RK45 method [74] implemented in the ODE solver scipy.integrate.solve_ivp from Scipy [75] to solve the above integral.
-Energy Computation.
-To compute the energy, we use Equation 5, which can be also computed using Equation 23.
-From the Boltzmann distribution $p({\mathbf{x}})=\frac{1}{Z}\exp[-\beta E(\mathbf{x})\bigl{]}$, we obtained $\log p({\mathbf{x}})\propto-\beta E(\mathbf{x})$.
-In our case the inverse temperature is $\beta=\frac{1}{2\sigma^{2}t}$.
-The equation for our VE-SDE thus becomes
-$E^{\text{DM}}({\mathbf{x}}_{t},t)=-2\sigma^{2}t\log p({\mathbf{x}}_{t},t)$ (24)
-To visualize the energy landscape, we observe the energy at $t=0.15$ which corresponds to $\beta=3.3$.
-Sample Clustering. To cluster the generated samples, we use the AgglomerativeClustering algorithm from scikit-learn, which is a part of SciPy [75].
-Appendix C Additional Details on the Likelihood of Diffusion Models
-For concreteness, we further explain the likelihood computation in Equation 23 within this section. Specifically, we want to highlight the formulation done by [25, 72] which resulted in Equation 23.
-Probability Flow ODE. Suppose we have the following forward process
+All models were trained for $800000$ iterations with the maximum batch size of $500$.
+Meanwhile, we used Euler-Maruyama discretization method [35], with 1000 discretized steps, to solve the reverse SDE (20) to generate samples.
+Sample Clustering. To cluster the generated samples, we used the AgglomerativeClustering algorithm from scikit-learn, a part of SciPy [68].
+A.3  Log-Likelihood of Diffusion Models
+For concreteness, we provide the likelihood computation of Eq. (24) within this section to highlight the formulation of Eq. (24) done by [62, 13].
+Probability Flow ODE. Suppose we have the following forward process:
 $\mathrm{d}{\mathbf{x}}_{t}=\mathbf{f}({\mathbf{x}}_{t},t)\mathrm{d}t+\mathbf{g}({\mathbf{x}}_{t},t)\mathrm{d}\mathbf{w}_{t}$ (26)
@@
 $\frac{\partial p_{t}({\mathbf{x}}_{t})}{\partial t}=-\sum^{N}_{i=1}\frac{\partial}{\partial x_{i}}\big[f_{i}({\mathbf{x}}_{t},t)p_{t}({\mathbf{x}}_{t})\big]+\frac{1}{2}\sum^{N}_{i=1}\sum^{N}_{j=1}\frac{\partial^{2}}{\partial x_{i}\partial x_{j}}\bigg[\sum^{N}_{k=1}g_{ij}({\mathbf{x}}_{t},t)g_{jk}({\mathbf{x}}_{t},t)p_{t}({\mathbf{x}}_{t})\bigg]$ (27)
-which corresponds to Fokker-Planck equation [76]:
+which corresponds to the Fokker-Planck equation [46]:
 $\begin{split}\frac{\partial p_{t}({\mathbf{x}}_{t})}{\partial t}&=-\sum^{N}_{i=1}\frac{\partial}{\partial x_{i}}\big[f_{i}({\mathbf{x}}_{t},t)p_{t}({\mathbf{x}}_{t})\big]+\frac{1}{2}\sum^{N}_{i=1}\frac{\partial}{\partial x_{i}}\bigg[\sum^{N}_{j=1}\frac{\partial}{\partial x_{j}}\bigg[\sum^{N}_{k=1}g_{ij}({\mathbf{x}}_{t},t)g_{jk}({\mathbf{x}}_{t},t)p_{t}({\mathbf{x}}_{t})\bigg]\bigg]\\ &=-\sum^{N}_{i=1}\frac{\partial}{\partial x_{i}}\big[f_{i}({\mathbf{x}}_{t},t)p_{t}({\mathbf{x}}_{t})\big]\\ &\quad\,\,+\frac{1}{2}\sum^{N}_{i=1}\frac{\partial}{\partial x_{i}}\bigg[p_{t}({\mathbf{x}}_{t})\,\nabla\cdot\big[\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big]+p_{t}({\mathbf{x}}_{t})\big[\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big]\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg]\\ &=-\sum^{N}_{i=1}\frac{\partial}{\partial x_{i}}\bigg\{f_{i}({\mathbf{x}}_{t},t)p_{t}({\mathbf{x}}_{t})\\ &\quad\,\,-\frac{1}{2}\bigg[\nabla\cdot\big[\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big]+\big[\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big]\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg]p_{t}({\mathbf{x}}_{t})\bigg\}\\ &=-\sum^{N}_{i=1}\frac{\partial}{\partial x_{i}}\big[\tilde{f}_{i}({\mathbf{x}}_{t},t)\big]p_{t}({\mathbf{x}}_{t})\end{split}$ (28)
 where $\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)={\mathbf{f}}({\mathbf{x}}_{t},t)-\frac{1}{2}\nabla\cdot\big[\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big]-\frac{1}{2}\big[\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big]\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$.
-With careful inspection of Equation 27, it is equal to the Liouville equation if the diffusion term $\tilde{\mathbf{g}}({\mathbf{x}},t)=0$ and essentially, it is the probability flow ODE where
-$\begin{split}\mathrm{d}{\mathbf{x}}_{t}&=\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)\mathrm{d}t+\tilde{\mathbf{g}}({\mathbf{x}}_{t},t)\mathrm{d}{\mathbf{w}}_{t}\\ &=\bigg{\{}{\mathbf{f}}({\mathbf{x}}_{t},t)-\frac{1}{2}\nabla\cdot\big{[}\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big{]}-\frac{1}{2}\big{[}\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big{]}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg{\}}\mathrm{d}t\end{split}$ (28)
-Using Equation 28, we can derive an appropriate probability flow ODE from the forward process (1) introduced in the main text. For example, we have the following equation
+With careful inspection of Eq.
+(28), it is equal to the Liouville equation if the diffusion term $\tilde{\mathbf{g}}({\mathbf{x}},t)=0$ and essentially, it is the probability flow ODE where
+$\begin{split}\mathrm{d}{\mathbf{x}}_{t}&=\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)\mathrm{d}t+\underset{0}{\underbrace{\tilde{\mathbf{g}}({\mathbf{x}}_{t},t)\mathrm{d}{\mathbf{w}}_{t}}}\\ &=\bigg\{{\mathbf{f}}({\mathbf{x}}_{t},t)-\frac{1}{2}\nabla\cdot\big[\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big]-\frac{1}{2}\big[\mathbf{g}({\mathbf{x}}_{t},t)\mathbf{g}({\mathbf{x}}_{t},t)^{\top}\big]\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg\}\mathrm{d}t\end{split}$ (29)
+Using the ODE (29), we can derive an appropriate probability flow ODE from the forward process detailed in the main text. For example, we have the following equation
 $\mathrm{d}{\mathbf{x}}_{t}=\underset{:=\,\tilde{\mathbf{f}}({\mathbf{x}}_{t},t)}{\underbrace{\bigg\{\mathbf{f}({\mathbf{x}}_{t},t)-\frac{1}{2}g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg\}}}\mathrm{d}t$ (30)
 for the toy model where $\mathbf{f}({\mathbf{x}}_{t},t)=0$.
-Log-likelihood. Furthermore, if we first take the logarithm of Equation 27 where
+Log-likelihood. Furthermore, if we first take the logarithm of Eq. (28) where
 $\begin{split}\frac{\partial\log p_{t}({\mathbf{x}}_{t})}{\partial t}=\frac{1}{p_{t}({\mathbf{x}}_{t})}\frac{\partial p_{t}({\mathbf{x}}_{t})}{\partial t}=-\sum^{N}_{i=1}\frac{\partial\tilde{f}_{i}({\mathbf{x}}_{t},t)}{\partial x_{i}}=\nabla\cdot\tilde{{\mathbf{f}}}({\mathbf{x}}_{t},t)\end{split}$ (31)
@@
 $\nabla\cdot\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)=\mathbb{E}_{p(\epsilon)}\big[\epsilon^{\top}\nabla_{{\mathbf{x}}}\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)\epsilon\big]$ (33)
-where $\nabla_{{\mathbf{x}}}\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)$ denotes the Jacobian of $\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)$ and the random variable $\epsilon$ satisfies $\mathbb{E}_{p(\epsilon)}[\epsilon]=0$ and $\text{Cov}_{p(\epsilon)}[\epsilon]=\mathbf{I}$.
-However, for the toy model in Section 3, due to its low dimensionality we find that it is possible to not utilize the Hutchinson trace estimator at all, and instead opt to compute the trace of the Jacobian without using the noise estimators — to get the actual trace of the Hessian of the energy.
-Appendix D Experimental Details on the Transition Mapping
-Dataset Point A Point B Total
-(Data Size) (Data Size) (Data Size)
-MNIST 4,000 32,000 60,000
-CIFAR10 2,000 16,000 50,000
-LSUN-CHURCH 2,000 16,000 126,227
-FASHION-MNIST 4,000 16,000 60,000
-Table 1: Table showing the critical points $A$ and $B$ of the memorization-generalization transition for each dataset.
-We use 30 linearly spaced points between $A$ and $B$ to finely characterize the transition plots shown in Figure 5.
-Points Selection.
-For the computation of Figure 5, we follow the experiment setup of [45] and conduct a sparse search starting at data size $500$ and doubling it all the way to the total data size $|\mathsf{S}|$, i.e., $K=500,1000,2000,\dots,|\mathsf{S}|$.
-We then identified two transitional critical points, $A$ and $B$, see Table 1, to conduct a more thorough search to elucidate the memorization-generalization transition.
-Point $A$ indicates the initial drop in memorization while $B$ signals the plateauing of memorization.
+where $\nabla_{{\mathbf{x}}}\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)$ denotes the Jacobian of $\tilde{{\mathbf{f}}}_{\theta}({\mathbf{x}}_{t},t)$ and the random variable $\epsilon$ satisfies $\mathbb{E}_{p(\epsilon)}[\epsilon]=0$ and $\text{Cov}_{p(\epsilon)}[\epsilon]=I$.
+However, for the toy example, due to its low dimensionality we found that it is possible to not utilize the Hutchinson trace estimator at all, and instead opted to compute the trace of the Jacobian without using the noise estimators.
+Appendix B Memorization-to-Generalization Transition
+B.1  Selection of Points
+For the computation of fig.˜4 in the main text, we followed the experiment setup of [74] and conducted an initial sparse search of the transition, starting at data size $500$ and doubling it all the way to the total data size $|\mathsf{S}|$, i.e., $K\in\{500,1000,2000,\dots,|\mathsf{S}|\}$.
+We then identified two transitional critical points, $A$ and $B$, see Table (1), to conduct a more thorough search to elucidate the memorization-to-generalization transition.
+Here, point $A$ indicates the initial drop in memorization while $B$ signals the plateauing of memorization.
 To capture the finer details of this transition, we performed a search of $30$ inclusive linearly spaced points, from point $A$ to point $B$.
-For regions outside of the transition, using linear spacing, we sample $5$ points from $|\mathsf{S}|=2$ to point $A$, and another $5$ points from $B$ to the total dataset size $|\mathsf{S}|$, inclusively.
-In other words, we train a separate model for each selected data size and generate the corresponding evaluation and synthetic sets — keeping the same configuration for all models, see Table 3.
-Then, we computed the memorization, spurious, and generalization fractions using equations (13), (14), and (15).
+For regions outside of the transition, using linear spacing, we sampled $5$ points from $|\mathsf{S}|=2$ to point $A$, and another $5$ points from $B$ to the total dataset size $|\mathsf{S}|$, inclusively.
+In other words, we trained a separate model for each selected data size and generated the corresponding evaluation and synthetic sets while keeping the same configuration for all models, see Table (3).
+Then, we computed the memorization, spurious, and generalization fractions using the detection metrics in Sec.
+(4).
 The selection process for the values $\delta_{m}$ and $\delta_{s}$ is explained below.
-Dataset $\delta_{m}$ $\delta_{s}$ $d(\cdot)$
-CIFAR10 0.03 0.018 LPIPS
-LSUN-CHURCH 0.12 0.05 LPIPS
-MNIST 4.5 3.5 $L_{2}-$distance
-FASHION-MNIST 4.5 2.5 $L_{2}-$distance
-Table 2: Table displaying memorized and spurious threshold values in the computation of the transition plots in Figure 5.
-Selection of the Distance Metric.
-For high-dimensional datasets, CIFAR10 [63] and LSUN-CHURCH [66], we utilized LPIPS [77] with the AlexNet [78] backbone, as the function $d(\cdot,\cdot)$ for both memorization (13) and spurious (14) detection metrics.
+Table 1: A table showing the critical points $A$ and $B$ of the memorization-generalization transition for each dataset.11footnotetext: Note: We use 30 linearly spaced points between $A$ and $B$ to finely characterize the memorization-to-generalization transition.
+Training Data Size
+Dataset Point AA Point $B$ Total
+MNIST 4000 32000 60000
+CIFAR10 2000 16000 50000
+LSUN-CHURCH 2000 16000 126227
+FASHION-MNIST 4000 16000 60000
+B.2  Selection of Distance Metric
+For high-dimensional datasets, CIFAR10 [37] and LSUN-CHURCH [75], we utilized LPIPS [76] with the AlexNet [38] backbone, as the function $d(\cdot,\cdot)$ for both memorization and spurious detection metrics.
 We selected this approach since it is a commonly used perceptual metric that compares the similarity between two images based on their feature representations.
@@
 For simpler datasets like MNIST [17] and FASHION-MNIST [73], where images are single-channel and less complex, we found that $L_{2}$-distance suffices for both memorization and spurious detection metrics.
-Selection of Threshold Values $\delta_{m}$ and $\delta_{s}$.
-The detection thresholds, $\delta_{m}$ and $\delta_{s}$, were set based on the chosen distance metric $d(\cdot,\cdot)$ and visual inspection of the distance histograms for bimodality as mentioned in Section 4 of the main text.
-Moreover, these thresholds were chosen to reflect the varying visual complexity and feature richness of the datasets, and their values for each dataset are shown in Table 2.
-For CIFAR10, we utilized the histograms obtained at $K=7310$, which demonstrate strong bimodality shape, to tune our threshold values to minimize the number of spurious samples classified as memorized samples and vice versa, via manual inspection.
-For manual inspection, a general rule we followed — it is better to have more false positives in the generalized set than the memorized set since spurious samples can be seen as early-generalized samples.
-Thus, we observe the least-$5$ spurious samples and ensure that they do not resemble memorized samples.
-We performed the same process for memorized samples.
-The entire process is then repeated for the other three datasets.
-We utilized $K=4896$ for LSUN-CHURCH, $K=21379$ for MNIST, and $K=7724$ for FASHION-MNIST to select $\delta_{m}$ and $\delta_{s}$.
-For more details on the selected histograms, please refer to Figs.
-(4), (14), (12), and (13).
-Dataset Initial Latent Channel Multipliers Number of Parameters Batch Size Training Iterations
-MNIST 128 (1, 2, 2) 24.5M 128 400,000
-CIFAR10 128 (1, 2, 2, 2) 35.7M 128 500,000
-CIFAR10 96 (1, 2, 2, 2) 20.1M 128 500,000
-CIFAR10 64 (1, 2, 2, 2) 8.9M 128 500,000
-LSUN-CHURCH 64 (1, 1, 2, 2, 4, 4) 27.4M 64 800,000
-LSUN-CHURCH 96 (1, 1, 2, 2, 4, 4) 61.7M 64 800,000
-LSUN-CHURCH 128 (1, 1, 2, 2, 4, 4) 109.7M 64 800,000
-FASHION-MNIST 128 (1, 2, 2) 24.5M 128 400,000
-Table 3: Table displaying both model and training configurations for each dataset.
-Model and Training Details.
-For each point in our transition plots in Figure 5, we train a DDPM-based diffusion model, where the score model is a PixelCNN++ based U-Net [79, 80].
-We keep the variances, $\beta_{\text{min}}=10^{-4}$ and $\beta_{\text{max}}$ = $2\times 10^{-2}$, timesteps $T=1000$, and learning rate $lr=$ $2\times 10^{-4}$ for all models and datasets.
-Each model has $2$ residual blocks [81] for each down- and up- sampling layer, while an attention block is placed at 16×16$\times$16 × resolution.
-We only modified the channel multipliers for each model based on the complexity of the dataset, see Table 3.
-If the dataset size is smaller than the specified batch size, we take the batch size to be equal to that small dataset size.
-For generation or inference, we use the exponential moving average (EMA) of each trained model, as delineated in [23], which was obtained with the decay value set as $0.9999$ during training.
+B.3  Selection of Threshold Values
+The detection thresholds, $\delta_{m}$ and $\delta_{s}$, were set based on the chosen distance metric $d(\cdot,\cdot)$ and visual inspection of the distance histograms for bimodality as mentioned in Sec.
+(4).
+Moreover, these thresholds were chosen to reflect the varying visual complexity and feature richness of the datasets, and their values for each dataset are shown in Table (2).
+Specifically, for each training data size $K$ of each dataset, we manually inspected the bimodality of the olive and gray histograms to select $\delta_{m}$ and $\delta_{s}$.
+Moreover, to better tune these thresholds at each $K$, we manually inspected the least $5$ spurious samples, ensuring that they do not resemble memorized samples, while doing the same for memorized samples.
+If the resemblance of one set to another is too strong, we decreased the corresponding threshold.
+Please refer to Figs.
+(10)-(12) below and fig.˜3 in the main text for examples of the histograms.
+Please note that the bimodality of the gray histograms in FASHION-MNIST might not be significant, but it is enough to select their respective $\delta_{s}$.
+B.4  Model and Training Details
+For each point in our transition plots of fig.˜4, we trained a DDPM-based diffusion model, where the score model is a PixelCNN++ based U-Net [64, 53].
+We kept the variances, $\beta_{\text{min}}=10^{-4}$ and $\beta_{\text{max}}$ = $2\times 10^{-2}$, timesteps $T=1000$, and learning rate $lr=$ $2\times 10^{-4}$ for all models and datasets.
+Each model has $2$ residual blocks [21] for each down-sampling and up-sampling layer, while an attention block is placed at $16\times$ resolution.
+We only modified the channel multipliers for each model based on the complexity of the dataset, see Table (3).
+If the training data size $K$ is smaller than the specified batch size, the batch size is set to be equal to $K$.
+For generation or inference, we used the exponential moving average of each trained model, as delineated in [22], which is obtained with the decay value set as $0.9999$ during training.
 Please note that we did not use random flipping in the training of our models, since we want our measurements to reflect the training data size at best as random flipping implicitly increases the number of patterns that the models see during training.
 However, we did use dropout (of value $0.1$) for the training of CIFAR10, MNIST, and FASHION-MNIST models.
+For LSUN-CHURCH dataset, the images were center-cropped and down-sampled to $64\times 64$ resolution.
 Lastly, for each of the training set $\mathsf{S}_{\alpha}$, where $\alpha=1,\dots,M$, they are split from the original dataset given a specific size, using the same random seed value of $3407$.
+B.5  Additional Results
 Figure 7: The fractions of memorized, spurious, and generalized samples in synthetic sets across different training sizes for CIFAR10 [37] and various model’s widths.
@@
 In contrast, as the model’s size increases, the transition happens at an later stage (or larger $K$), where it is able to memorize more training data.
-The numbering, e.g., $64$, $96$, and $128$, denotes the initial latent dimension of the U-Net, see Table 3.
-Additional Experiments.
+The numbering, e.g., $64$, $96$, and $128$, denotes the initial latent dimension of the U-Net, see Table (3).
 Inspired by [74, 20], we extend our investigation on the memorization-generalization transition by studying the effects of varying the parameter size of the diffusion model for CIFAR10 [37] and LSUN-CHURCH [75].
 The results of this experiment are recorded in Figs.
-(9) and (10), respectively.
-Specifically, we trained two additional sets of 38 diffusion models, with the initial latent dimensions of 64 and 96 respectively, and repeat the same experimental setup delineated above, see also Table 3.
+(7)-(8) for these two datasets respectively.
+Specifically, we trained two additional sets of 38 diffusion models, with the initial latent dimensions of 64 and 96 respectively, and repeated the same experimental setup delineated above, see also Table (3).
 For CIFAR10 and the U-Net (64) models, we identified the critical points, $A$ and $B$, to be $500$ and $8000$ respectively.
@@
 To select $\delta_{m}$ and $\delta_{s}$, we followed the same process detailed above by observing the distance histograms to see where bimodality occurs with respect to the distance value, and manually inspecting the images to ensure there is a minimal amount of false positives in memorized and spurious samples.
-For U-Net (64) models, the threshold values $\delta_{m}$ and $\delta_{s}$ are both selected to be $0.02$ based on the bimodality of the distance histograms observed at $K=2568$ — meanwhile, for U-Net (96), the threshold values $\delta_{m}$ and $\delta_{s}$ are chosen as $0.03$ and $0.02$ according to distance histograms observed at $K=3931$.
 Similarly, for the LSUN-CHURCH models, we identified the critical points, $A$ and $B$, to be $1000$ and $8000$ for the U-Net (64) models; and for the larger U-Net (128) models, $4000$ and $16000$ as the critical points.
-The threshold values, $\delta_{m}$ and $\delta_{s}$, are both selected to be $0.04$ after observing the bimodality of the distances at $K=3413$ — meanwhile, for U-Net (128), these values are selected to be $\delta_{s}=0.04$ and $\delta_{m}=0.1$ after observing the distance histograms at $K=5241$.
+Please refer to Table (2) for the threshold values.
 Figure 8: The fractions of memorized, spurious, and generalized samples in synthetic sets across different training sizes for LSUN-CHURCH [75] and various model’s widths.
@@
 In contrast, as the model’s size increases, the transition happens at an later stage (or larger $K$), where it is able to memorize more training data.
-The numbering, e.g., $64$, $96$, and $128$, denotes the initial latent dimension of the U-Net, see Table 3.
-With the smaller U-Net (64) models, we observe the memorization-generalization transition happens at an earlier stage with respect to the larger U-Net (128) models, where the peak of spurious samples appears at $K=2568$ instead of $K=5862$ for CIFAR10 and $K=1000$ instead of $K=4000$ for LSUN-CHURCH.
-Similarly, for CIFAR10 and its U-Net (96) models, the peak of spurious samples happens at an early stage of $K=3931$ instead.
+The numbering, e.g., $64$, $96$, and $128$, denotes the initial latent dimension of the U-Net, see Table (3).
+With the smaller U-Net (64) models, we observe the memorization-generalization transition happens at an earlier stage with respect to the larger U-Net (128) models, where the peak of spurious samples appears at $K=2568$ instead of $K=5862$ for CIFAR10, and $K=4862$ instead of $K=6086$ for LSUN-CHURCH.
 Overall, these results indicate that as the parameter size of the diffusion model increases, the memorization capacity of the model is increased and thus, generalization is delayed since the model can memorize more of its training data denoted by the appearance of spurious patterns happening at a larger training data size $K$.
-In contrast, when the memorization capacity is reduced, generalization is initiated at an earlier stage (or smaller training data size) with respect to when the parameter size of the model is larger.
-Most importantly, the emergence of spurious patterns remains consistent despite the variations in the parameter size of the diffusion model.
+In contrast, when the memorization capacity is reduced, generalization is initiated at an earlier stage (or smaller training data size) given a larger parameterized model.
+Most importantly, the emergence of spurious patterns remains consistent despite the variations in the parameter size of the diffusion model: these patterns quickly rise and fall at the boundaries of the memorization-to-generalization transition.
 Figure 9: Memorization-Generalization transition plots for the CIFAR10 dataset.
-The left figure depicts the memorization-generalization transition computed with 4×4$\times$4 × synthetic sets for various training data sizes.
-Meanwhile, right figure shows the transition computed using 8×8$\times$8 × synthetic sets for the same training data sizes.
-With the larger synthetic sets, more spurious patterns are detected at the peak ($K=5862$).
+The column (a) figure depicts the memorization-generalization transition computed with $4\times$ synthetic sets for various training data sizes.
+Meanwhile, the column (b) figure shows the transition computed using $8\times$ synthetic sets for the same training data sizes.
+With the larger synthetic sets, more spurious patterns are detected around the peak of the transition.
 However, the overall trend of these patterns quickly rise, during the co-existence of memorization and generalization, and fall as generalization becomes the dominant phase, is maintained.
-Both figures use the same memorized and spurious threshold values detailed in Table 2.
-Additionally, to verify the consistency of the memorization-generalization transition, we explored the influence of larger synthetic sets and recorded our results in Figure 11.
-Specifically, we repeated our experiment done in Figure 5 for CIFAR10 using 8×8$\times$8 × synthetic sets — which was originally done using 4×4$\times$4 × synthetic sets.
-Based on Figure 11, we observe a slight increase in the fraction of spurious samples when using the larger 8×8$\times$8 × synthetic sets.
+Both figures use the same memorized and spurious threshold values detailed in Table (2).
+Additionally, to verify the consistency of the memorization-generalization transition, we explored the influence of larger synthetic sets and recorded our results in Fig.
+(9).
+Specifically, we repeated our experiment done in Fig.
+(4) for CIFAR10 using $8\times$ synthetic sets, which was originally done using $4\times$ synthetic sets.
+Based on these results, we observe a slight increase in the fraction of spurious samples when using the larger $8\times$ synthetic sets.
 Specifically, at $K=5862$, the fraction increases from $0.04$ to roughly $0.06$ with respect to the synthetic set size.
-Nonetheless, the trend of memorization-generalization, with the emergence of spurious samples, is still maintained regardless of the multiplier for the synthetic size — highlighting the consistency of the spurious phenomenon in diffusion models in relation to the training data size.
-Hardware Details.
-The training of the models detailed in Table 3 was done using NVIDIA Tesla V100 GPUs, where a single GPU was used to train each model.
-Each GPU has 32GB of memory and is linked with IBM Power9 processors which clocked at 3.15 GHz.
-Each CIFAR10 model took roughly 18, 20, and 22 hours to train for different initial latent dimensions, e.g., 64, 96, 128, respectively.
-Meanwhile, each LSUN-CHURCH model requires roughly 48, 56, and 103 hours of training with respect to the model’s initial latent dimension.
-In contrast, FASHION-MNIST and MNIST models require approximately 17 hours to be fully trained (per model).
-Please note that these durations are with respect to the required training iterations for each model and their corresponding dataset detailed in Table 3.
-Please note that we utilized the same hardware for our experiments as well.
-Appendix E Additional Visualizations of Different Sample Types Across the Transition
-E.1 LSUN-CHURCH
+Nonetheless, the trend of memorization-generalization, with the emergence of spurious samples, is still maintained regardless of the multiplier for the synthetic size.
+This aspect highlights the consistency of the spurious phenomenon in diffusion models in relation to the varying of training data size.
+Table 2: Table displaying the range of the memorized and spurious threshold values in the computation of the transition plots in Fig.
+(4) of the main text.11footnotetext: The full set of threshold values is provided alongside with the code.
+Range [min, max]
+Dataset Initial Latent Dim. $\delta_{m}$ $\delta_{s}$
+MNIST 128 $[2,4.5]$ $[2,4]$
+FASHION-MNIST 128 $[2,2.5]$ $[2.5,3]$
+64 $[0.01,0.04]$ $[0.015,0.025]$
+CIFAR10 96 $[0.015,0.04]$ $[0.015,0.025]$
+128 $[0.015,0.035]$ $[0.015,0.025]$
+64 $[0.06,0.15]$ $[0.01,0.1]$
+LSUN-CHURCH 96 $[0.1,0.125]$ $[0.05,0.06]$
+128 $[0.1,0.12]$ $[0.04,0.06]$
+Table 3: A table displaying both model and training configurations for each dataset.
+Configurations
+Dataset Initial Latent Dim. Channel Multipliers Num. of Param. Batch Size Training Iterations
+MNIST 128 (1, 2, 2) 24.5M 128 400,000
+FASHION-MNIST 128 (1, 2, 2) 24.5M 128 400,000
+64 (1, 2, 2, 2) 8.9M 128 500,000
+CIFAR10 96 (1, 2, 2, 2) 20.1M 128 500,000
+128 (1, 2, 2, 2) 35.7M 128 500,000
+64 (1, 1, 2, 2, 4, 4) 27.4M 64 800,000
+LSUN-CHURCH 96 (1, 1, 2, 2, 4, 4) 61.7M 64 800,000
+128 (1, 1, 2, 2, 4, 4) 109.7M 64 800,000
+B.6  Additional Visualizations of Distance Histogram
 Figure 10: Different sample types across the memorization-to-generalization transition for LSUN-CHURCH [75], defined by the spurious and memorized thresholds, $\delta_{s}$ and $\delta_{m}$.
@@
 The inset shows amplified spurious fraction (green curve).
-E.2 FASHION-MNIST
 Figure 11: Different sample types across the memorization-to-generalization transition for FASHION-MNIST [73], defined by the spurious and memorized thresholds, $\delta_{s}$ and $\delta_{m}$.
@@
 The inset shows amplified spurious fraction (green curve).
-E.3 MNIST
 Figure 12: Different sample types across the memorization-to-generalization transition for MNIST [17], defined by the spurious and memorized thresholds, $\delta_{s}$ and $\delta_{m}$.
@@
 The inset shows amplified spurious fraction (green curve).
-Appendix F Additional Visualizations of the Three Sample Types
-F.1 Memorized Samples
+B.7  Additional Examples of the Three Distinct Sample Types
 Figure 13: Visualizations of additional memorized patterns and their top-4 nearest neighbors for different datasets.
 The top row illustrates nearest neighbors from the training set while the bottom row depicts those from the synthetic set.
-Memorized samples are duplicates of the training set.
-During the strong memorization phase, duplicates are also found within the synthetic set.
-Note, even though our memorized detection metric (Eq.
-13) does not utilize the synthetic set, we are showing the nearest neighbors obtained from it, for consistency.
-F.2 Spurious Samples
+Memorized samples are duplicates of the training set $\mathsf{S}$.
+During the strong memorization phase, duplicates are also found within the synthetic set $\mathsf{S}^{\prime}$.
+Note, even though our memorized detection metric does not utilize the synthetic set, we are showing the nearest neighbors obtained from it, for consistency.
 Figure 14: Visualizations of additional spurious patterns and their top-4 nearest neighbors for different datasets.
@@
 Spurious patterns are demonstrated to arise from the onset of generalization where the mixing of training data points begins.
-Since the model’s generalization is at its infancy, duplicates of spurious patterns sometimes appear several times in the synthetic set $S^{\prime}$, much like memorized patterns.
+Since the model’s generalization is at its infancy, duplicates of spurious patterns sometimes appear several times in the synthetic set $\mathsf{S}^{\prime}$, much like memorized patterns.
 Additionally, these samples lack the uniqueness to be considered as generalized samples as the model has yet to fully learn the underlying data distribution.
-F.3 Generalized Samples
 Figure 15: Visualization of generalized patterns and their top-4 nearest neighbors for different datasets.
-The top row illustrates nearest neighbors from the training set while the bottom row depicts those from the synthetic set.
+The top row illustrates nearest neighbors from the training set $\mathsf{S}$ while the bottom row depicts those from the synthetic set $\mathsf{S}^{\prime}$.
 Generalized samples are novel samples, which have little to no resemblance to their nearest neighbors in training and synthetic sets.
-Appendix G Experimental Details on the Basin of Attraction Analysis Across the Transition
-Algorithm Details.
-To obtain the results recorded in the above Figs.
-(18) and (19), we first computed the critical time $t_{c}$ given an image ${\mathbf{x}}_{0}={\mathbf{y}}$.
-Specifically, this is done by using Algorithm (1), where $t_{c}$ is obtained using the deterministic DDIM [24] sampler to denoise the perturbed version ${\mathbf{x}}_{t}$ of ${\mathbf{x}}_{0}$ for each step in $t=1,\dots,T$, given $N^{\prime}$ number of denoising steps starting backward from $t$ to 00.
+Appendix C The Volume of the Basin of Attraction
+C.1  Algorithm Details
+To obtain the average log-volume results recorded in fig.˜4 of the main text, we first computed the critical time $t_{c}$ given an image $\hat{{\mathbf{x}}}={\mathbf{x}}_{0}$.
+Specifically, this is done by using Alg.
+(1), where $t_{c}$ is obtained using the deterministic DDIM [60] sampler to denoise the perturbed version ${\mathbf{x}}_{t}$ of ${\mathbf{x}}_{0}$ for each step in $t=1,\dots,T$, given $N^{\prime}$ number of denoising steps starting backward from $t$ to 0.
 To save computational time, the algorithm utilizes a stride constant $s$ to skip over some of the timesteps $t=1,1+s,1+2s,\dots,T-s-1$.
@@
 To account for the variations in noise, the image ${\mathbf{x}}_{0}$ is perturbed for $M$ number of times at each step $t$.
-To check whether if the critical time $t_{c}$ is reached, the algorithm utilizes a distance metric $d$ — which is LPIPS [77] equipped with the vgg-16 backbone [82] for all of the datasets — checking if a recovered image $\hat{{\mathbf{x}}}_{0}$ obtained from the DDIM sampler [24] with $N^{\prime}$ denoising steps is similar to the original pattern ${\mathbf{x}}_{0}$ within an error threshold $\delta_{d}\in(0,1)$.
-Unlike the experiment on the memorization-generalization transition detailed in Section 5, we can afford to trade computational time for better accuracy in this experiment by using a larger backbone for LPIPS.
-Please note that the metric LPIPS is bounded between 0 (most similar) and 1 (least similar).
-To keep track of the critical time $t_{c}$ for all perturbations, we allow it to be a vector of integers $t_{c}\in\mathbb{N}^{M}$, where $M$ is the number of perturbations or trials for the image ${\mathbf{x}}_{0}$.
-To ensure that $t_{c}$ is correct, we have a binary stopping vector $\hat{S}\in\{0,1\}^{M}$, which keeps track of whether each perturbation of ${\mathbf{x}}_{0}$ is no longer recoverable, such that we only update an entry of the vector $t_{c}$ if and only if the corresponding entry in $\hat{S}$ is a zero (not stopped).
-Since the entire process of finding $t_{c}$ is a rough search over a set of timesteps, we require a “breaking condition” for the search.
-In Algorithm 1, this condition is determined by the value $p\in[0,1]$, which is a probability corresponding to the recovery probability of the image ${\mathbf{x}}_{0}$ with respect to its $M$ perturbations.
-To compute $p$, we calculate the distance vector $\mathbf{d}=d(\hat{{\mathbf{x}}}_{0},{\mathbf{x}}_{0})$ where $\hat{{\mathbf{x}}}_{0}$ is the predicted version of ${\mathbf{x}}_{0}$ obtained from denoising ${\mathbf{x}}_{t}$.
-Then, we compute the recovery binary vector $\mathbf{r}\in[0,1]^{M}$ by applying the indicator function onto $\mathbf{d}$ using the threshold $\delta_{d}$.
-Lastly, we calculate the probability value $p=\frac{1}{M}\sum^{M}_{i=1}\mathbf{r}_{i}$.
+[…]
+Since the entire process of finding $t_{c}$ is a rough search over a set of timesteps, we require a ‘breaking condition’ for the search.
+In Alg.
+(1), this condition is determined by the value $p\in[0,1]$, which is a probability corresponding to the recovery probability of the image ${\mathbf{x}}_{0}$ with respect to its $M$ perturbations.
+To compute $p$, we calculated the distance vector $\mathbf{d}=d(\hat{{\mathbf{x}}}_{0},{\mathbf{x}}_{0})$ where $\hat{{\mathbf{x}}}_{0}$ is the predicted version of ${\mathbf{x}}_{0}$ obtained from denoising ${\mathbf{x}}_{t}$.
+Then, we computed the recovery binary vector $\mathbf{r}\in[0,1]^{M}$ by applying the indicator function onto $\mathbf{d}$ using the threshold $\delta_{d}$.
+Lastly, we calculated the probability value $p=\frac{1}{M}\sum^{M}_{i=1}\mathbf{r}_{i}$.
 If this value $p$ is below the probability threshold $\delta_{p}$, we stop the search for critical time across all trials.
-Experimentation Details.
-For each training data size $K$, we used a maximum of the top 512 samples belonging to each of the sample types — e.g., training data, memorized, spurious, and generalized samples — to find their critical time $t_{c}$ using Algorithm 1.
-Note, with the exception of the training data points, we sorted samples belonging to each set — identified via the detection metrics (13), (14) and (15) — by their distance from least to greatest.
-We computed the log volume using the following equation
-$V({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}}=\frac{\pi^{\frac{N}{2}}}{\Gamma(\frac{N}{2}+1)}R({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}})^{N}$ (33)
-where $\Gamma(\cdot)$ is the Euler’s gamma function [83] and $N$ is the dimensionality of the image.
-The radius $R({\mathbf{x}}_{0},{\mathbf{x}}_{t_{c}})$ is the Euclidean norm of the difference between the original image ${\mathbf{x}}_{0}$ and its recoverable perturbation ${\mathbf{x}}_{t_{c}}$.
+Inputs
+  1.5pt Pattern $\hat{{\mathbf{x}}}$
+  1.5pt Time stride $s$
+  1.5pt Score model $s_{\theta}$
+  1.5pt DDIM sampler $\Phi$
+  1.5pt Error threshold $\delta_{d}$
+  1.5pt Distance metric $d$
+  1.5pt Denoising steps $N^{\prime}$
+  1.5pt Total diffusion time $T$
+  1.5pt Number of trials $M$
+  1.5pt Probability threshold $\delta_{p}$
+  1.5pt Forward process $\mathcal{F}$
+end
+Compute $t_{c}$
+  5pt ${\mathbf{x}}_{0}\leftarrow\text{duplicate}(\hat{{\mathbf{x}}},M)$
+ // Duplicate pattern for $M$ times
+  5pt $t_{c}\leftarrow\mathbf{0}$
+ // Initialize critical time tracker
+  5pt $\hat{S}\leftarrow\mathbf{0}$
+ // Track trials that have stopped
+  5pt for $t=1,1+s,1+2s,\dots,T-s-1$ do
+     5pt $\epsilon\sim\mathcal{N}(0,I)$
+    // Sample noise
+     5pt
+    ${\mathbf{x}}_{t}\leftarrow\mathcal{F}({\mathbf{x}}_{0},\epsilon,t)$
+    // Forward process
+     5pt
+    $\hat{{\mathbf{x}}}_{0}\leftarrow\Phi(s_{\theta},{\mathbf{x}}_{t},t,N^{\prime})$
+    // DDIM denoising
+     5pt
+    $\mathbf{d}\leftarrow d(\hat{{\mathbf{x}}}_{0},{\mathbf{x}}_{0})$
+    // Compute distance per trial
+     8pt
+    $\mathbf{r}_{i}\leftarrow\begin{cases}1&\text{if }\mathbf{d}_{i}\leq\delta_{d}\\ 0&\text{otherwise}\end{cases}$ for $i=1,\dots,M$
+    // Binary recovery vector
+     8pt
+    $p\leftarrow\frac{1}{M}\sum_{i=1}^{M}\mathbf{r}_{i}$
+    // Recovery probability
+     8pt
+    if $p<\delta_{p}$ then
+       return $t_{c}$
+    else
+       $\hat{S}\leftarrow\text{update}(\hat{S},\mathbf{r})$
+[…]
+      end if
+   end for
+end
+Output
+  5pt Critical time $t_{c}\in\mathbb{N}^{M}$
+end
+Algorithm 1 Compute critical time $t_{c}$ for a pattern
+C.2  Experimental Details
+For each training data size $K$, we used a maximum of the top 512 samples belonging to each of the sample types: training data, memorized, spurious, and generalized samples.
+Note, with the exception of the training data points, we sorted samples belonging to each set, identified via our detection metrics, by their distance from least to greatest.
+We computed the logarithmic version of the following $N$-ball volume equation:
+$V(\hat{{\mathbf{x}}},{\mathbf{x}}_{t_{c}})=\frac{\pi^{\frac{N}{2}}}{\Gamma(\frac{N}{2}+1)}R(\hat{{\mathbf{x}}},{\mathbf{x}}_{t_{c}})^{N}$ (34)
+using the $t_{c}$ we computed from Alg.
+(1) where $\Gamma(\cdot)$ is the Euler’s gamma function [55] and $N$ is the dimensionality of the image.
+The radius $R(\hat{{\mathbf{x}}},{\mathbf{x}}_{t_{c}})$ is the Euclidean norm of the difference between a given image $\hat{{\mathbf{x}}}\in\mathbb{R}^{N}$ and its recoverable perturbation ${\mathbf{x}}_{t_{c}}$ computed from the forward process.
 For Figs.
-(6), (18), and (19), we plotted the average log-volume value per each training data size $K$.
-Meanwhile, for Algorithm 1, we fixed the stopping probability $\delta_{p}$ as $0.8$, the number of perturbations $M$ as $20$, the DDIM steps $N^{\prime}$ as $10$ and the stride constant $s$ as $10$ for all datasets.
+(16)-(17), we plotted the average log-volume value calculated for each training data size $K$ per dataset.
+Meanwhile, for Alg.
+(1), we fixed the stopping probability $\delta_{p}$ as $0.8$, the number of perturbations $M$ as $20$, the DDIM steps $N^{\prime}$ as $10$ and the stride constant $s$ as $10$ for all datasets.
 At some training data size $K$, in which there are very few samples in either memorized, spurious or generalized set, we opted to not include the statistics of such a set, partly due to its large variance.
@@
 For CIFAR10 and LSUN-CHURCH, we utilized the threshold $\delta_{d}$ as $0.03$ and $0.12$ respectively.
-We kept these threshold values similar to the memorized threshold values $\delta_{m}$ detailed in Table 2, since they have proven to be effective at separating the memorized samples from non-memorized ones in the previous experiment.
-However, since it is difficult to select the threshold value for the $L_{2}$-distance metric — because the range of the metric is between 00 and $\infty$ — we opted to use LPIPS to experiment with MNIST and FASHION-MNIST using the same threshold $\delta_{d}=0.1$ for both.
+We kept these threshold values similar to the memorized threshold values $\delta_{m}$ detailed in Table (2), since they have proven to be effective at separating the memorized samples from non-memorized ones in the previous experiment.
+However, since it is difficult to select an appropriate threshold value for the $L_{2}$-distance metric (due to its range being between 0 and $\infty$), we opted to use LPIPS to experiment with MNIST and FASHION-MNIST using the same threshold $\delta_{d}=0.1$ for both.
 To find this optimal threshold value, we selected three data sizes, where each belongs to one of the three phases: memorization, spurious, or generalization.
@@
 Specifically, we evaluated these thresholds using $K\in\{4000,11724,60000\}$ for MNIST, and $K\in\{4000,8965,60000\}$ for FASHION-MNIST.
-We simply perform Algorithm 1 for those three data sizes on a small batch of $32$ samples for the three sample types, and manually check whether the recovered samples (using the obtained critical time $t_{c}$ to denoise from) are similar to the original images and if this observation is consistent across those three data sizes.
-G.1 Algorithm for Computing the Critical Time
-1
-2
-3Inputs
-4        Pattern ${\mathbf{y}}$
-5        Score model $s_{\theta}$
-6        DDIM sampler $\Phi$
-7        Error threshold $\delta_{d}$
-8        Distance metric $d$
-9        Denoising steps $N^{\prime}$
-10        Total diffusion time $T$
-11        Time stride $s$
-12        Number of trials $M$
-13        Probability threshold $\delta_{p}$
-14        Cumulative variances $\bar{\alpha}$
-15        Variances $\beta$
-16Compute $t_{c}$
-        ${\mathbf{x}}_{0}\leftarrow\text{duplicate}({\mathbf{y}},M)$
-         // Duplicate pattern ${\mathbf{y}}$ for $M$ times
-        $t_{c}\leftarrow\mathbf{0}$
-         // Initialize critical time tracker
-        $\hat{S}\leftarrow\mathbf{0}$
-         // Track trials that have stopped
-17        for $t=1,1+s,1+2s,\dots,T-s-1$ do
-               ${\mathbf{z}}_{t}\sim\mathcal{N}(0,\mathbf{I})$
-                // Sample noise
-18
-              ${\mathbf{x}}_{t}\leftarrow\sqrt{\bar{\alpha}_{t}}{\mathbf{x}}_{0}+\sqrt{1-\bar{\alpha}_{t}}{\mathbf{z}}_{t}$
-                // Forward diffusion
-19
-              $\hat{{\mathbf{x}}}_{0}\leftarrow\Phi(s_{\theta},{\mathbf{x}}_{t},t,\bar{\alpha},\beta,N^{\prime})$
-                // DDIM denoising
-20
-              $\mathbf{d}\leftarrow d(\hat{{\mathbf{x}}}_{0},{\mathbf{x}}_{0})$
-                // Compute distance per trial
-21
-              $\mathbf{r}_{i}\leftarrow\begin{cases}1&\text{if }\mathbf{d}_{i}\leq\delta_{d}\\ 0&\text{otherwise}\end{cases}$ for $i=1,\dots,M$
-                // Binary recovery vector
-22
-              $p\leftarrow\frac{1}{M}\sum_{i=1}^{M}\mathbf{r}_{i}$
-                // Recovery probability
-23
-24              if $p<\delta_{p}$ then
-25                      return $t_{c}$
-26              else
-27                      $\hat{S}\leftarrow\text{update-stop}(\hat{S},\mathbf{r})$
-28                      $t_{c}\leftarrow\text{update-time}(t,t_{c},\hat{S})$
-29
-30
-31Output
-32        Critical time $t_{c}\in\mathbb{N}^{M}$
-Algorithm 1 Compute Critical Time $t_{c}$ for a single pattern ${\mathbf{y}}$
-G.2 The Basin of Attraction Across the Transition for All Datasets
-Figure 18: Average log-volume of the hyperspheric basin of attraction of different sample types as the training data size $K$ grows, in CIFAR10 and LSUN-CHURCH computed using Equation 33 with up to 512 samples per type.
+We simply performed Alg.
+(1) for those three data sizes on a small batch of $32$ samples for the three sample types, and manually checked whether the recovered samples (using the obtained critical time $t_{c}$ to denoise from) are similar to the original images and if this observation is consistent across those three data sizes.
+C.3  Average Log-Volume Results
+Figure 16: Average log-volume of the hyperspheric basin of attraction of different sample types as the training data size $K$ grows, in CIFAR10 [37] and LSUN-CHURCH [75] using Eq.
+(34) with up to 512 samples per type.
 When fewer than 512 samples are available, all points are used.
-The log-volume of training samples and the transition plot from Figure 5 are included to confirm the shrinking basins of attraction in diffusion models.
-Variations in the critical time $t_{c}$ during the transition, which affect the radius $R$, lead to distinct log-volume trends across sample types, see the recorded critical times in Figure 20.
+The average log-volume of training samples and the transition plot are included to confirm the shrinking basins of attraction in diffusion models.
+Variations in the critical time $t_{c}$ during the transition, which affect the radius $R$, lead to distinct log-volume trends across sample types, see the recorded critical times in Fig.
+(18).
 Shaded regions indicate standard deviation of the log-volume.
-Figure 19: Average log-volume of the hyperspheric basin of attraction of different sample types as the training data size $K$ grows, in FASHION-MNIST and MNIST, computed using Equation 33 with up to 512 samples per type.
+Figure 17: Average log-volume of the hyperspheric basin of attraction of different sample types as the training data size $K$ grows, in FASHION-MNIST [73] and MNIST [17], computed using Eq.
+(34) with up to 512 samples per type.
 When fewer than 512 samples are available, all points are used.
-The log-volume of training samples and the transition plot from Figure 5 are included to confirm the shrinking basins of attraction in diffusion models.
-Variations in critical time $t_{c}$ and radius $R$ lead to distinct log-volume trends across sample types, see the recorded critical times in Figure 21.
+The average log-volume of training samples and the transition plot are included to confirm the shrinking basins of attraction in diffusion models.
+Variations in critical time $t_{c}$ and radius $R$ lead to distinct log-volume trends across sample types, see the recorded critical times in Fig.
+(19).
 Shaded regions indicate standard deviation of the log-volume.
-G.3 Average Critical Time Across The Transition
-Figure 20: Average critical time, obtained using Algorithm 1, with a maximum of 512 samples across the memorization-generalization transition for each sample type in CIFAR10 and LSUN-CHURCH.
-The average critical times for the training data samples at each training data size $K$ are included to illustrate the overall decrease in the time and also the amount of perturbation in which the diffusion model can effectively recover the original sample within some error.
+C.4  Average Critical Time Results
+Figure 18: Average critical time, obtained using Alg.
+(1), with a maximum of 512 samples across the memorization-generalization transition for each sample type in CIFAR10 [37] and LSUN-CHURCH [75] The average critical times for the training data samples at each training data size $K$ are included to illustrate the overall decrease in the time and also the amount of perturbation in which the diffusion model can effectively recover the original sample within some error.
 Each sample type exhibits a distinctive critical time at each training data size $K$.
@@
 The transition plot of each dataset is shown here to visually guide where the memorization-generalization transition occurs.
-Figure 21: Average critical time, obtained using Algorithm 1, with a maximum of 512 samples across the memorization-generalization transition for each sample type in FASHION-MNIST and MNIST.
-The average critical times for the training data samples at each training data size $K$ are included to illustrate the overall decrease in the time and also the amount of perturbation in which the diffusion model can effectively recover the original sample within some error.
+Figure 19: Average critical time, obtained using Alg.
+(1), with a maximum of 512 samples across the memorization-generalization transition for each sample type in FASHION-MNIST [73] and MNIST [17] The average critical times for the training data samples at each training data size $K$ are included to illustrate the overall decrease in the time and also the amount of perturbation in which the diffusion model can effectively recover the original sample within some error.
 Each sample type exhibits a distinctive critical time at each training data size $K$.
@@
 The transition plot of each dataset is shown here to visually guide where the memorization-generalization transition occurs.
-G.4 Visual Demonstration of the Shrinkage of the Basins of Attraction
-Experiment Details.
-For Figure 7, we randomly selected a subset of training images that are shared across all 38 diffusion models, each trained with a different training set size.
-To determine an appropriate timestep $t\approx t_{c}$, we referred to Figure 20, which plots the average critical time $t_{c}$ as a function of training data size $K$.
-From this plot, we identified a plateau point in the memorization fraction curve, occurring around $K=3931$, and selected $t_{c}\approx 450$ for CIFAR10.
-Using the timestep starting at $450$, each image was perturbed via the forward process and then denoised using the DDIM sampler [24] with 25 steps across all models (starting backward from the indicated timestep).
-We manually examined the denoised outputs across training sizes to assess whether the resulting generations effectively illustrate the shrinkage of the basin of attraction.
-If the output appeared overly memorized (or too close to the original), we increased the timestep by increments of 10 and repeated the process to obtain a more illustrative transition.
-Lastly, in the case of LSUN-CHURCH, we selected the initial timestep to be $550$ accordingly to Figure 20.
-Appendix H Additional Details on the Relative Energy Computation and its Experimentation
-In this section, we provide details on the formulation of the relative energy done by [38] and an overview on the experiment done in Section 7 of the main text.
-Relative Energy Formulation. Consider the SDE for the reverse process in Equation 2 of the main text
-$\mathrm{d}{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})]\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (34)
-We can re-express the above equation with the potential function $u(x,t)$
-$-\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)=\mathbf{f}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$ (35)
+Appendix D The Curvature of the Energy
+D.1  Overview
+Following [3, 66], we utilized their stable version of algorithm for computing the singular values of the Jacobian of the score function, detailed in Alg.
+(2).
+The motivations for the algorithm are first detailed in [63] and later expanded by [3, 66].
+Here we recite the motivations laid out in [66].
+Consider a generative diffusion model with $p_{0}({\mathbf{x}})$ defined on a $d$-dimensional manifold $\mathcal{M}$, where $\mathcal{M}=\mathcal{M}_{0}$ denotes the manifold at $t=0$.
+We can define a time-dependent set of points:
+$\mathcal{M}_{t}=\{{\mathbf{x}}^{*}|\tilde{s}_{\mathcal{M}}({\mathbf{x}}^{*},t)=0,\,\,\text{with}\,\,J_{\mathcal{M}}({\mathbf{x}}^{*},t)\,\,\text{n.s.d.}\}$ (35)
+which [66] name the stable latent set of the diffusion process.
+The negative semi-definiteness (n.s.d.) is a stability condition of the Jacobian matrix $J_{\mathcal{M}}({\mathbf{x}}^{*},t)$ of the support score $\tilde{s}_{\mathcal{M}}({\mathbf{x}}^{*},t)$, defined as the score function obtained from the uniform data distribution $\tilde{p}_{0}({\mathbf{x}})=\frac{1}{|\mathcal{M}|}\delta_{\mathcal{M}}({\mathbf{x}})$.
+Using the relationship $s_{\theta}({\mathbf{x}},t)=-\frac{\epsilon_{\theta}({\mathbf{x}},t)}{\sigma_{t}}$, the diffusing particles typically explore shells of a radius that concentrates on $\sigma_{t}$.
+Then, for a small perturbation $\mathbf{p}$} around a state $\mathbf{x}$$^{*}$ on the latent manifold $\mathcal{M}_{t}$, the score function is well approximated by linearization:
+$s({\mathbf{p}},t)\approx J({\mathbf{x}}^{*},t){\mathbf{p}}=\sum_{j}({\mathbf{v}}_{j}\cdot{\mathbf{p}})\lambda_{j}({\mathbf{x}}^{*},t){\mathbf{v}}_{j}$ (36)
+where ${\mathbf{v}}_{j}$ and $\lambda_{j}$ are the respective $j$-th eigenvector and eigenvalue of the negative Jacobian $J({\mathbf{x}}^{*},t)$.
+Small perturbations, which aligned with the tangent space of $\mathcal{M}_{t}$, correspond to small eigenvalues; while orthogonal perturbations correspond to high eigenvalues, as the score tends to push the stochastic dynamics towards its fixed-points.
+Hence, based on these relationships, the spectrum of eigenvalues can provide detailed information regarding the local geometry of the stable latent set (35).
+Moreover, as already indicated by [63], one can estimate the dimensionality of the manifold from the location of a drop, or a sharp change, in the sorted spectrum of eigenvalues.
+The full range of the spectra results is shown in Figs.
+(20)-(23).
+D.2  Experimental Details
+Using Alg.
+(2), we computed the singular values of the Jacobian of the learnt score function $s_{\theta}$ given a small time $t_{0}$ for all datasets shown in Fig.
+(5A) of the main text.
+For each sample type, we utilized a maximum of 512 samples and computed the singular values at $t_{0}=2$ for all datasets and training data sizes.
+The singular values are sorted from greatest to least, and the top-5 values are ignored to avoid the case of very large singular values skewing the visualizations.
+Meanwhile, for the spectra results of the training data points shown in the bottom row of Fig.
+(5A), we utilized $\min(K,2048)$ number of training samples for all datasets and training data sizes.
+In this visualization, we utilized the rainbow color scheme and plotted the singular values for small training data sizes, starting from the purple color, to larger ones, ending at the red color.
+Finally, regarding our experiment with Stable Diffusion [50], following [72, 30] we utilized its version 1.4 from the package Diffusers [69].
+The candidate memorized and spurious prompts we utilized are originally found by [71, 70].
+Additional examples are shown in Fig.
+(24).
+Meanwhile, for candidate generalized samples, their prompts are obtained from LAION2B-Aesthetic [54].
+Instead of using Alg.
+(2), we opted for Alg.
+(3) and computed the singular values of the Jacobian of the score using $t_{0}=5$, due to the lack of computational resources.
+Please note that according to [66] Alg.
+(2) produces a much smaller number of extreme singular values than the later algorithm.
+For generation of the samples, we utilized DDIM [60] sampler with different noise vectors initialized from a fixed seed.
+For the visualizations, we skipped the top-50 singular values and set the limit for the y-axis as 1500.
+Inputs
+   Target sample ${\mathbf{x}}_{0}\in\mathbb{R}^{N}$
+  2.5pt Time $t_{0}$
+  2.5pt Score model $s_{\theta}$
+  2.5pt Forward process $\mathcal{F}$
+end
+[…]
+// Create an empty matrix
+1exCompute matrix $\mathsf{M}$
+ for $i=1,2,\dots,4N$ do
+     5pt $\epsilon\sim\mathcal{N}(0,I)$
+    // Sample noise
+    ${\mathbf{x}}^{+}\leftarrow\mathcal{F}({\mathbf{x}}_{0},\epsilon,t_{0})$
+    // Left perturbation
+    ${\mathbf{x}}^{-}\leftarrow\mathcal{F}({\mathbf{x}}_{0},-\epsilon,t_{0})$
+    // Right perturbation
+    $\mathsf{M}^{(i)}\leftarrow\frac{s_{\theta}({\mathbf{x}}^{+},t_{0})-s_{\theta}({\mathbf{x}}^{-},t_{0})}{2}$
+    // Add a new column to $\mathsf{M}$
+   end for
+end
+Output
+  5pt $\{s_{i}\}^{N}_{i=1}\leftarrow\text{SVD}(\mathsf{M})$
+ // Compute singular values
+end
+Algorithm 2 Compute singular values of the Hessian of the energy using central difference for a pattern
+Inputs
+   Target sample ${\mathbf{x}}_{0}\in\mathbb{R}^{N}$
+  2.5pt Time $t_{0}$
+  2.5pt Score model $s_{\theta}$
+  2.5pt Forward process $\mathcal{F}$
+end
+[…]
+// Create an empty matrix
+1exCompute matrix $\mathsf{M}$
+ for $i=1,2,\dots,4N$ do
+     5pt $\epsilon\sim\mathcal{N}(0,I)$
+    // Sample noise
+    ${\mathbf{x}}^{+}\leftarrow\mathcal{F}({\mathbf{x}}_{0},\epsilon,t_{0})$
+    // Perturbation
+    $\mathsf{M}^{(i)}\leftarrow s_{\theta}({\mathbf{x}}^{+},t_{0})$
+    // Add a new column to $\mathsf{M}$
+   end for
+end
+Output
+  5pt $\{s_{i}\}^{N}_{i=1}\leftarrow\text{SVD}(\mathsf{M})$
+ // Compute singular values
+end
+Algorithm 3 Compute singular values of the Hessian of the energy for a pattern
+D.3  Curvature Results
+Figure 20: Average singular value spectra of the score’s Jacobian for memorized, spurious, and generalized samples from diffusion models trained on the LSUN-CHURCH [75] dataset, shown across 38 increasing training data sizes $K$.
+The plots reveal a clear geometric hierarchy in the intermediate training regime, where memorized samples consistently exhibit the highest curvature (largest singular values), followed by spurious, and finally generalized samples.
+As the model enters the full generalization phase at large $K$, the average spectra for all three sample types become very distinguishable.
+Shaded region is the standard deviation of the singular values of different samples of each category.
+Figure 21: Average singular value spectra of the score’s Jacobian for memorized, spurious, and generalized samples from diffusion models trained on the FASHION-MNIST [73] dataset, shown across 38 increasing training data sizes $K$.
+The plots reveal a clear geometric hierarchy in the intermediate training regime, where memorized samples consistently exhibit the highest curvature (largest singular values), followed by spurious, and finally generalized samples.
+As the model enters the full generalization phase at large $K$, the average spectra for all three sample types become very distinguishable.
+Shaded region is the standard deviation of the singular values of different samples of each category.
+Figure 22: Average singular value spectra of the score’s Jacobian for memorized, spurious, and generalized samples from diffusion models trained on the MNIST [17] dataset, shown across 38 increasing training data sizes $K$.
+The plots reveal a clear geometric hierarchy in the intermediate training regime, where memorized samples consistently exhibit the highest curvature (largest singular values), followed by spurious, and finally generalized samples.
+As the model enters the full generalization phase at large $K$, the average spectra for all three sample types become very distinguishable.
+Shaded region is the standard deviation of the singular values of different samples of each category.
+Figure 23: Average singular value spectra of the score’s Jacobian for memorized, spurious, and generalized samples from diffusion models trained on the CIFAR10 [37] dataset, shown across 38 increasing training data sizes $K$.
+In contrast to the clear separation as seen in other datasets, the spectra for all three sample types remain heavily overlapped throughout the entire transition from memorization to generalization.
+However, as the model enters the full generalization phase at large $K$, the average spectra for all three sample types become less indistinguishable.
+Shaded region is the standard deviation of the singular values of different samples of each category.
+D.4  Additional Examples of Stable Diffusion
+Figure 24: Additional examples of candidate memorized, spurious, and generalized samples from Stable Diffusion trained on LAION dataset [54], each corresponding to a distinct curvature signature.
+The candidate memorized sample has much larger singular values, while the candidate generalized sample has much smaller ones.
+The selected possible spurious samples have larger singular values than the candidate generalized sample, but much less in contrast to the candidate memorized sample.
+The common trait between candidate memorized and spurious samples is that both are stable attractors, as demonstrated by their repeated similar generations given different initial points (or noise vectors) and the conditioning on text-prompt.
+The y-axis is clipped at the value of 1500 to better contrast the shown examples’ spectra.
+Appendix E Relative Energy Gap
+E.1  Overview
+In this section, we provide an additional experiment on the relative energy gap, which measures the gap between the energy depth of training data points and each of the three distinct sample types, using the method for approximating the relative potential from [49].
+Consider the reverse SDE of the main text:
+$\mathrm{d}{\mathbf{x}}_{t}=[{\mathbf{f}}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})]\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t}.$ (37)
+We can re-express the above equation with the potential function $u(x,t)$:
+$\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)=\mathbf{f}({\mathbf{x}}_{t},t)-g(t)^{2}\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})$ (38)
 by using the drift term. By moving the minus sign and taking the integration of Eq. (38) with respect to ${\mathbf{x}}_{t}$, we have
-[…]
-which is the potential function $u({\mathbf{x}}_{t},t)$ of the generative process. Substituting this potential function (36) into the reverse SDE (2), we have the following generative process
-$\mathrm{d}{\mathbf{x}}_{t}=-\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (37)
-For our experimental setting, where we use the DDPM [23] setting, we have the following variance preserving forward SDE
-$\mathrm{d}{\mathbf{x}}_{t}=-\frac{1}{2}\beta(t){\mathbf{x}}_{t}\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}{\mathbf{w}}_{t}$ (38)
+$u({\mathbf{x}}_{t},t)=g(t)^{2}\log p_{t}({\mathbf{x}}_{t})+\int^{0}_{{\mathbf{x}}_{t}}f({\mathbf{z}},t)d{\mathbf{z}},$ (39)
+which is the potential function $u({\mathbf{x}}_{t},t)$ of the generative process. Substituting this potential function (39) into the reverse SDE, we have the following generative process:
+$\mathrm{d}{\mathbf{x}}_{t}=-\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)\mathrm{d}t+g(t)\mathrm{d}\bar{{\mathbf{w}}}_{t}.$ (40)
+For our experimental setting, where we used the DDPM [22] setting, we have the following variance preserving forward SDE:
+$\mathrm{d}{\mathbf{x}}_{t}=-\frac{1}{2}\beta(t){\mathbf{x}}_{t}\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}{\mathbf{w}}_{t},$ (41)
 and the corresponding reverse (or generative) process SDE
-$\mathrm{d}{\mathbf{x}}_{t}=\bigg{[}-\frac{1}{2}\beta(t){\mathbf{x}}_{t}-\beta(t)\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg{]}\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (39)
-With Equation 37, we have
-$\mathrm{d}{\mathbf{x}}_{t}=-\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}\bar{{\mathbf{w}}}_{t}$ (40)
-and by using Equation 35, the potential for our experimental setting equates to
-[…]
+$\mathrm{d}{\mathbf{x}}_{t}=\bigg[-\frac{1}{2}\beta(t){\mathbf{x}}_{t}-\beta(t)\nabla_{{\mathbf{x}}_{t}}\log p_{t}({\mathbf{x}}_{t})\bigg]\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}\bar{{\mathbf{w}}}_{t}.$ (42)
+With Eq. (40), we have
+$\mathrm{d}{\mathbf{x}}_{t}=-\nabla_{{\mathbf{x}}_{t}}u({\mathbf{x}}_{t},t)\mathrm{d}t+\sqrt{\beta(t)}\mathrm{d}\bar{{\mathbf{w}}}_{t},$ (43)
+and by using Eq. (38), the potential for our experimental setting equates to
+$\begin{split}u({\mathbf{x}}_{t},t)&=g(t)^{2}\log p_{t}({\mathbf{x}}_{t})+\int_{{\mathbf{x}}_{t}}^{0}f({\mathbf{z}},t)d{\mathbf{z}}\\ &=\beta(t)\log p_{t}({\mathbf{x}}_{t})+\frac{1}{2}\beta(t)\int_{{\mathbf{x}}_{t}}^{0}{\mathbf{z}}_{t}\mathrm{d}{\mathbf{z}}\\ &=\beta(t)\log p_{t}({\mathbf{x}}_{t})-\frac{1}{4}\beta(t){\mathbf{x}}_{t}^{2}.\end{split}$ (44)
 Assume we have two images, ${\mathbf{x}}_{1}(t)$ and ${\mathbf{x}}_{2}(t)$, where ${\mathbf{x}}_{1}$ is a target image and ${\mathbf{x}}_{2}(t)$ is a reference image at time $t$.
@@
 We can obtain such a trajectory via circular interpolation between ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{2}$ following
-$\tilde{{\mathbf{x}}}(\alpha)=\cos(\alpha){\mathbf{x}}_{1}+\sin(\alpha){\mathbf{x}}_{2}$ (42)
+$\tilde{{\mathbf{x}}}(\alpha)=\cos(\alpha){\mathbf{x}}_{1}+\sin(\alpha){\mathbf{x}}_{2},$ (45)
 and compute the relative potential (up to a constant) as a function of $\alpha$ by using
-$\tilde{u}(\alpha)=u(\tilde{{\mathbf{x}}}(\alpha))=\int^{\pi/2}_{0}\nabla u(\tilde{{\mathbf{x}}}(\alpha))\cdot{\mathbf{v}}\,\mathrm{d}\alpha$ (43)
+$\tilde{u}(\alpha)=u(\tilde{{\mathbf{x}}}(\alpha))=\int^{\pi/2}_{0}\nabla u(\tilde{{\mathbf{x}}}(\alpha))\cdot{\mathbf{v}}\,\mathrm{d}\alpha,$ (46)
 where ${\mathbf{v}}=-\sin(a){\mathbf{x}}_{1}+\cos(a){\mathbf{x}}_{2}$.
-See Figure 22 for a visualization of this circular interpolation between two samples.
+See Fig.
+(25) for a visualization of this circular interpolation between two samples.
 For discrete times, we instead estimate the relative energy as
 $\tilde{u}(\bar{\alpha}_{L})\approx\sum^{L}_{i=1}\nabla u(\tilde{{\mathbf{x}}}(\bar{\alpha}_{i}))\cdot{\mathbf{v}}\Delta\alpha$ (47)
-given that $\bar{\alpha}=\{0,\dots,\frac{\pi}{2}\}$ and the potential $u(\cdot)$ is obtained from the above Equation 41.
-Experimentation Details.
-In our experimentation of Section 7, we utilized a maximum of the top 2048 samples for each of three sample types (i.e., memorized, spurious, and generalized sets), and the training data set corresponding at each training data size $K$.
-Note, with the exception of the training data set, the samples in the three sets are sorted from least to greatest, based on their distance corresponding to their respective detection metric, e.g., Eq.
-(13), (14) or (15).
-We computed the relative energy of each set following Algorithm 2, which is strictly following the formulations detailed above.
+given that $\bar{\alpha}=\{0,\dots,\frac{\pi}{2}\}$ and the potential $u(\cdot)$ is obtained from the above Eq. (44).
+E.2  Experimental Details
+In our experimentation, we utilized a maximum of the top 512 samples for each of three sample types, memorized, spurious, and generalized, to compute the relative energy following Eq.
+(47), at each training data size $K$.
+For training data points, we instead utilized $\min(K,2048)$ number of samples.
+Note, with the exception of the training data set, the samples in the three sets are sorted from least to greatest, based on their distance corresponding to their respective detection metric (from the main text).
+We computed the relative energy of each set following Eq.
+(4), which is strictly following the formulations detailed above.
 Note, we randomly selected and utilized an image from the entire training set as part of the relative energy calculation.
-To define the average relative energy gap, let $\tilde{u}_{\text{target}}\in\mathbb{R}^{M}$ be a set of $M$ relative potentials computed from using either a memorized, spurious, or generalized set and following Equation 44.
-Please note that $M$ is capped at $2048$ samples.
-Since at some training data size $K$, certain sample type sets might not have less than $2048$ samples.
-For such cases, we use the entirety of that set to compute the relative energy of its samples.
-Meanwhile, let $\tilde{u}_{\text{data}}\in\mathbb{R}^{M^{\prime}}$ be the set of relative potentials computed from the training dataset at size $K$, where $M^{\prime}=2048$ for $K\geq 2048$.
+To define the average relative energy gap, let $\tilde{u}_{\text{target}}\in\mathbb{R}^{M}$ be a set of $M$ relative potentials computed from using either a memorized, spurious, or generalized set and following Eq.
+(47).
+Please note that $M$ is capped at $512$ for the three sample types.
+Meanwhile, let $\tilde{u}_{\text{data}}\in\mathbb{R}^{M^{\prime}}$ be the set of relative potentials computed from the training dataset at size $K$ where $M^{\prime}=\min(K,2048)$.
 The average relative energy gap is defined as
 $\Delta(\tilde{u}_{\text{target}},\tilde{u}_{\text{data}})=\frac{1}{M}\sum^{M}_{i=1}\big(\tilde{u}_{\text{target}}\big)_{i}-\frac{1}{M^{\prime}}\sum^{M^{\prime}}_{j=1}\big(\tilde{u}_{\text{data}}\big)_{j}=\mu(\tilde{u}_{\text{target}})-\mu(\tilde{u}_{\text{data}})$ (48)
-Using Equation 45, we computed the average relative energy gap and recorded our results in Figs.
-(8) and (23).
-The shaded regions reported in these figures represent the standard deviations computed from using $\tilde{u}_{\text{target}}$, and the average relative energy gap as the mean value.
+Using Eq.
+(48), we computed the average relative energy gap and recorded our results in Fig.
+(26).
+The shaded regions reported in this figure represent the standard deviations computed from using $\tilde{u}_{\text{target}}$, and the average relative energy gap as the mean value.
 Since at some training data size $K$, there might be very few samples in one of the three sets, we opted not to include the relative energy gap of such sets, partly due to their large variance.
 Similar to the basin of attraction experiment, we determine whether a set is too small if the number of its samples is less than $0.1\%$ of its corresponding synthetic set.
-Finally, for our interpolation values $\bar{\alpha}=\{0,\dots,\frac{\pi}{2}\}$ detailed in Algorithm 2 and used in Equation 44, we generate $20$ points starting from 0 to $\frac{\pi}{2}$ using linear spacing.
-Moreover, since the relative energy computation requires a reference image, we studied the effects of having different reference images on the relative energy gap in Figs.
-(24), (25), (26), and (27).
-Overall, there is a slight difference in the values of the relative energy gap across the memorization-generalization transition for all datasets.
-However, the general trend is still maintained.
-Specifically, the gap for memorized samples is initially close to zero and quickly diverges as the training data size $K$ increases.
-Spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase.
-However, the potential of spurious states quickly diverges as the later stages of generalization are initiated and consequently, where the gap between generalized and training data samples’ relative energy significantly decreases.
+Finally, for our interpolation values $\bar{\alpha}=\{0,\dots,\frac{\pi}{2}\}$ used for Eq.
+(47), we generate $20$ points starting from 0 to $\frac{\pi}{2}$ using linear spacing.
 Figure 25: Circular interpolation of a target image ${\mathbf{x}}_{1}$ and a reference image ${\mathbf{x}}_{2}$ over $\alpha\in\{0,\dots,\frac{\pi}{2}\}$ for different datasets.
-H.1 Algorithm for Computing Relative Energy
-1
-2
-3Inputs
-4        Target pattern ${\mathbf{x}}_{1}\in\mathbb{R}^{N}$
-5        Reference pattern ${\mathbf{x}}_{2}\in\mathbb{R}^{N}$
-6        Score model $s_{\theta}$
-7        Interpolation values $\bar{\alpha}$
-8        Initial variance $\beta_{0}$
-9Assumption: ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{2}$ are samples at $t=0$
-10 Compute $\tilde{u}(\bar{\alpha}_{L})$
-11        $\tilde{u}(\bar{\alpha}_{L})\leftarrow 0$
-12
-       $\bar{\alpha}\leftarrow\{0,\cdots,\frac{\pi}{2}\}$
+E.3  Discussion
+Since the relative energy computation requires a reference image, we also studied the effects of having different reference images on the relative energy gap in Figs.
+(27)-(30).
+Overall, the results for the relative energy gap across the memorization-generalization transition for all datasets are shown in Fig.
+(26).
+As the training data size $K$ increases, the gap for memorized samples is initially close to zero and quickly diverges.
+Spurious samples’ relative potential resembles that of the training data samples during the intermediate phase of the memorization-generalization phase, where their potential is the lowest among the three sample types.
+However, the potential of spurious states quickly diverges as the later stages of generalization are initiated, and consequently, the gap between generalized and training data samples’ relative energies became the most narrowed among the three distinct sample types – further emphasizing the changing in the energy landscape as the DM undergoes its transition from memorization to generalization as $K$ increases.
+Inputs
+   Target pattern ${\mathbf{x}}_{1}\in\mathbb{R}^{N}$
+  2.5pt Reference pattern ${\mathbf{x}}_{2}\in\mathbb{R}^{N}$
+  2.5pt Score model $s_{\theta}$
+  2.5pt Interpolation values $\bar{\alpha}$
+  2.5pt Initial variance $\beta_{0}$
+end
+1exAssumption: ${\mathbf{x}}_{1}$ and ${\mathbf{x}}_{2}$ are samples at $t=0$
+1exCompute $\tilde{u}(\bar{\alpha}_{L})$
+  5pt $\bar{\alpha}\leftarrow\{0,\cdots,\frac{\pi}{2}\}$
  // Linear Spacing from 0 to $\frac{\pi}{2}$
-13
-14       $\Delta\alpha\leftarrow\bar{\alpha}_{1}-\bar{\alpha}_{0}$
-15
-16       for $\alpha\in\bar{\alpha}$ do
-17               $\tilde{{\mathbf{x}}}\leftarrow\cos(\alpha){\mathbf{x}}_{1}+\sin(\alpha){\mathbf{x}}_{2}$
-18               $\tilde{{\mathbf{v}}}\leftarrow-\sin(a){\mathbf{x}}_{1}+\cos(a){\mathbf{x}}_{2}$
-19               $\nabla_{\tilde{{\mathbf{x}}}}\log p_{0}(\tilde{{\mathbf{x}}})\leftarrow-s_{\theta}(\tilde{{\mathbf{x}}},0)/\sqrt{\beta_{0}}$
-               $\nabla_{\tilde{{\mathbf{x}}}}u(\tilde{{\mathbf{x}}})\cdot{\mathbf{v}}\,\Delta\alpha\leftarrow\big{(}-\frac{1}{2}\beta_{0}\tilde{{\mathbf{x}}}-\beta_{0}\nabla_{\tilde{{\mathbf{x}}}}\log p_{0}(\tilde{{\mathbf{x}}})\big{)}\cdot{\mathbf{v}}\,\Delta\alpha$
-                // See Equation 39
-20
+  5pt
+ $\Delta\alpha\leftarrow\bar{\alpha}_{1}-\bar{\alpha}_{0}$
+  5pt
+ $\tilde{u}(\bar{\alpha}_{L})\leftarrow 0$
+  8pt
+ for $\alpha\in\bar{\alpha}$ do
+     5pt $\tilde{{\mathbf{x}}}\leftarrow\cos(\alpha){\mathbf{x}}_{1}+\sin(\alpha){\mathbf{x}}_{2}$
+     5pt $\tilde{{\mathbf{v}}}\leftarrow-\sin(a){\mathbf{x}}_{1}+\cos(a){\mathbf{x}}_{2}$
+     5pt $\nabla_{\tilde{{\mathbf{x}}}}\log p_{0}(\tilde{{\mathbf{x}}})\leftarrow-s_{\theta}(\tilde{{\mathbf{x}}},0)/\sqrt{\beta_{0}}$
+     8pt $\nabla_{\tilde{{\mathbf{x}}}}u(\tilde{{\mathbf{x}}})\cdot{\mathbf{v}}\,\Delta\alpha\leftarrow\big(-\frac{1}{2}\beta_{0}\tilde{{\mathbf{x}}}-\beta_{0}\nabla_{\tilde{{\mathbf{x}}}}\log p_{0}(\tilde{{\mathbf{x}}})\big)\cdot{\mathbf{v}}\,\Delta\alpha$
+    // Eq. (42)
     $\tilde{u}(\alpha)\leftarrow\sum^{N}_{j=1}\bigg(\nabla_{\tilde{{\mathbf{x}}}}u(\tilde{{\mathbf{x}}})\cdot{\mathbf{v}}\,\Delta\alpha\bigg)_{j}$
-                // Sum over $N$ pixels
-21
+    // Sum over NN
     $\tilde{u}(\bar{\alpha}_{L})\leftarrow\tilde{u}(\bar{\alpha}_{L})+\tilde{u}(\alpha)$
-                // See Equation 44
-22
-23
-24Output
-25        Relative energy $\tilde{u}(\bar{\alpha}_{L})$
-Algorithm 2 Compute the Discretized Relative Energy between a target ${\mathbf{x}}_{1}$ and a reference ${\mathbf{x}}_{2}$
-H.2 Average Relative Energy Gap of Sample Types for All Datasets
-Figure 23: Average energy gap of each sample type and the training samples for the four datasets as the training data size $K$ grows, computed using Equation 44.
+    // Eq. (47)
+   end for
+end
+Output
+  5pt Relative energy $\tilde{u}(\bar{\alpha}_{L})$
+end
+Algorithm 4 Compute the discretized relative energy between a target and a reference sample
+E.4  Average Relative Energy Gap Results
+Figure 26: Average energy gap of each sample type and the training samples for the four datasets as the training data size $K$ grows, computed using Eq.
+(47).
 Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself.
-A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points.
+A maximum of 512 samples per type is used at each $K$, or the full set is used instead if there are fewer than 512 points.
 Shaded regions show standard deviation of the gap values.
@@
 However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
-H.3 Average Relative Energy Gap using Different Reference Images for All Datasets
-Figure 24: Average energy gap of each sample type and the training samples of CIFAR10 as the training data size $K$ grows, computed using Equation 44 and four different reference images.
+E.5  Average Relative Energy Gap Results with Different References
+Figure 27: Average energy gap of each sample type and the training samples of CIFAR10 [37] as the training data size $K$ grows, computed using Eq.
+(47) and four different reference images.
 Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself.
-A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points.
+A maximum of 512 samples per type is used at each $K$, or the full set is used instead if there are fewer than 512 points.
 Shaded regions show standard deviation of the gap values.
@@
 However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
-Figure 25: Average energy gap of each sample type and the training samples of LSUN-CHURCH as the training data size $K$ grows, computed using Equation 44 and four different reference images.
+Figure 28: Average energy gap of each sample type and the training samples of LSUN-CHURCH [75] as the training data size $K$ grows, computed using Eq.
+(47) and four different reference images.
 Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself.
-A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points.
+A maximum of 512 samples per type is used at each $K$, or the full set is used instead if there are fewer than 512 points.
 Shaded regions show standard deviation of the gap values.
@@
 However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
-Figure 26: Average energy gap of each sample type and the training samples of FASHION-MNIST as the training data size $K$ grows, computed using Equation 44 and four different reference images.
+Figure 29: Average energy gap of each sample type and the training samples of FASHION-MNIST [73] the training data size $K$ grows, computed using Eq.
+(47) and four different reference images.
 Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself.
-A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points.
+A maximum of 512 samples per type is used at each $K$, or the full set is used instead if there are fewer than 512 points.
 Shaded regions show standard deviation of the gap values.
@@
 However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
-Figure 27: Average energy gap of each sample type and the training samples of MNIST as the training data size $K$ grows, computed using Equation 44 and four different reference images.
+Figure 30: Average energy gap of each sample type and the training samples of MNIST [17] as the training data size $K$ grows, computed using Eq.
+(47) and four different reference images.
 Gaps are measured relative to training samples, whose energy is set to zero since there is no gap to itself.
-A maximum of 2048 samples per type is used at each $K$, or the full set is used instead if there are fewer than 2048 points.
+A maximum of 512 samples per type is used at each $K$, or the full set is used instead if there are fewer than 512 points.
 Shaded regions show standard deviation of the gap values.
@@
 However, the potential quickly diverges as the later stages of generalization are initiated and consequently, the gap between generalized and training data samples’ relative energy significantly decreases.
+Appendix F Hardware Details
+The training of the models detailed in Table (3) was done using NVIDIA Tesla V100 GPUs, where a single GPU was used to train each model.
+Each GPU has 32GB of memory and is linked with Power9 processors, clocking at 3.15 GHz maximum.
+Each CIFAR10 model took roughly 18, 20, and 22 hours to train for different initial latent dimensions, e.g., 64, 96, 128, respectively.
+Meanwhile, each LSUN-CHURCH model requires roughly 48, 56, and 103 hours of training with respect to the model’s initial latent dimension.
+In contrast, FASHION-MNIST and MNIST models require approximately 17 hours to be fully trained (per model).
+Please note that these durations are with respect to the required training iterations for each model and their corresponding dataset detailed in Table (3).
+Also, note that we utilized the same hardware for our analyses.